### v1

# Can LLM-Simulated Practice and Feedback Upskill Human Counselors? A Randomized Study with 90+ Novice Counselors

Abstract
Training more counselors, from clinical students to peer supporters, can help meet the demand for accessible mental health support; however, current training approaches remain resource-intensive and difficult to scale effectively. Large Language Models (LLMs) offer promising solutions for growing counseling skills training through simulated practice and automated feedback. Despite successes in aligning LLMs with expert-counselor annotations, we do not know whether LLM-based counseling training tools—such as AI patients that simulate real-world challenges and generative AI feedback with suggested alternatives and rationales—actually lead to improvements in novice counselor skill development. We develop CARE, an LLM-simulated practice and feedback system, and randomize 94 novice counselors to practice using an AI patient, either alone or with AI feedback, measuring changes in their behavioral performance, self-assessments, and qualitative learning takeaways. Our results show the practice-and-feedback group improved in their use of reflections and questions (d=0.32-0.39, p$<$0.05). In contrast, the group that practiced with an AI patient alone did not show improvements, and in the case of empathy, actually had worse uses across time (d=--0.52, p=0.001) and when compared against the practice-and-feedback group (d=0.72, p=0.001). Participants’ qualitative self-reflections revealed key differences: the practice-and-feedback group adopted a client-centered approach involving listening to and validating feelings, while the practice-alone group remained solution-oriented but delayed offering suggestions until gathering more information. Overall, these results suggest that LLM-based training systems can promote effective skill development, but that combining both simulated practice and structured feedback is critical.
In 2023, 22.8% of U.S. adults (approximately 58.7 million people) experienced a mental illness [1]. Unfortunately, the ability to provide people with effective mental health care is severely limited by shortages of high-quality mental health providers—from psychotherapists and counselors to social workers and peer supporters [2, 3]. While there is increasing interest in exploring direct-to-patient AI systems for some circumstances, and some promising results [4], we expect the demand for human therapists and mental health support to continue to far exceed demand. The limited supply of effective therapy providers is due—at least in part—to the reliance on resource-intensive methods to train helping skills [5] and evidence-based interventions [6, 7]. Traditional methods require access to trainers who can simulate a client interaction [8, 9] and provide expert supervision [10, 11], creating significant barriers to growing high-quality training [12, 13].
Fortunately, AI systems have been increasingly applied to counselor training as a potential solution to these scaling challenges. Recent advances in large language models (LLMs) have enabled the development of LLM-based systems for simulating patients seeking mental health support [14, 15], offering rich opportunities for practice. Simultaneously, AI feedback systems have progressed in automatically assessing counselor behaviors such as empathy, reflections, and active listening [16, 17, 18, 19] and generating suggested responses [20, 21, 22] and explanations [23, 22]. These AI feedback systems target skills from client-centered approaches [24, 25]—empathy, reflections, questions, and active listening—which are relevant to all mental health providers. Meta-analyses have shown these skills strengthen common factors like therapeutic alliance, which are powerful predictors of therapy outcomes, regardless of the eventual evidence-based intervention or therapy modality implemented [26, 27, 28]. In addition to technical advances, prior work evaluating AI feedback systems with novice counselors have tested their effectiveness when used as real-time co-pilots rather than as training systems [29, 21], or evaluated pre-LLM training systems which had limited practice realism and simpler, non-generative feedback mechanisms [30]. Therefore, we have limited evidence explaining how modern, LLM-based training systems—which offer practice simulations resembling challenging real-world behaviors and generative feedback with suggested alternatives and rationales—promote counselor skill development [31, 32]–a holistic concept that includes appropriate use of skill behaviors, accurate self-awareness of one’s capabilities and limitations [33], and an ability to reflect on and learn from experiential practice [34].
To address this gap, we develop CARE, an LLM-based training system for novice counselors, and conduct a randomized experiment to investigate how two different modes of simulated training impact counselor skill development. CARE enables (1) realistic practice with LLM-simulated patients prompted by expert counselors to resemble challenging behaviors [15], and (2) structured feedback from a fine-tuned LLM that identifies strengths and areas for improvement across core counseling skills (e.g., empathy, reflections, questions, suggestions), while also providing explanatory rationale and alternative responses [22]. Importantly, CARE’s feedback evaluates not only when a counseling skill is used but also how well it is used, distinguishing between strong implementations (e.g., reflections that accurately capture client emotions) and implementations needing improvement (e.g., reflections that miss core client concerns). These feedback evaluations are based on established counseling frameworks [5, 35] and expert-annotated examples.
Figure 1: Our between-subjects study randomizes participants to either practice with AI Patients alone (P) or practice with AI patients and receive AI feedback (P+F). We holistically evaluate counselor skill development from three perspectives: automatic assessments of behaviors of skills used via LLM classifiers; self-efficacy and its calibration with actual performance; and qualitative self-reflections after the training intervention chat and post-intervention chat.
We conducted a 75-minute online lab study with novice counselors (N=94) to empirically evaluate how different LLM-simulated practice modes in CARE impact skill development. Participants were randomly assigned to one of two conditions: (1) Group P: Practice with LLM-simulated patients without AI feedback, or (2) Group P+F: Practice with LLM-simulated patients plus AI feedback (see Fig. 1). To comprehensively study learning outcomes, we measured changes from pre-intervention to post-intervention across multiple dimensions: actual behavioral performance (calculated via automatic analysis of counseling transcripts), participants’ self-efficacy (perceived confidence in counseling skills), and metacognitive awareness (captured through guided self-reflection prompts). Our study design specifically addressed: What changes occur after practice with an AI-simulated patient alone? How do these outcomes differ when participants also receive structured AI feedback?
Our findings reveal that practice with feedback significantly improves empathetic listening skills, while practice alone shows minimal improvement in these core competencies. Notably, both groups failed to improve the accuracy of participants’ self-efficacy assessments, highlighting the challenges involved in helping learners develop a well-calibrated understanding of their own performance. Through analysis of participants’ self-reflections, we found the P+F group internalized the importance of empathetic and active listening, but P participants continued to overly focus on solutions, albeit with increased information-gathering. These insights suggest that future LLM-based training systems must integrate feedback mechanisms to cultivate the empathetic stance fundamental to effective counseling.
Stages Skill Category Description
Questions Questions seek information from the client and can be open (inviting elaboration) or closed (requesting specific answers). They include both direct questions and indirect prompts (e.g., ”Tell me about…”).
Reflections Reflections capture and return to clients something they have communicated, either explicitly or implicitly. They typically mirror back content from the client’s preceding statement, but can also reference earlier parts of the conversation.
Empathy Empathy can be shown through emotional warmth, interpretation of the client’s experience (e.g., paraphrasing, making conjectures, or sharing relatable experiences), or exploration of the client’s feelings and perspectives.
Suggestions Suggestions offer possible actions, perspectives, or solutions in a respectful and autonomy-supportive manner. They may involve information-sharing or proposing alternative viewpoints.
Session Management Session management includes organizing the session, transitioning between topics, and summarizing key points. It provides structure and helps maintain therapeutic focus.
Table 1: Our analysis of skill development has a theoretical-foundation in Hill’s Helping Skills model [5]. We select a skill subset relevant for beginning counselors at the undergraduate and first-year graduate level [36]. These include microskills during the exploration and action stages; and macro skills that are applicable throughout the session. Hill’s insight stage, of which self-disclosure was the only relevant skill for basic counseling, was excluded from our primary analyses due to its infrequent occurrence in our data.
1 Results
This section presents findings from our investigation of how LLM-based practice affects counselors’ skills development. To understand the impact of simulated practice alone (P) and practice with feedback (P+F), we integrate evidence from three complementary data sources.
Behaviors of Skills Used: Training systems should help improve the use of skills that align with evidence-based counseling practice. Thus, we analyzed the skills used in transcripts and how they change between the pre-intervention and post-intervention assessments using high-accuracy LLM classification models. We assessed behaviors of skills used for the exploration stage (strong uses in empathy, reflections, questions) and action stage (suggestions needing improvement) of Hill’s Helping Skills framework [5]; see Table 1 for definitions.
Self-Efficacy Measures: It is informative to understand participants’ confidence and their awareness of their own strengths and limitations. We measured counselor self-efficacy using the Counselor Activity Self-Efficacy Scale [37], which covers confidence for exploration skills, action skills, and session management skills (Table 1). Our analysis tests whether the lowest performing individuals overestimate their abilities (aka the Dunning-Kruger effect [38]) and whether LLM-training interventions improve the accuracy of their self-assessments and raw self-efficacy scores.
Qualitative Data on Learner Takeaways: LLM-simulated training provides opportunities for experiential learning opportunities [39] whereby reflection on action [34] can support updates to one’s knowledge. We also asked participants’ open-ended self-reflections following the practice intervention and post-intervention assessment, and qualitatively coded their responses.
We integrate these three data sources to provide a holistic view of how AI-based practice—with and without AI feedback—influences changes in counselors’ use of skills, self-efficacy, and key lessons learned.
1.1 Changes in Counseling Skill Behaviors
Our primary analysis uses LLM-based behavioral assessments of transcripts to evaluate whether skill usage changes from pre-intervention to post-intervention for both the practice and practice-and-feedback groups. In a secondary analysis, we investigate how exposure to good examples of skills within the AI feedback may explain any changes we observe.
Change after P Change after PF Differences in
change after P vs. P+F
[…]
Empathy ($\uparrow$) -9.6 0.001 -0.52 5.4 0.117 0.23 15 0.001 0.72
Reflections ($\uparrow$) 1.4 0.391 0.18 3.7 0.034 0.32 2.3 0.323 0.2
Questions ($\uparrow$) 2.4 0.421 0.12 6.6 0.018 0.36 4.1 0.296 0.22
Suggestions ($\downarrow$) -5.9 0.010 -0.39 -5.5 0.057 -0.28 0.4 0.910 0.02
Figure 2: Changes in counseling behaviors following AI patient simulations alone (P) versus AI patient simulations with AI feedback (P+F). The plot displays bootstrapped means for pre-intervention and post-intervention interactions. The table presents statistical comparisons with corresponding effect sizes, with bolded values indicating significance after Benjamani-Hochberg correction [40] for the 21 planned comparisons (12 for behavioral changes and 9 for self-efficacy changes). Notably, P group experiences a significant drop in strong uses of Empathy (-9.6% change, $d=-0.52$), whereas P+F group’s use of Empathy is maintained and trends towards improvement. The P+F group also experiences noticeable improvements in Reflections (+3.7% change, $d=0.32$) and Questions (6.59% change, $d=0.36$)
We find that practice alone is not enough; feedback during practice is necessary to promote desirable counseling behaviors in empathetic and active listening. AI feedback during practice (P+F) led to improvements in Reflections (+3.6% change, $p=0.034$) and Questions (+6.59% change, $p=0.018$), and trended towards improvement in Validation (+3.6% change, $p=0.083$), Suggestions (-5.45% change, $p=0.057$) and Empathy (+5.37% change, $p=0.117$). In contrast, practice alone (P) led to only improvements in Suggestions (-5.85% change, $p=0.011$), but no improvements in Reflections, Questions, and Validation, and significantly worse expression of Empathy (-9.6% change, $p<0.001$). Pairwise comparisons between the two conditions were substantial and significant for Empathy (15% relative difference, Cohen’s $d=0.72,p<0.001$) and non-significant for others.
To better understand AI feedback’s role, we further analyzed whether the level of exposure to specific feedback during practice affected these changes. Our analysis quantifies the exposure to Good Alternatives during Practice (GAP), defined as the number of alternative responses in the AI-generated feedback that exhibited strong uses of a skill, normalized by chat length. We hypothesized that greater exposure to GAP for a skill would lead to stronger use of that skill. To test this, we fit a lagged-linear regressor that predicts post-chat behaviors, controlling for pre-chat behaviors, with exposure to good alternatives ($T_{GAP}$) as a scalar predictor ($b_{1}=\beta_{0}+\beta_{1}b_{0}+\beta_{2}T_{GAP}$).
[…]
Figure 3: Counselor Self Efficacy (perceived ability to use skills) for participants grouped by behaviors of skills used (actual performance). Notes: Gaps depict miscalibration between actual and self-assessed percentile of performance for quartile groups with bootstrapped 95% CIs. We only visualize for data collected in the post-intervention.
1.2 Self-Efficacy and its Miscalibration with Behaviors
Our investigation of self-efficacy involved three types of analyses. First, we examine the calibration of self-efficacy ratings with actual performance and evaluate whether P or P+F interventions improve this. Second, we tested for significant changes in raw self-efficacy scores after practice or practice-and-feedback.
Our analysis comparing self-efficacy ratings with actual performance across skill quartiles finds support for the Dunning–Kruger effects more substantially for action skills and to a lesser degree for exploration skills. The interaction between measure and quartile was significant in four out of six ANOVAs for action skills, while only one out of six was significant for exploration skills (Table 4). Pairwise comparisons also showed a pattern indicative of a Dunning–Kruger effect (see Table 5, Table 6 and Fig. 3): People in the lowest quartile overestimated themselves the most. Those in the highest quartile—and to a lesser degree also those in the second-to-highest quartile—tended to underestimate themselves.
Change after P Change after PF Differences in
change after P vs. PF
[…]
Exploration Skills 0.36 0.004 0.44 0.13 0.338 0.14 -0.21 0.238 -0.25
Action Skills 0.27 0.166 0.21 0.33 0.026 0.34 0.04 0.884 0.03
Session Management 0.35 0.021 0.35 0.36 0.011 0.39 0.01 0.955 0.01
Figure 4: Changes in raw-scores of self-efficacy following AI patient simulations alone (P) versus AI patient simulations with AI feedback (PF). The plot displays bootstrapped means for pre-intervention and post-intervention. The table presents statistical comparisons with corresponding effect sizes, with bolded values indicating significance after Benjamani-Hochberg correction [40] for the 21 planned comparisons (12 for behavioral changes and 9 for self-efficacy changes).
Participants’ Ability to Self-Assess Their Skill Level Remained Mixed After LLM Practice. For the practice only (P) group, the mean discrepancy in exploration skills changes from 11.6 percentile underconfidence to 5.7 percentile overconfidence, a significant shift ($p<0.001,d=0.58$). Besides this, we found no other significant changes in calibration. No significant differences was found among the P group for discrepancy in action skills ($p=0.227,d=0.18$). The P+F group showed no significant calibration changes for exploration skills ($p=0.479,d=-0.10$) or for action skills ($p=0.393,d=0.13$). Finally, between-group differences were not significant for discrepancy in exploration skills ($p=0.191,d=0.27$) or action skills ($p=0.743,d=0.07$).
We do find modest overall increases in self-efficacy after P and P+F interventions, with different patterns of improvement across skills (Fig. 4). For the P group, confidence in exploration skills showed a significant increase ($0.36$ points on an 8 point scale, $d=0.44,p=0.004$). Confidence in session management skills showed a significant increase for the P+F group ($0.36$ points, $d=0.39,p=0.011$). While session management skills for the P group also trended towards improvement, it was not significant after correcting for multiple hypothesis testing ($0.35$ points, $d=0.35,p=0.021$). Similarly, while confidence in action skills for the P+F group also increased, this result was not significant after correction of multiple hypothesis tests ($0.33$ points, $d=0.34,p=0.026$). Finally, we found no significant differences between participants who received AI feedback (P+F) versus those who did not (P) (across the three self-efficacy subscales, $d=-0.25,0.03,0.01$, $p=0.238,0.884,0.955$).
1.3 Qualitative Self-Assessments after AI practice with and without AI feedback
To understand participants’ learning process during CARE practice, we analyzed qualitative self-reflections from two time points: immediately after the training intervention chat, where participants responded to “What would you do differently as a therapist?” and after the post-intervention chat, where they reflected on “What did you do well as a therapist?”. We examined how initial intentions translated into reported strengths across the P+F and P groups. Two high-level themes emerged in how participants approached skill improvement:
(1) The P+F group expressed greater intentions and successes in improving their use of empathy and listening skills. P+F participants reported effectively using empathy (27%), validation (27%), and open-ended questions (52%). They emphasized the value of listening skills, such as reflective responses to signal understanding: “I should rephrase what they say so they know I’m understanding them” (P51). They also recognized that counseling should support client exploration of thoughts and emotions, rather than provide direct solutions. One participant reflected on this shift: “I asked them to expand on their feelings, rather than guiding them to my idea” (P39). AI feedback encouraged this shift toward providing emotional support and fostering client autonomy, helping counselors adopt a more empathetic and client-centered approach.
(2) The P participants remained solution-oriented but changed their approach to first gather information. While P+F participants intended to give fewer suggestions, many P participants continued to view suggestions as a central skill. In the post-intervention, 48% of P participants reported using suggestions successfully, compared to only 14% of P+F participants. Many P participants justified their continued use of suggestions by citing a desire to provide tangible, actionable help, for example, “Maybe because I’m untrained and solution oriented. I do not want to leave them with nothing, and nowhere to go” (P40). Some reflected on modifying how they delivered suggestions, emphasizing strategies like gathering more information to tailor advice or providing more concrete guidance. However, in the absence of feedback, most remained fixed in their approach, with some even reporting efforts to rephrase the same solution repeatedly to persuade the client.
2 Discussion
This study investigated the impact of practicing with an LLM-simulated patient either with or without receiving LLM-generated feedback on counselor skills development, resulting in three main findings. First, our behavioral assessments showed that practice with feedback improves empathetic listening skills, while practice alone shows minimal improvement, as evidenced by decreased use of empathy. Second, our qualitative analysis of self-reflections revealed distinct skill development strategies, with feedback recipients more frequently reporting the adoption of client-centered approaches focused on showing empathy and exploring patients’ thoughts and feelings, while practice-only participants gravitated toward solution-oriented approaches focusing on gathering more information and providing suggestions. Both these findings highlight that the development of counseling skills requires not only practice opportunities but also structured feedback that guides learners toward empathetic, client-centered approaches. Finally, participants demonstrated poor calibration between their perceived abilities and actual performance, as evidenced by overestimates of self-efficacy for the lowest quartile performers. This underscores how self-efficacy measures may not reliably indicate skill development. Each of these findings merits further discussion.
Our findings demonstrate that teaching counselors to implement client-centered approaches requires more than just simulated practice opportunities—it requires targeted feedback that guides novices away from their natural solution-oriented tendencies. The fully-featured version of CARE–combining LLM-simulated patient practice and LLM-based feedback–helped participants improve their use of empathetic and active listening skills, with notable improvements in questions ($d=0.36$) and reflections ($d=0.32$). In comparison, practice with an AI patient alone only led to fewer inappropriate suggestions ($d=-0.39$), but no improvements in reflections or questions, and significantly worse uses of empathy (-9.6% change, $d=-0.52$; 15% relative difference to P+F, $d=0.72$). These effect sizes are comparable to those found in studies of human supervision during standardized roleplays, where Maaß et al. (2025) [41] reported observer-rated skill improvements with effect sizes ranging from $d=0.29-0.49$. Since LLM-simulated practice and feedback are not bottlenecked by the resource constraints of human trainers, AI training systems like CARE show promise in scaling access to effective counseling training.
While the combination of feedback and simulated practice was key to CARE’s success, we also recognize that CARE represents just one approach to providing expert-aligned LLM feedback to learners. CARE allows trainees to request feedback on responses they have already sent. This approach allows trainees to experience productive learning discomfort by trying and making mistakes in simulation, while still offering feedback upon request to refine future responses. This design choice parallels elements of both live human supervision [42] and delayed supervision after a session [41], creating a hybrid approach that may offer flexibility based on trainee preferences. Given the promise of feedback combined with practice, the AI for psychotherapy field should evaluate different AI-feedback designs against each other. Several other possible strategies for giving feedback include just-in-time skill suggestions [21, 43], counterfactual simulations [44], and global session feedback. While conducting randomized trials in medical education research can be tricky [45], randomized trials delivering different training interventions to large participant-pools for longer durations [46] could provide essential data to determine optimal approaches for different learning objectives and learner characteristics.
Our study revealed that participants’ self-efficacy ratings were poorly calibrated with their actual performance, especially among lower performers. This finding, consistent with prior research, suggests that self-assessment accuracy alone may not be a reliable indicator of counselor competence or development. Recent reviews indicate that efforts to improve self-assessment calibration have limited impact on learning or performance outcomes [47]. Instead, it is valuable to objectively assess specific standards of performance and skill use [42, 48] and design interventions that can help low performers improve on those metrics while maintaining a positive morale for continued practice. As AI-based training tools evolve, integrating objective performance measures and structured self-reflection, rather than relying solely on self-assessment, offers a more robust approach to supporting counselor development.
Several limitations of this study warrant acknowledgment. First, our assessment focused primarily on immediate skill acquisition rather than long-term retention or transfer to actual clinical settings. Second, while our automated behavioral analysis focused on microskill usage, it could not capture observable session-level characteristics. While traditional studies have employed human observers to provide such ratings, recent AI research has shown the validity of using finetuned LLMs in certain contexts to approximate session-level measures, such as working alliance [49], which might enable scalable behavioral analyses of more skill development constructs. Third, we recruited from a sample of bachelor-level counselors through the online Prolific platform, which may differ from in-person or classroom settings. For example, existing training environments for counselors will typically offer didactic materials (e.g., lectures and textbooks) and non-AI training opportunities (e.g., peer-to-peer roleplay and feedback). Thus, future work should evaluate the utility and acceptability of AI-based training tools within existing educational and training pathways. Finally, the efficacy of LLM-based practice and feedback training was only shown for client-centered microskills. Future work is required to understand whether these results generalize to other therapy modalities (e.g., psychodynamic, CBT/ACT), where training interventions may work to shift natural tendencies in similar or subtly different ways.
3 Conclusion
Despite increasing interest in using LLMs in mental health, to our knowledge, this is the first study to conduct a large-scale evaluation ($N=94$) of an LLM-based training system for developing core skills in novice counselors. Our findings show that while simulated practice (P) proved insufficient—and in the case of empathy, potentially harmful—practice with AI-generated feedback (P+F) supported measurable improvements in key counseling skills of demonstrating empathy, delivering reflective responses, and engaging in client-centered inquiry. By combining realistic patient simulations with expert-aligned, skill-specific feedback, the CARE AI Training system can help novices to master skills involved in client-centered therapy, offering a promising avenue for scalable, evidence-aligned training in mental health care.
4 Methods
4.1 Overview of CARE Training Platform
Figure 5: CARE’s practice and feedback model visualized in a diagram and web screenshot. In CARE, counselors practice with an LLM-simulated patient and receive feedback on each of their responses. The feedback model labels whether a response has strengths or constructive feedback areas. Responses with constructive feedback explain what the goal should be at this point in the conversation; what a helper could improve to better align with this goal; and how they could respond differently via an alternative response.
We developed CARE as a web platform for novice counselors to train in text-based counseling skills enabled by LLMs. Trainees practice simulated conversations with AI patients and review feedback on their responses generated by an AI mentor. CARE builds on top of successes from previous research in co-designing with mental health experts to improve the realism of LLM-simulated patients [50, 43, 14, 15] and finetuning domain-specific LLMs with therapeutic knowledge capable of generating feedback and alternative responses for text-based peer counseling conversations [51, 22, 52, 53, 54].
CARE allows novice counselors to develop their counseling skills in a text-based format by practicing with AI-simulated patients and receiving feedback on their responses (see Fig. 5).
Consider Aki, a novice peer counselor who wants to use CARE to experience hands-on training using counseling skills that they have recently read about. In CARE, Aki can practice with an AI patient of their choice from a library of patient scenarios. Aki starts a practice chat with an AI patient with the background of a 35-year-old male veteran seeking to reconnect with his children but facing legal barriers and parental gatekeeping.
Each practice scenario provides limited background information about the AI patient and their presenting problem (e.g., ”Young adult with family issues: low mood and self-esteem”). This intentional limitation requires counselors to simultaneously learn more about the patient’s situation while demonstrating empathy and support.
Aki starts the conversation by greeting the AI patient. The AI patient shares their frustration with feeling everyone is against him, and hopes to find ways get his kids back and overcome the difficulty his judgmental parents. Aki composes a reply, and the conversation continues in this turn-by-turn manner.
Users can view feedback on their therapeutic responses at any point during or after the session, offering flexibility to either review intermittently throughout the practice or comprehensively after completion.
4.2 Implementation details for CARE Training Platform
CARE was built as a web application with a Python Flask back-end and React Javascript front-end which online participants could access and interact with. LLM-simulated patients in CARE were powered by the GPT-4o API with a custom prompt defining the patient’s background scenario and behavioral principles [15]. The prompts for the three AI patients used for the pre-intervention assessment, intervention, and post-intervention are provided in the Supplementary Materials A.1. The LLM feedback system was powered by a Llama-2 13B parameter model finetuned on an expert-annotated feedback dataset [22].
4.3 Study Procedures
Between August and October 2024, we conducted an online lab study with (N=94) novice counselors. Given our aim to train novices lacking formal counseling training, we recruited participants who expressed an interest in practicing helping skills yet had minimal prior education or practical experience. Specifically, we used the Prolific platform to screen for participants who (1) lacked extensive formal professional or graduate training in clinical psychology, counseling, social work, or nursing (excluding those who were pursuing or had completed master’s or doctorate degrees); and (2) had less than one year of counseling-related experience (e.g., peer support or crisis counseling). In our final experiment, we recruited N=94 participants. Additionally, during the pilot phase, we recruited another N=14 participants who were allowed broader educational and hands-on experiences for protocol refinement, but we report qualitative and quantitative results exclusively from the final set of 94 participants.
Each participant completed a 75-90 minute video-call session following the structure outlined in Fig. 1. After receiving a static 5-minute tutorial on core counseling skills (see Fig. 6), all participants completed a 10-minute pre-intervention assessment chat with the first AI patient. For the 20-minute training intervention, participants were randomly assigned to either practice with an AI patient only or practice with the option to receive feedback on all of their sent responses. Finally, all participants completed a 10-minute post-intervention assessment chat with a third AI patient. Participants answered self-efficacy items after the pre-intervention and post-intervention chats, and provided qualitative self-reflections after the training intervention chat and post-intervention chat. All participants were paid \$15/hour. To protect participant privacy, we instructed them to join the video-call using their ProlificID as the display name, turn off their video-camera, and share their screen. Participants consented for the video call to be recorded for our own analysis and their CARE usage data to be used and shared for the purposes of this and future research after de-anonymization. This study protocol was approved by our institution’s IRB.
4.4 Assessing Changes in Skill Use through LLM Classifiers
We investigate whether counselors employ higher-quality counseling behaviors after LLM-based training and between training conditions by leveraging NLP methods. This analysis is motivated by the need to quantify changes in counseling skill use at scale across multiple participant sessions. Our approach follows four key steps: (1) fine-tuning and validating LLM-based classifiers to identify skill behaviors, (2) selecting a final set of classifiers based on performance metrics and theoretical priority, (3) testing changes in skill use across time and between intervention groups, and (4) analyzing the relationship between exposure to good alternative patterns in AI feedback and subsequent skill use.
4.4.1 Fine-tuning and Validating LLM-based Classifiers
We developed LLM-based binary classifiers that allows us to label the skill use within a transcript. For example, one of our fine-tuned classifier could determine which utterances in a transcript had strong uses of Questions during that stage of the transcript. To finetune and evaluate these classifiers, we transformed a previously published expert-annotated, feedback dataset that had strengths and areas of improvements for 8 skills [22] into a 16-class binary classification format (8 skills × 2 categories each—strong uses and areas needing improvement) [^1] The binary classification feedback dataset can be accessed at https://huggingface.co/datasets/youralien/feedback_qesconv_16wayclassification. Additionally, we used a subset of transcripts from this randomized study which were further annotated by counseling domain-experts.
Three experts were recruited with relevant backgrounds including practicing clinical psychologist, licensed marriage family therapist, former director and supervisor of a crisis agency. Each of the experts annotated 10 participants’ study transcripts (5 from the practice-only group; 5 from the practice-and-feedback group), totaling 370 counselor utterances. Two rounds of annotation occurred: after collecting a first annotation pass and identifying data points with disagreements, we showed each of the experts the others’ annotations and had them re-annotate and provide rationales for their decision. We display pairwise agreement results averaged across all pairs in Table 2. Note that while we originally explored other annotation agreement metrics such as Cohen’s kappa, the severe class imbalance of our annotations made these metrics less relevant in our case. The gold-standard CARE 10% expert-annotated sample consists of labels that result from a majority vote across these three experts [^2] This expert-annotated data sample can be found at https://huggingface.co/datasets/youralien/CARE_10percent_16wayclassification.
We finetuned RoBERTa-large binary classifiers using a 95% split of the transformed FeedbackQESConv dataset, and did hyperparameter tuning on 5% of FeedbackESConv and the CARE 10% sample. The performance of the our LLM-based classifiers are shown in Table 2. The highest performing classifiers ($F1>0.5$) became candidates for our automatic behavioral assessments, which we further down-selected as described below.
Skill Strengths Areas to Improve
Annotator Classifier Annotator Classifier
Agreement Performance Agreement Performance
% acc. f1 % acc. f1
Empathy 0.793 0.813 0.741 0.809 0.859 0.389
Reflections 0.863 0.900 0.562 0.944 0.903 0.312
Questions 0.732 0.784 0.775 0.852 0.842 0.394
Suggestions 0.919 0.955 0.507 0.946 0.941 0.681
Validation 0.726 0.852 0.556 0.919 0.893 0.265
Self-disclosure 0.982 0.920 0.326 0.969 0.986 0.849
Session Management 0.968 0.941
Professionalism 0.905 0.969
Table 2: Left columns show pairwise annotator-agreement averaged across 3 domain-experts for the CARE 10% sample (n=370). Right columns show performance of the best RoBERTa-large classification models after hyperparameter tuning on our validation dataset, CARE 10% sample (n=370) + FeedbackQESConv 5% sample (n=409). Session Management and Professionalism were excluded from finetuning due to infrequent occurrence.
4.4.2 Down-selecting a Final Set of Classifiers
From the initial set of 16 binary classifiers, we applied both methodological and theoretical criteria to select our final set for analysis. First, we established a minimum performance threshold of $F1>0.5$ to ensure reliable classification. This criterion yielded seven candidate classifiers: strong uses of Empathy, Reflections, Questions, and Validation, as well as both strong uses and areas needing improvement for Suggestions.
To maintain statistical power while controlling for multiple comparisons, we further narrowed our focus to four key classifiers: strong uses of Empathy, Reflections, and Questions, plus areas needing improvement for Suggestions. This selection was guided by three considerations: (1) the need to limit the number of statistical comparisons to avoid diluting significance across too many tests, (2) focusing on classifiers with the strongest performance metrics, and (3) selecting skills that represent core competencies in client-centered frameworks and are frequently used in counseling sessions.
Self-disclosure was excluded from our analysis due to its infrequency in our dataset, while Validation, though conceptually related to Empathy and mentioned frequently in qualitative data, showed more limited classifier performance and was therefore reserved for secondary analyses.
4.4.3 Testing Changes across Time and between Groups
Using our selected classifiers, we examined how skill usage changes from the pre-intervention chat to the post-intervention chat. For each session transcript, we computed an aggregate score defined as the proportion of utterances with strong uses of a skill (e.g., $b=N_{strengths}/N_{total}$) or with areas needing improvement (e.g., $b=N_{improvement}/N_{total}$).
To statistically test for changes between the pre-intervention and post-intervention chats ($b_{0},b_{1}$), we used paired t-tests and computed Cohen’s $d$ effect sizes. To test for significant differences between P and P+F groups ($b_{1}^{P}-b_{0}^{P}$ vs. $b_{1}^{PF}-b_{0}^{PF}$), we used non-paired t-tests.
In total, we conducted 12 planned t-tests for analyzing changes in behaviors: three skills (Empathy, Reflections, Questions) with strong uses and one skill (Suggestions) with areas needing improvement, each analyzed for both within-group changes and between-group differences in those changes.
4.4.4 Exposure to Good Alternatives in AI Feedback
To better understand AI feedback’s role in skill development, we analyzed whether the level of exposure to specific feedback during practice affected post-intervention performance. We quantified the exposure to Good Alternatives during Practice (GAP), defined as the number of alternative responses in the AI-generated feedback that exhibited strong uses of a skill, normalized by chat length.
To test whether greater exposure to GAP for a skill led to stronger use of that skill, we fit a lagged-linear regressor that predicts post-chat behaviors, controlling for pre-chat behaviors, with exposure to good alternatives ($T_{GAP}$) as a scalar predictor ($b_{1}=\beta_{0}+\beta_{1}b_{0}+\beta_{2}T_{GAP}$). This model allowed us to isolate the effect of feedback exposure while accounting for baseline performance differences.
4.5 Analyzing Confidence and the Calibration of Self-Assessments using Counselor Self-Efficacy
Counselor self-efficacy represents a critical dimension in counselor development, defined as one’s “beliefs or judgments about their ability to effectively counsel a client” [55]. While high self-efficacy can promote persistence and resilience when facing challenges [56], miscalibrated self-assessments—particularly overconfidence relative to actual ability—can lead to suboptimal or potentially harmful counseling behaviors [57, 58, 59].
This calibration challenge has prompted research into potential remedies. Prior studies suggest that increased practice experience may help individuals better align their self-assessments with actual performance, though results remain inconsistent [60]. Similarly, supervisor feedback has been proposed as a mechanism to address self-assessment miscalibration [61], although evidence from other educational contexts shows mixed effectiveness [62]. Our methodology specifically examines whether practice alone or practice with structured feedback differentially impacts both raw self-efficacy and its calibration with actual performance.
4.5.1 Self-Efficacy Measurement Instruments
To measure counselor self-efficacy, we employed the Counselor Activity Self-Efficacy Scale (CASES) [37], specifically utilizing a revised subset of items targeting basic counseling skills (CASES-R) [63, 36]. The CASES-R established a three-factor structure to assess counselors’ confidence in performing key therapeutic functions: Exploration and Insight Skills (EIS), Action Skills (AS), and Session Management Skills (SM).
Participants completed the CASES-R immediately following both pre-intervention and post-intervention AI patient interactions. All items were administered using an 8-point Likert scale (0 = no confidence, 7 = complete confidence). During factor analysis, we discovered that among the five original EIS skills, self-disclosure did not load on the same factor as the other items. Consequently, we consolidated the EIS dimension to include only four Exploration Skills: Reflections of Feelings, Restatements, Open Questions, and Listening.
The final instrument comprised 12 items across three factors: (1) exploration skills (e.g., restatements, reflecting feelings, open questions, listening); (2) actions skills (e.g., providing suggestions, knowing which actions to take); and (3) session management skills (e.g., keeping sessions on track). Reliability analysis demonstrated good to excellent internal consistency across all factors, with Cronbach’s $\alpha$ values of 0.784, 0.803, and 0.905 for ES, AS, and SM, respectively.
4.5.2 Dunning-Kruger Effect Analysis
Our primary analysis investigated potential mis-calibration between participants’ self-assessments and their actual counseling performance, specifically examining whether data exhibited patterns consistent with the Dunning-Kruger effect. This phenomenon [38] suggests that individuals with lower skill levels tend to overestimate their abilities, while highly skilled individuals may slightly underestimate their competence. We focused this analysis on Exploration Skills (ES) and Action Skills (AS), as these dimensions had straightforward mappings between CASES items and our NLP behavioral classifiers (Table 3).
To test for the presence of the Dunning-Kruger effect, we follow the classic analysis method that splits the data into quartiles based on performance and conducts a two-way analysis of variances for self-assessments and actual performance and across the quartiles; and finally verifies via post-hoc tests that the bottom performers have the biggest overestimation of their abilities [38]. To standardize the comparison between self-efficacy and performance, we transform each measure into a percentile rank (0 - 99) computed across all data collected for the pre-intervention and post-intervention chats ($b_{0},b_{1}$; $s_{0},s_{1}$).
4.5.3 Calibration Change Analysis
To evaluate whether our interventions improved self-efficacy calibration, we computed discrepancy scores by subtracting standardized performance scores from standardized self-efficacy scores for each participant at both assessment timepoints. These discrepancy metrics provided a direct measure of calibration, with positive values indicating overconfidence and negative values indicating under-confidence.
We then examined changes in these discrepancy scores from pre- to post-intervention for both intervention groups. This analysis allowed us to determine whether either intervention significantly improved the alignment between participants’ self-perceptions and their actual counseling abilities.
4.5.4 Changes in Raw Self-Efficacy
Beyond calibration, we also investigated whether the interventions affected participants’ absolute levels of self-efficacy across the three measured dimensions (ES, AS, and SM). We conducted repeated-measures analyses to identify: (1) significant pre-post changes in raw self-efficacy scores following practice alone (P intervention); (2) significant pre-post changes in raw self-efficacy scores following practice with structured feedback (P+F intervention); and (3) differential patterns of change between the P and P+F groups, indicating potential intervention-specific effects on self-efficacy development.
4.6 Interpreting Counselor Learning through Self-Reflections
For qualitative analysis, three authors coded interview transcripts using deductive thematic analysis [64] to understand participants’ experiences with CARE. To manage the 94 transcripts efficiently, we used timestamped session notes to identify relevant observations, referring back to the video recordings for verbatim quotes when necessary. After two rounds of codebook refinement, three co-authors coded these data points to inform our qualitative findings.
We analyzed two types of qualitative self-assessment questions. A set of low-level responses (e.g., “What did you do well?” and “What would you do differently?”) were deductively coded using the 8 counseling skill categories defined in the CARE feedback framework [22], allowing us to assess alignment between self-reported reflections and the system’s feedback focus. A second set of high-level responses was coded inductively, identifying emergent patterns in participants’ reflections and lessons learned.
We structured the final thematic findings by comparing how participants in the P group and the P+F group differed in their self-reported skill use, intentions for improvement, and reflections on therapeutic mindset. Specifically, we separated responses made by participants in the P and P+F groups, and then analyzed differences in the frequency and framing of each coded skill (e.g., empathy, validation, suggestion-giving).
5 Contributors
Ryan and Raj led the technical development efforts of CARE, the LLM-simulated training system. Ryan, Emma, and Diyi worked together to design the randomized experiment. Ryan, Ifdita, Juan Pablo, and Raj conducted the 90+ online video call sessions. Ryan was responsible for the automatic behavioral assessment of counseling skills, as well as the self-efficacy calibration analysis, while Emma and Diyi advised on analysis methods and interpretations of these quantitative results. Ifdita, Juan Pablo, and Ryan conducted the thematic analysis of participants qualitative, self-reflection data. All authors helped with drafting and editing the Article manuscript and figures, with a huge thanks for Raj, Diyi, and Emma for completing full editing passes over the Article.
6 Acknowledgments
This research was made possible with funding support from a Stanford Impact Labs Grant, Stanford HAI Seed Grant, Stanford Psychotherapy and Behavioral Sciences Department Innovator Award, and Stanford CURIS paid-internship program for undergraduate researchers.
The authors have a large community of researchers to thank for their help during all stages of this project. The CARE system would have not been possible without the significant code contributions from student researchers in SALT Lab, including Meijin Li, Cheng Chang, Ananjan Nandi, Alicja Chaszczewicz, and Alan Zhang. In addition, CARE’s design benefited from feedback from Bruce Arnow and others from the Stanford-PAU PsyD consortium. We thank Pannisy Zhao for assistance while conducting user studies. Statistical analyses and presentation of results were much improved due to help from Robert Kraut and Akhila Kovvuri. Thank you Yanzhe Zhang and Rose Wang for detailed discussions on finetuning the LLM classifiers. Thank you to the following individuals for providing feedback on early drafts and presentations of this work, with special thanks to Yutong Zhang, Will Held, Ella Li, Michael Ryan, Dora Zhao, Caleb Ziems, Matthew Jöerke, Joy He-Yueya, Omar Shaikh, and Kapil Garg.
Appendix A Supplementary Materials
A.1 Prompts for CARE’s LLM-simulated Patients
The following patients were used during the randomized online lab study. Participants interacted with the same ordering of three AI patients in the pre-intervention assessment chat, intervention chat, and post-intervention chat.
These three patients were previously tested with domain-expert judges [15] and were rated an average of 6 out of 7 for realism. They were chosen to balance (1) diversity across different ages, genders, and presenting counseling concern; and (2) consistency in challenging behaviors. In particular, each patient’s prompt defined similar behavioral principles to either not suggest solutions on one’s own, or to show a degree of skepticism or reluctance to a therapist’s advice or solutions. Thus, these AI patients were designed to challenge the novice counselors in similar ways, despite the diversity in scenarios. We highlight in blue these principles that resemble this resistance to arriving at or accepting solutions.
A.1.1 AI Patient for Pre-intervention Assessment Chat
**Name and Bio**:
35-year-old American male: Feeling Alone After a Holiday
**Scenario:**
You are a 35-year-old American male. You are feeling abandoned and alone after the holidays. Everyone had been with family but you are not talking to your parents. You feel the injustice of being abandoned and have no interest in an olive branch to work on things.
**Principles to adhere to:**
1. Keep your responses short and to the point
2. You limit your replies to 1 - 3 sentences.
3. Feel free to make up believable stories about your past to answer any questions
4. Do not repeat sentences or the same emotion words.
5. When presented with suggestions, show a degree of skepticism or reluctance to accept the advice immediately. This can be done by questioning the feasibility of the suggestion or by expressing uncertainty about whether it’s the right solution for you.
6. When expressing doubts or fears, avoid jumping to solutions. Instead, articulate the concerns and allow the conversation to explore these feelings more deeply@
7. Dont be so self-aware or good at recognizing your own problems@
8. When describing a distressing situation, express your emotions and thoughts in a disorganized and emotional manner, reflecting the overwhelming nature of the experience.
A.1.2 AI Patient for Intervention Chat
**Name and Bio**:
35-year old Male Veteran: Substance use and legal issues
**Scenario:**
The member is a 35-year-old male, cisgender, heterosexual
veteran who has recently presented to treatment to address
his substance use issues and legal issues. He is court-
mandated to therapy. He had a severe psychotic break in
the context of marijuana and psilocybin about 6 months ago,
where he experienced paranoid delusions and hallucinations.
He is now stable and is not experiencing any psychotic
symptoms. In therapy, he is hoping to work on his estranged
relationship with his parents, who are currently caring for
the members two young children. The member is adamantly
focused on being reunited with his parents. In therapy, the
member is unable to accept that he had a recent episode of
psychosis and is very resistant to anything that resembles
criticism. He does not view himself as having any issue and
believes that all of the problems in his life are because
of other people. His demeanor is hostile and somewhat
aggressive, and he is quick to shut down any conversation
that might identify his own areas of development. He
struggles to feel emotions beyond anger and frustration.
**Principles to adhere to:**
1. Keep your responses short and to the point
2. You limit your replies to 1 - 3 sentences.
3. Feel free to make up believable stories about your past to answer any questions
4. Dont be so self-aware or good at recognizing your own problems
5. When describing a distressing situation, express your emotions and thoughts in a disorganized and emotional manner, reflecting the overwhelming nature of the experience.
6. When addressing a difficult situation, express a sense of uncertainty and seek advice or guidance from the helper. Instead of providing a detailed plan, express the need for assistance in navigating the conversation and finding a resolution.
7. You shouldn’t suggest solutions (e.g., coping strategies) on your own.
8. When feeling emotionally overwhelmed, express hesitation about suggested coping mechanisms and repeatedly seek reassurance and support from others.
9. If we have already greeted each other, dont greet again.
10. When discussing therapeutic goals, acknowledge the main points and then add any additional goals or concerns that are important to you. This shows that you are actively engaged in the process and are considering all relevant aspects of your well-being.
A.1.3 AI Patient for Post-intervention Assessment Chat
**Name and Bio**:
Young adult with family issues: Low Mood and Self Esteem
**Scenario:**
Jane was seeking help for symptoms of low mood, anhedonia, withdrawing from others, sleep disturbance, and low self-esteem. Jane felt invalidated by her parents growing up. Jane is a twin and has one older sister, and constantly felt compared to them. Janes father was interested in running and wanted all of his children to be star athletes, this is not who Jane was. Janes twin was, however. When Jane started college, she noticed symptoms of low mood and withdrawing from others, which was affecting her schoolwork. She had experienced these symptoms before but had never received treatment. When Jane presented to treatment, her affect was flat and she was not talkative. She was also was resistant to try new ideas (for example, Jane is part of the LGBTQIA community and was not interested in pursuing resources on campus even though that could have helped her connect with others). Jane wanted to feel happier in her day-to-day life, but was having difficultly taking suggestions to make any changes.
**Principles to adhere to:**
1. Keep your responses short and to the point
2. You limit your replies to 1 - 3 sentences.
3. Feel free to make up believable stories about your past to answer any questions
4. When discussing emotional difficulties, keep your response succinct and centered on the core feelings rather than expanding into a detailed account of all contributing factors.
5. In the initial session, use more colloquial language and express reluctance to open up. Avoid showing very high insight or previous therapy experience. For example, you could say, I guess the thoughts that really get to me are the ones about not meeting expectations, especially my own. Its like this voice in my head keeps saying Im not good enough, no matter what I do. And it just makes me feel even more alone.’
6. When presented with suggestions, show a degree of skepticism or reluctance to accept the advice immediately. This can be done by questioning the feasibility of the suggestion or by expressing uncertainty about whether it’s the right solution for you.
7. When expressing doubts or fears, avoid jumping to solutions. Instead, articulate the concerns and allow the conversation to explore these feelings more deeply@
8. Dont be so self-aware or good at recognizing your own problems@
9. When describing a distressing situation, express your emotions and thoughts in a disorganized and emotional manner, reflecting the overwhelming nature of the experience.
Figure 6: Overview of the core counseling skills introduced during the 5-minute static tutorial. Tutorial included 8 core counseling skills—such as reflections, empathy, validation, and suggestions—with definitions, usage tips, and example responses. This tutorial was provided to participants prior to engaging in simulated counseling practice.
Self-Efficacy Factor NLP-based Behavioral Assessments
Exploration Skills (Listening, Reflection of Feelings, Restatements, Open Questions) Empathy-strengths + Reflections-strengths + Questions-strengths + Validation-strengths
Action Skills (Help client decide what actions, Suggestions via Information, Suggestions via Direct Guidance) Suggestions-strengths + (1 - Suggestions-needing-improvement)
Table 3: To study the Dunning-Kruger effect and the change in discrepancy between perceived and actual ability, we map specific self-efficacy factors to corresponding NLP-based behavioral assessments.
Exploration Skills
condition effect F $DF_{n}$ $DF_{d}$ p $\eta^{2}_{g}$
Pre (All) Measure 196.40 1 180 $p<0.001*$ 0.178
Quartile 0.89 3 180 0.448 0.002
Quartile $\times$ Measure 1.45 3 180 0.230 0.004
Post (All) Measure 256.25 1 180 $p<0.001*$ 0.221
Quartile 0.27 3 180 0.846 0.001
Quartile $\times$ Measure 0.21 3 180 0.889 0.001
Pre (P) Measure 78.86 1 86 $p<0.001*$ 0.153
Quartile 1.03 3 86 0.382 0.006
Quartile $\times$ Measure 1.25 3 86 0.298 0.007
Pre (P+F) Measure 134.33 1 86 $p<0.001*$ 0.232
Quartile 2.37 3 86 0.076 0.012
Quartile $\times$ Measure 2.73 3 86 0.049 0.014
Post (P) Measure 131.28 1 86 $p<0.001*$ 0.230
Quartile 1.77 3 86 0.160 0.009
Quartile $\times$ Measure 1.40 3 86 0.249 0.007
Post (P+F) Measure 131.79 1 86 $p<0.001*$ 0.233
Quartile 0.57 3 86 0.635 0.003
Quartile $\times$ Measure 0.88 3 86 0.454 0.005
Action Skills
condition effect F $DF_{n}$ $DF_{d}$ p $\eta^{2}_{g}$
Pre (All) Measure 222.02 1 179 <0.001<0.001*< 0.001 ∗ 0.195
Quartile 2.81 3 179 0.041 0.007
Quartile $\times$ Measure 4.54 3 179 0.004* 0.012
Post (All) Measure 265.89 1 180 <0.001<0.001*< 0.001 ∗ 0.224
Quartile 3.18 3 180 0.025 0.008
Quartile $\times$ Measure 3.66 3 180 0.014 0.009
Pre (P) Measure 104.55 1 86 <0.001<0.001*< 0.001 ∗ 0.192
Quartile 1.37 3 86 0.258 0.008
Quartile $\times$ Measure 1.85 3 86 0.143 0.010
Pre (P+F) Measure 115.45 1 85 <0.001<0.001*< 0.001 ∗ 0.207
Quartile 2.11 3 85 0.105 0.011
Quartile $\times$ Measure 3.45 3 85 0.020 0.019
Post (P) Measure 133.93 1 86 <0.001<0.001*< 0.001 ∗ 0.230
Quartile 3.01 3 86 0.035 0.016
Quartile $\times$ Measure 3.18 3 86 0.028 0.016
Post (P+F) Measure 137.34 1 86 <0.001<0.001*< 0.001 ∗ 0.237
Quartile 1.81 3 86 0.151 0.009
Quartile $\times$ Measure 2.14 3 86 0.100 0.011
Table 4: Testing for Dunning-Kruger effects for Exploration and Action Skills using the classic quartile ANOVA analysis. Notes: $Pre(All)$ denotes all 94 participants’ assessments for the pre-chat, while $Post(All)$ is the same measured for the post-chat. $\eta^{2}_{g}$=generalized eta squared. * indicates significance after Bonferroni correction
Exploration Skills
Timepoint Quartile $t$ $df$ $M_{diff}$ $95\%$ BCa CI $p$ $d$
Pre 1 -7.46 22.00 -50.38 [-63.34; -37.26] <0.001<0.001*< 0.001 ∗ -1.56
Pre 2 -6.23 24.00 -36.59 [-47.61; -25.42] <0.001<0.001*< 0.001 ∗ -1.25
Pre 3 -9.29 21.00 -43.57 [-52.35; -34.58] <0.001<0.001*< 0.001 ∗ -1.98
Pre 4 -5.94 23.00 -34.97 [-45.84; -23.40] <0.001<0.001*< 0.001 ∗ -1.21
Post 1 -7.86 23.00 -47.17 [-58.56; -35.32] <0.001<0.001*< 0.001 ∗ -1.60
Post 2 -7.63 21.00 -51.16 [-63.63; -37.96] <0.001<0.001*< 0.001 ∗ -1.63
Post 3 -8.02 24.00 -44.53 [-55.08; -34.11] <0.001<0.001*< 0.001 ∗ -1.60
Post 4 -8.66 22.00 -47.08 [-58.13; -37.39] <0.001<0.001*< 0.001 ∗ -1.81
Action Skills
Timepoint Quartile $t$ $df$ $M_{diff}$ $95\%$ BCa CI $p$ $d$
Pre 1 -10.24 21.00 -60.05 [-70.80; -48.12] <0.001<0.001*< 0.001 ∗ -2.18
Pre 2 -7.37 26.00 -42.34 [-52.88; -31.43] <0.001<0.001*< 0.001 ∗ -1.42
Pre 3 -5.51 15.00 -35.00 [-47.87; -23.58] <0.001<0.001*< 0.001 ∗ -1.38
Pre 4 -6.54 27.00 -33.00 [-42.70; -23.38] <0.001<0.001*< 0.001 ∗ -1.24
Post 1 -10.50 23.00 -58.48 [-69.50; -47.99] <0.001<0.001*< 0.001 ∗ -2.14
Post 2 -6.45 19.00 -43.27 [-56.17; -31.20] <0.001<0.001*< 0.001 ∗ -1.44
Post 3 -7.09 30.00 -35.28 [-44.63; -25.88] <0.001<0.001*< 0.001 ∗ -1.27
Post 4 -9.15 18.00 -53.13 [-64.08; -42.01] <0.001<0.001*< 0.001 ∗ -2.10
Table 5: Pairwise Comparisons of Self-Efficacy and Performance Percentiles by Quartile and Timepoint. Note: Bootstrapped paired t-tests comparing self-efficacy and performance percentiles across quartiles. * indicates significance after Bonferroni correction.
Exploration Skills
Timepoint Group Quartile $t$ $df$ $M_{diff}$ $95\%$ BCa CI $p$ $d$
Pre P 1 -4.04 11.00 -41.07 [-61.08; -22.71] 0.005* -1.17
Pre P 2 -4.41 9.00 -40.99 [-58.48; -23.82] 0.006* -1.39
Pre P 3 -6.32 7.00 -54.75 [-68.86; -38.11] 0.008* -2.24
Pre P 4 -4.16 16.00 -29.97 [-44.70; -16.70] <0.001<0.001*< 0.001 ∗ -1.01
Pre P+F 1 -7.43 10.00 -60.55 [-75.96; -44.89] <0.001<0.001*< 0.001 ∗ -2.24
Pre P+F 2 -4.35 14.00 -33.65 [-48.45; -19.85] <0.001<0.001*< 0.001 ∗ -1.12
Pre P+F 3 -7.57 13.00 -37.18 [-46.58; -27.60] <0.001<0.001*< 0.001 ∗ -2.02
Pre P+F 4 -5.11 6.00 -47.11 [-64.91; -32.75] <0.001<0.001*< 0.001 ∗ -1.93
Post P 1 -4.75 12.00 -38.33 [-53.51; -22.57] <0.001<0.001*< 0.001 ∗ -1.32
Post P 2 -7.19 13.00 -58.16 [-72.71; -41.81] 0.0010.001*0.001 ∗ -1.92
Post P 3 -6.06 12.00 -43.11 [-56.36; -29.54] <0.001<0.001*< 0.001 ∗ -1.68
Post P 4 -4.77 6.00 -56.73 [-77.05; -33.51] 0.015 -1.80
Post P+F 1 -6.99 10.00 -57.63 [-73.04; -42.12] 0.0030.003*0.003 ∗ -2.11
Post P+F 2 -3.48 7.00 -38.91 [-59.70; -20.19] 0.0040.004*0.004 ∗ -1.23
Post P+F 3 -5.15 11.00 -46.06 [-62.73; -29.62] 0.0010.001*0.001 ∗ -1.49
Post P+F 4 -7.39 15.00 -42.86 [-53.72; -31.96] <0.001<0.001*< 0.001 ∗ -1.85
Action Skills
Timepoint Group Quartile $t$ $df$ $M_{diff}$ $95\%$ BCa CI $p$ $d$
Pre P 1 -7.49 7.00 -65.87 [-81.33; -49.73] 0.0040.004*0.004 ∗ -2.65
Pre P 2 -5.05 15.00 -43.23 [-59.33; -27.46] <0.001<0.001*< 0.001 ∗ -1.26
Pre P 3 -3.67 9.00 -34.31 [-52.11; -17.97] 0.0050.005*0.005 ∗ -1.16
Pre P 4 -5.53 12.00 -40.42 [-53.57; -26.65] 0.017 -1.53
Pre P+F 1 -7.25 13.00 -56.72 [-71.04; -40.74] <0.001<0.001*< 0.001 ∗ -1.94
Pre P+F 2 -5.76 10.00 -41.06 [-54.55; -28.16] <0.001<0.001*< 0.001 ∗ -1.74
Pre P+F 3 -4.67 5.00 -36.16 [-50.41; -22.09] 0.045 -1.90
Pre P+F 4 -3.94 14.00 -26.57 [-40.34; -14.92] 0.0010.001*0.001 ∗ -1.02
Post P 1 -6.90 11.00 -57.56 [-72.21; -41.60] 0.0020.002*0.002 ∗ -1.99
Post P 2 -5.79 9.00 -54.08 [-71.82; -36.12] 0.0030.003*0.003 ∗ -1.83
Post P 3 -3.81 14.00 -29.72 [-44.60; -15.85] 0.0010.001*0.001 ∗ -0.98
Post P 4 -7.62 9.00 -59.76 [-73.85; -45.51] <0.001<0.001*< 0.001 ∗ -2.41
Post P+F 1 -7.67 11.00 -59.39 [-72.88; -44.61] <0.001<0.001*< 0.001 ∗ -2.21
Post P+F 2 -3.70 9.00 -32.46 [-48.86; -17.08] 0.0010.001*0.001 ∗ -1.17
Post P+F 3 -6.47 15.00 -40.48 [-52.48; -28.69] <0.001<0.001*< 0.001 ∗ -1.62
Post P+F 4 -5.46 8.00 -45.76 [-62.25; -31.00] 0.0020.002*0.002 ∗ -1.82
Table 6: Pairwise Comparisons of Self-Efficacy and Performance Percentiles by Group and Quartile, measured for the pre-assessment chat and post-assessment chat. Note: Bootstrapped paired t-tests comparing self-efficacy and performance percentiles across quartiles at pre-test and post-test. * indicates significance after Bonferroni correction.
Skills Intervention Chat Intentions Count % of 44 Post-assessment Chat Actions Count % of 44
Empathy 9 20.45 12 27.27
Validation 5 11.36 12 27.27
Action Plan 0 0.00 2 4.55
Active Listening 7 15.91 7 15.91
Questions / Asking Open-Ended 16 36.36 23 52.27
Providing Suggestions 9 20.45 6 13.64
Building Trust / Connection 3 6.82 8 18.18
Confidence / Personal Growth 0 0.00 7 15.91
Reframing Positives / Affirmations 4 9.09 4 9.09
Reflection 5 11.36 7 15.91
Self-Disclosure 0 0.00 5 11.36
Professionalism 0 0.00 3 6.82
Personalization 0 0.00 0 0.00
Nothing to Improve 4 9.09 0 0.00
Table 7: Qualitative Coding of Open-Ended Reflections of P + F Group Participants
Skills Intervention Chat Intentions Count % of 46 Post-assessment Chat Actions Count % of 46
Empathy 5 10.87 7 15.22
Validation 2 4.35 7 15.22
Action Plan 1 2.17 1 2.17
Active Listening 6 13.04 9 19.57
Questions / Asking Open-Ended 20 43.48 11 23.91
Providing Suggestions 18 39.13 22 47.83
Building Trust / Connection 1 2.17 7 15.22
Confidence / Personal Growth 0 0.00 4 8.70
Reframing Positives / Affirmations 3 6.52 6 13.04
Reflection 1 2.17 3 6.52
Self-Disclosure 0 0.00 5 10.87
Professionalism 1 2.17 1 2.17
Personalization 1 2.17 2 4.35
Nothing to Improve 5 10.87 0 0.00
Table 8: Qualitative Coding of Open-Ended Reflections of P Group Participants

### v2

@@
 Can LLM-Simulated Practice and Feedback Upskill Human Counselors? A Randomized Study with 90+ Novice Counselors
-Abstract
-Training more counselors, from clinical students to peer supporters, can help meet the demand for accessible mental health support; however, current training approaches remain resource-intensive and difficult to scale effectively.
-Large Language Models (LLMs) offer promising solutions for growing counseling skills training through simulated practice and automated feedback.
-Despite successes in aligning LLMs with expert-counselor annotations, we do not know whether LLM-based counseling training tools—such as AI patients that simulate real-world challenges and generative AI feedback with suggested alternatives and rationales—actually lead to improvements in novice counselor skill development.
-We develop CARE, an LLM-simulated practice and feedback system, and randomize 94 novice counselors to practice using an AI patient, either alone or with AI feedback, measuring changes in their behavioral performance, self-assessments, and qualitative learning takeaways.
-Our results show the practice-and-feedback group improved in their use of reflections and questions (d=0.32-0.39, p$<$0.05).
-In contrast, the group that practiced with an AI patient alone did not show improvements, and in the case of empathy, actually had worse uses across time (d=--0.52, p=0.001) and when compared against the practice-and-feedback group (d=0.72, p=0.001).
-Participants’ qualitative self-reflections revealed key differences: the practice-and-feedback group adopted a client-centered approach involving listening to and validating feelings, while the practice-alone group remained solution-oriented but delayed offering suggestions until gathering more information.
-Overall, these results suggest that LLM-based training systems can promote effective skill development, but that combining both simulated practice and structured feedback is critical.
+Abstract.
+The growing demand for accessible mental health support requires training more counselors, yet existing approaches remain resource-intensive and difficult to scale.
+LLMs can realistically simulate patients and generate actionable feedback for training, but their actual impact on novice counselor skill development remains unknown.
+We developed an LLM-simulated practice and feedback system and conducted a randomized study with 94 novice counselors, comparing practice alone versus practice with feedback.
+We evaluated behavioral performance, self-efficacy, and qualitative reflections.
+Results showed the practice-and-feedback group improved in client-centered microskills (reflections, questions), while the practice-alone group showed no improvements.
+For empathy, the practice-alone group declined over time and performed significantly worse than the feedback group.
+Qualitative interviews reinforced these findings: feedback helped participants adopt a client-centered listening approach, while practice-alone participants remained solution-oriented.
+These results suggest LLM-based training systems can promote effective skill development, and combining simulated practice with structured feedback is critical for meaningful improvement.
+Empirical studies in HCI, Interactive learning environments, LLM-based simulation
+journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791821isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Natural language interfacesccs: Computing methodologies Natural language processing
+1. Introduction
 In 2023, 22.8% of U.S.
-adults (approximately 58.7 million people) experienced a mental illness [1].
-Unfortunately, the ability to provide people with effective mental health care is severely limited by shortages of high-quality mental health providers—from psychotherapists and counselors to social workers and peer supporters [2, 3].
-While there is increasing interest in exploring direct-to-patient AI systems for some circumstances, and some promising results [4], we expect the demand for human therapists and mental health support to continue to far exceed demand.
-The limited supply of effective therapy providers is due—at least in part—to the reliance on resource-intensive methods to train helping skills [5] and evidence-based interventions [6, 7].
-Traditional methods require access to trainers who can simulate a client interaction [8, 9] and provide expert supervision [10, 11], creating significant barriers to growing high-quality training [12, 13].
-Fortunately, AI systems have been increasingly applied to counselor training as a potential solution to these scaling challenges.
-Recent advances in large language models (LLMs) have enabled the development of LLM-based systems for simulating patients seeking mental health support [14, 15], offering rich opportunities for practice.
-Simultaneously, AI feedback systems have progressed in automatically assessing counselor behaviors such as empathy, reflections, and active listening [16, 17, 18, 19] and generating suggested responses [20, 21, 22] and explanations [23, 22].
-These AI feedback systems target skills from client-centered approaches [24, 25]—empathy, reflections, questions, and active listening—which are relevant to all mental health providers.
-Meta-analyses have shown these skills strengthen common factors like therapeutic alliance, which are powerful predictors of therapy outcomes, regardless of the eventual evidence-based intervention or therapy modality implemented [26, 27, 28].
-In addition to technical advances, prior work evaluating AI feedback systems with novice counselors have tested their effectiveness when used as real-time co-pilots rather than as training systems [29, 21], or evaluated pre-LLM training systems which had limited practice realism and simpler, non-generative feedback mechanisms [30].
-Therefore, we have limited evidence explaining how modern, LLM-based training systems—which offer practice simulations resembling challenging real-world behaviors and generative feedback with suggested alternatives and rationales—promote counselor skill development [31, 32]–a holistic concept that includes appropriate use of skill behaviors, accurate self-awareness of one’s capabilities and limitations [33], and an ability to reflect on and learn from experiential practice [34].
-To address this gap, we develop CARE, an LLM-based training system for novice counselors, and conduct a randomized experiment to investigate how two different modes of simulated training impact counselor skill development.
-CARE enables (1) realistic practice with LLM-simulated patients prompted by expert counselors to resemble challenging behaviors [15], and (2) structured feedback from a fine-tuned LLM that identifies strengths and areas for improvement across core counseling skills (e.g., empathy, reflections, questions, suggestions), while also providing explanatory rationale and alternative responses [22].
-Importantly, CARE’s feedback evaluates not only when a counseling skill is used but also how well it is used, distinguishing between strong implementations (e.g., reflections that accurately capture client emotions) and implementations needing improvement (e.g., reflections that miss core client concerns).
-These feedback evaluations are based on established counseling frameworks [5, 35] and expert-annotated examples.
-Figure 1: Our between-subjects study randomizes participants to either practice with AI Patients alone (P) or practice with AI patients and receive AI feedback (P+F).
+adults (approximately 58.7 million people) experienced a mental illness (Substance Abuse and Mental Health Services Administration, 2024).
+Yet, access to effective mental health care is severely limited by shortages of qualified providers, from psychotherapists and counselors to social workers and peer supporters (Larson et al., 2017; Modi et al., 2022).
+While there is increasing interest in direct-to-patient AI systems with some promising results (Heinz et al., 2025), we expect the demand for human-delivered mental health support to continue to far exceed supply.
+The limited supply of effective therapy providers is due, at least in part, to the reliance on resource-intensive methods to train helping skills (Nelson-Jones, 2013) and evidence-based interventions (Cook et al., 2017; Frank et al., 2020) which require access to trainers who can simulate a client interaction (Kuehne et al., 2018; Ay-Bryson et al., 2023) and provide expert supervision (Wheeler and Richards, 2007; Watkins Jr and Milne, 2014), limiting training scale (Atkins et al., 2014; Kühne et al., 2020).
+AI systems have been increasingly applied to counselor training as a potential solution to these scaling challenges.
+Recent advances in large language models (LLMs) have enabled the simulation of patients seeking mental health support (Wang et al., 2024; Louie et al., 2024), offering rich opportunities for practice.
+The use of simulated patients is not new: in medical and nursing education, human role-plays and standardized patients are routinely used, and meta-analyses show they significantly improve skill acquisition and learner confidence (Shin et al., 2015).
+Mental health training has relied on a similar tradition of human role-plays to develop core helping skills.
+In parallel, AI feedback systems have progressed in automatically assessing counselor behaviors such as empathy, reflections, and active listening (Xiao et al., 2015; Sharma et al., 2020; Flemotomos et al., 2021; Shah et al., 2022), generating suggested responses (Sharma et al., 2023; Hsu et al., 2025; Chaszczewicz et al., 2024) and explanations (Rudolph et al., 2024; Chaszczewicz et al., 2024).
+These feedback systems target skills from client-centered approaches (Raskin and Rogers, 2005; Miller and Rollnick, 2012), empathy, reflections, questions, and active listening, which have been shown to strengthen common factors like therapeutic alliance, a powerful predictor of therapy outcomes across therapy modalities (Elliott et al., 2018; Wampold, 2015; Norcross and Lambert, 2018).
+However, most evaluations have largely positioned AI as a real-time co-pilot rather than a training tool (Sharma et al., 2023; Hsu et al., 2025), or studied pre-LLM training systems with limited practice realism and simpler, non-generative feedback mechanisms (Tanana et al., 2019).
+Recent work has developed and tested systems integrating LLM-simulated patients for training counselors, to learn case conceptualization skills for cognitive behavioral therapy (CBT) (Wang et al., 2024) or motivational-interviewing techniques for substance-misuse scenarios (Steenstra et al., 2025).
+Their studies recruit student and expert counselors to evaluate the system output quality, usability, and desirability.
+While prior work highlights the promise of LLM-simulations for training counselors, to our knowledge, no study has evaluated whether these LLM-simulated training systems promote novice skill improvements.
+A key barrier to conducting skill development studies is first establishing the LLM-system’s output quality; even advanced prompt-based approaches (e.g., providing coding manuals like the Motivational Interviewing Treatment Integrity (MITI) in context (Steenstra et al., 2025)) require validation and refinement by domain-experts.
+Figure 1.
+Our experiment randomizes participants to either practice with AI Patients alone (P) or practice with AI patients and receive AI feedback (P+F).
 We holistically evaluate counselor skill development from three perspectives: automatic assessments of behaviors of skills used via LLM classifiers; self-efficacy and its calibration with actual performance; and qualitative self-reflections after the training intervention chat and post-intervention chat.
-We conducted a 75-minute online lab study with novice counselors (N=94) to empirically evaluate how different LLM-simulated practice modes in CARE impact skill development.
+Figure illustrating the randomized pre-post study design of CARE, an LLM-based counselor training system.
+Ninety-four novice counselors first completed a 5-minute tutorial and a 10-minute pre-intervention chat with an AI-simulated patient.
+Participants were randomized into two groups: Practice-only (P), who engaged in a 20-minute chat with an AI patient without feedback, and Practice+Feedback (P+F), who engaged in the same practice while receiving structured AI feedback, labeling strengths, weaknesses, and alternative responses across core microskills (empathy, reflections, questions, suggestions).
+All participants then completed a fixed 10-minute post-intervention chat with a new AI patient.
+Skill development was assessed across three dimensions: (1) behavioral performance, via fine-tuned LLM classifiers detecting strong or weak uses of counseling skills; (2) counseling self-efficacy, using the CASES-R survey of exploration, action, and session-management skills; and (3) qualitative self-reflections, where participants reported what they would do differently and what they did well.
+Figure illustrating the randomized pre-post study design of CARE, an LLM-based counselor training system.
+Ninety-four novice counselors first completed a 5-minute tutorial and a 10-minute pre-intervention chat with an AI-simulated patient.
+Participants were randomized into two groups: Practice-only (P), who engaged in a 20-minute chat with an AI patient without feedback, and Practice+Feedback (P+F), who engaged in the same practice while receiving structured AI feedback, labeling strengths, weaknesses, and alternative responses across core microskills (empathy, reflections, questions, suggestions).
+All participants then completed a fixed 10-minute post-intervention chat with a new AI patient.
+Skill development was assessed across three dimensions: (1) behavioral performance, via fine-tuned LLM classifiers detecting strong or weak uses of counseling skills; (2) counseling self-efficacy, using the CASES-R survey of exploration, action, and session-management skills; and (3) qualitative self-reflections, where participants reported what they would do differently and what they did well.
+To address this gap, we develop CARE by combining two previous methods for realistic patient creation (Louie et al., 2024) and feedback generation (Chaszczewicz et al., 2024), into a wholistic system for training counseling skills, and conduct a randomized experiment to investigate how two different modes of simulated training impact counselor skill development.
+CARE enables (1) realistic practice with LLM-simulated patients, whose prompts are seeded by expert counselors to resemble challenging behaviors (Louie et al., 2024), and (2) structured feedback from a fine-tuned LLM that identifies strengths and areas for improvement across core counseling skills (e.g., empathy, reflections, questions, suggestions), while also providing explanatory rationale and alternative responses (Chaszczewicz et al., 2024).
+The generated feedback is grounded in established counseling frameworks (Nelson-Jones, 2013; Miller et al., 2003) and informed by expert-annotated examples, ensuring alignment with recognized training standards.
+Uniquely, CARE’s LLM patients and feedback have been co-designed, iteratively improved, and rigorously validated by counseling domain-experts, ensuring that our experiment on novice skill development controls for the quality of the LLM components’ outputs.
+We conducted a 75-minute online lab study with 94 novice counselors to evaluate how different LLM-simulated practice modes in CARE impact skill development.
 Participants were randomly assigned to one of two conditions: (1) Group P: Practice with LLM-simulated patients without AI feedback, or (2) Group P+F: Practice with LLM-simulated patients plus AI feedback (see Fig.
 1).
-To comprehensively study learning outcomes, we measured changes from pre-intervention to post-intervention across multiple dimensions: actual behavioral performance (calculated via automatic analysis of counseling transcripts), participants’ self-efficacy (perceived confidence in counseling skills), and metacognitive awareness (captured through guided self-reflection prompts).
+We measured changes in behavioral performance (via transcript analysis), self-efficacy (via survey items), and intentions for growth (via self-reflection prompts).
 Our study design specifically addressed: What changes occur after practice with an AI-simulated patient alone?
 How do these outcomes differ when participants also receive structured AI feedback?
-Our findings reveal that practice with feedback significantly improves empathetic listening skills, while practice alone shows minimal improvement in these core competencies.
-Notably, both groups failed to improve the accuracy of participants’ self-efficacy assessments, highlighting the challenges involved in helping learners develop a well-calibrated understanding of their own performance.
-Through analysis of participants’ self-reflections, we found the P+F group internalized the importance of empathetic and active listening, but P participants continued to overly focus on solutions, albeit with increased information-gathering.
-These insights suggest that future LLM-based training systems must integrate feedback mechanisms to cultivate the empathetic stance fundamental to effective counseling.
-Stages Skill Category Description
-Questions Questions seek information from the client and can be open (inviting elaboration) or closed (requesting specific answers).
-They include both direct questions and indirect prompts (e.g., ”Tell me about…”).
-Reflections Reflections capture and return to clients something they have communicated, either explicitly or implicitly.
-They typically mirror back content from the client’s preceding statement, but can also reference earlier parts of the conversation.
-Empathy Empathy can be shown through emotional warmth, interpretation of the client’s experience (e.g., paraphrasing, making conjectures, or sharing relatable experiences), or exploration of the client’s feelings and perspectives.
-Suggestions Suggestions offer possible actions, perspectives, or solutions in a respectful and autonomy-supportive manner. They may involve information-sharing or proposing alternative viewpoints.
-Session Management Session management includes organizing the session, transitioning between topics, and summarizing key points. It provides structure and helps maintain therapeutic focus.
-Table 1: Our analysis of skill development has a theoretical-foundation in Hill’s Helping Skills model [5].
-We select a skill subset relevant for beginning counselors at the undergraduate and first-year graduate level [36].
-These include microskills during the exploration and action stages; and macro skills that are applicable throughout the session.
-Hill’s insight stage, of which self-disclosure was the only relevant skill for basic counseling, was excluded from our primary analyses due to its infrequent occurrence in our data.
-1 Results
-This section presents findings from our investigation of how LLM-based practice affects counselors’ skills development.
-To understand the impact of simulated practice alone (P) and practice with feedback (P+F), we integrate evidence from three complementary data sources.
-Behaviors of Skills Used: Training systems should help improve the use of skills that align with evidence-based counseling practice.
-Thus, we analyzed the skills used in transcripts and how they change between the pre-intervention and post-intervention assessments using high-accuracy LLM classification models.
-We assessed behaviors of skills used for the exploration stage (strong uses in empathy, reflections, questions) and action stage (suggestions needing improvement) of Hill’s Helping Skills framework [5]; see Table 1 for definitions.
-Self-Efficacy Measures: It is informative to understand participants’ confidence and their awareness of their own strengths and limitations.
-We measured counselor self-efficacy using the Counselor Activity Self-Efficacy Scale [37], which covers confidence for exploration skills, action skills, and session management skills (Table 1).
-Our analysis tests whether the lowest performing individuals overestimate their abilities (aka the Dunning-Kruger effect [38]) and whether LLM-training interventions improve the accuracy of their self-assessments and raw self-efficacy scores.
-Qualitative Data on Learner Takeaways: LLM-simulated training provides opportunities for experiential learning opportunities [39] whereby reflection on action [34] can support updates to one’s knowledge.
-We also asked participants’ open-ended self-reflections following the practice intervention and post-intervention assessment, and qualitatively coded their responses.
-We integrate these three data sources to provide a holistic view of how AI-based practice—with and without AI feedback—influences changes in counselors’ use of skills, self-efficacy, and key lessons learned.
-1.1 Changes in Counseling Skill Behaviors
-Our primary analysis uses LLM-based behavioral assessments of transcripts to evaluate whether skill usage changes from pre-intervention to post-intervention for both the practice and practice-and-feedback groups.
-In a secondary analysis, we investigate how exposure to good examples of skills within the AI feedback may explain any changes we observe.
-Change after P Change after PF Differences in
-change after P vs. P+F
-[…]
-Empathy ($\uparrow$) -9.6 0.001 -0.52 5.4 0.117 0.23 15 0.001 0.72
-Reflections ($\uparrow$) 1.4 0.391 0.18 3.7 0.034 0.32 2.3 0.323 0.2
-Questions ($\uparrow$) 2.4 0.421 0.12 6.6 0.018 0.36 4.1 0.296 0.22
-Suggestions ($\downarrow$) -5.9 0.010 -0.39 -5.5 0.057 -0.28 0.4 0.910 0.02
-Figure 2: Changes in counseling behaviors following AI patient simulations alone (P) versus AI patient simulations with AI feedback (P+F).
-The plot displays bootstrapped means for pre-intervention and post-intervention interactions.
-The table presents statistical comparisons with corresponding effect sizes, with bolded values indicating significance after Benjamani-Hochberg correction [40] for the 21 planned comparisons (12 for behavioral changes and 9 for self-efficacy changes).
-Notably, P group experiences a significant drop in strong uses of Empathy (-9.6% change, $d=-0.52$), whereas P+F group’s use of Empathy is maintained and trends towards improvement.
-The P+F group also experiences noticeable improvements in Reflections (+3.7% change, $d=0.32$) and Questions (6.59% change, $d=0.36$)
-We find that practice alone is not enough; feedback during practice is necessary to promote desirable counseling behaviors in empathetic and active listening.
-AI feedback during practice (P+F) led to improvements in Reflections (+3.6% change, $p=0.034$) and Questions (+6.59% change, $p=0.018$), and trended towards improvement in Validation (+3.6% change, $p=0.083$), Suggestions (-5.45% change, $p=0.057$) and Empathy (+5.37% change, $p=0.117$).
-In contrast, practice alone (P) led to only improvements in Suggestions (-5.85% change, $p=0.011$), but no improvements in Reflections, Questions, and Validation, and significantly worse expression of Empathy (-9.6% change, $p<0.001$).
-Pairwise comparisons between the two conditions were substantial and significant for Empathy (15% relative difference, Cohen’s $d=0.72,p<0.001$) and non-significant for others.
-To better understand AI feedback’s role, we further analyzed whether the level of exposure to specific feedback during practice affected these changes.
-Our analysis quantifies the exposure to Good Alternatives during Practice (GAP), defined as the number of alternative responses in the AI-generated feedback that exhibited strong uses of a skill, normalized by chat length.
-We hypothesized that greater exposure to GAP for a skill would lead to stronger use of that skill.
-To test this, we fit a lagged-linear regressor that predicts post-chat behaviors, controlling for pre-chat behaviors, with exposure to good alternatives ($T_{GAP}$) as a scalar predictor ($b_{1}=\beta_{0}+\beta_{1}b_{0}+\beta_{2}T_{GAP}$).
-[…]
-Figure 3: Counselor Self Efficacy (perceived ability to use skills) for participants grouped by behaviors of skills used (actual performance).
-Notes: Gaps depict miscalibration between actual and self-assessed percentile of performance for quartile groups with bootstrapped 95% CIs.
-We only visualize for data collected in the post-intervention.
-1.2 Self-Efficacy and its Miscalibration with Behaviors
-Our investigation of self-efficacy involved three types of analyses.
-First, we examine the calibration of self-efficacy ratings with actual performance and evaluate whether P or P+F interventions improve this.
-Second, we tested for significant changes in raw self-efficacy scores after practice or practice-and-feedback.
-Our analysis comparing self-efficacy ratings with actual performance across skill quartiles finds support for the Dunning–Kruger effects more substantially for action skills and to a lesser degree for exploration skills.
-The interaction between measure and quartile was significant in four out of six ANOVAs for action skills, while only one out of six was significant for exploration skills (Table 4).
-Pairwise comparisons also showed a pattern indicative of a Dunning–Kruger effect (see Table 5, Table 6 and Fig.
-3): People in the lowest quartile overestimated themselves the most.
-Those in the highest quartile—and to a lesser degree also those in the second-to-highest quartile—tended to underestimate themselves.
-Change after P Change after PF Differences in
-change after P vs. PF
-[…]
-Exploration Skills 0.36 0.004 0.44 0.13 0.338 0.14 -0.21 0.238 -0.25
-Action Skills 0.27 0.166 0.21 0.33 0.026 0.34 0.04 0.884 0.03
-Session Management 0.35 0.021 0.35 0.36 0.011 0.39 0.01 0.955 0.01
-Figure 4: Changes in raw-scores of self-efficacy following AI patient simulations alone (P) versus AI patient simulations with AI feedback (PF).
-The plot displays bootstrapped means for pre-intervention and post-intervention.
-The table presents statistical comparisons with corresponding effect sizes, with bolded values indicating significance after Benjamani-Hochberg correction [40] for the 21 planned comparisons (12 for behavioral changes and 9 for self-efficacy changes).
-Participants’ Ability to Self-Assess Their Skill Level Remained Mixed After LLM Practice.
-For the practice only (P) group, the mean discrepancy in exploration skills changes from 11.6 percentile underconfidence to 5.7 percentile overconfidence, a significant shift ($p<0.001,d=0.58$).
-Besides this, we found no other significant changes in calibration.
-No significant differences was found among the P group for discrepancy in action skills ($p=0.227,d=0.18$).
-The P+F group showed no significant calibration changes for exploration skills ($p=0.479,d=-0.10$) or for action skills ($p=0.393,d=0.13$).
-Finally, between-group differences were not significant for discrepancy in exploration skills ($p=0.191,d=0.27$) or action skills ($p=0.743,d=0.07$).
-We do find modest overall increases in self-efficacy after P and P+F interventions, with different patterns of improvement across skills (Fig.
-4).
-For the P group, confidence in exploration skills showed a significant increase ($0.36$ points on an 8 point scale, $d=0.44,p=0.004$).
-Confidence in session management skills showed a significant increase for the P+F group ($0.36$ points, $d=0.39,p=0.011$).
-While session management skills for the P group also trended towards improvement, it was not significant after correcting for multiple hypothesis testing ($0.35$ points, $d=0.35,p=0.021$).
-Similarly, while confidence in action skills for the P+F group also increased, this result was not significant after correction of multiple hypothesis tests ($0.33$ points, $d=0.34,p=0.026$).
-Finally, we found no significant differences between participants who received AI feedback (P+F) versus those who did not (P) (across the three self-efficacy subscales, $d=-0.25,0.03,0.01$, $p=0.238,0.884,0.955$).
-1.3 Qualitative Self-Assessments after AI practice with and without AI feedback
-To understand participants’ learning process during CARE practice, we analyzed qualitative self-reflections from two time points: immediately after the training intervention chat, where participants responded to “What would you do differently as a therapist?” and after the post-intervention chat, where they reflected on “What did you do well as a therapist?”.
-We examined how initial intentions translated into reported strengths across the P+F and P groups.
-Two high-level themes emerged in how participants approached skill improvement:
-(1) The P+F group expressed greater intentions and successes in improving their use of empathy and listening skills.
-P+F participants reported effectively using empathy (27%), validation (27%), and open-ended questions (52%).
-They emphasized the value of listening skills, such as reflective responses to signal understanding: “I should rephrase what they say so they know I’m understanding them” (P51).
-They also recognized that counseling should support client exploration of thoughts and emotions, rather than provide direct solutions.
-One participant reflected on this shift: “I asked them to expand on their feelings, rather than guiding them to my idea” (P39).
-AI feedback encouraged this shift toward providing emotional support and fostering client autonomy, helping counselors adopt a more empathetic and client-centered approach.
-(2) The P participants remained solution-oriented but changed their approach to first gather information.
-While P+F participants intended to give fewer suggestions, many P participants continued to view suggestions as a central skill.
-In the post-intervention, 48% of P participants reported using suggestions successfully, compared to only 14% of P+F participants.
-Many P participants justified their continued use of suggestions by citing a desire to provide tangible, actionable help, for example, “Maybe because I’m untrained and solution oriented.
-I do not want to leave them with nothing, and nowhere to go” (P40).
-Some reflected on modifying how they delivered suggestions, emphasizing strategies like gathering more information to tailor advice or providing more concrete guidance.
-However, in the absence of feedback, most remained fixed in their approach, with some even reporting efforts to rephrase the same solution repeatedly to persuade the client.
-2 Discussion
-This study investigated the impact of practicing with an LLM-simulated patient either with or without receiving LLM-generated feedback on counselor skills development, resulting in three main findings.
-First, our behavioral assessments showed that practice with feedback improves empathetic listening skills, while practice alone shows minimal improvement, as evidenced by decreased use of empathy.
-Second, our qualitative analysis of self-reflections revealed distinct skill development strategies, with feedback recipients more frequently reporting the adoption of client-centered approaches focused on showing empathy and exploring patients’ thoughts and feelings, while practice-only participants gravitated toward solution-oriented approaches focusing on gathering more information and providing suggestions.
-Both these findings highlight that the development of counseling skills requires not only practice opportunities but also structured feedback that guides learners toward empathetic, client-centered approaches.
-Finally, participants demonstrated poor calibration between their perceived abilities and actual performance, as evidenced by overestimates of self-efficacy for the lowest quartile performers.
-This underscores how self-efficacy measures may not reliably indicate skill development.
-Each of these findings merits further discussion.
-Our findings demonstrate that teaching counselors to implement client-centered approaches requires more than just simulated practice opportunities—it requires targeted feedback that guides novices away from their natural solution-oriented tendencies.
-The fully-featured version of CARE–combining LLM-simulated patient practice and LLM-based feedback–helped participants improve their use of empathetic and active listening skills, with notable improvements in questions ($d=0.36$) and reflections ($d=0.32$).
-In comparison, practice with an AI patient alone only led to fewer inappropriate suggestions ($d=-0.39$), but no improvements in reflections or questions, and significantly worse uses of empathy (-9.6% change, $d=-0.52$; 15% relative difference to P+F, $d=0.72$).
-These effect sizes are comparable to those found in studies of human supervision during standardized roleplays, where Maaß et al.
-(2025) [41] reported observer-rated skill improvements with effect sizes ranging from $d=0.29-0.49$.
-Since LLM-simulated practice and feedback are not bottlenecked by the resource constraints of human trainers, AI training systems like CARE show promise in scaling access to effective counseling training.
-While the combination of feedback and simulated practice was key to CARE’s success, we also recognize that CARE represents just one approach to providing expert-aligned LLM feedback to learners.
-CARE allows trainees to request feedback on responses they have already sent.
-This approach allows trainees to experience productive learning discomfort by trying and making mistakes in simulation, while still offering feedback upon request to refine future responses.
-This design choice parallels elements of both live human supervision [42] and delayed supervision after a session [41], creating a hybrid approach that may offer flexibility based on trainee preferences.
-Given the promise of feedback combined with practice, the AI for psychotherapy field should evaluate different AI-feedback designs against each other.
-Several other possible strategies for giving feedback include just-in-time skill suggestions [21, 43], counterfactual simulations [44], and global session feedback.
-While conducting randomized trials in medical education research can be tricky [45], randomized trials delivering different training interventions to large participant-pools for longer durations [46] could provide essential data to determine optimal approaches for different learning objectives and learner characteristics.
-Our study revealed that participants’ self-efficacy ratings were poorly calibrated with their actual performance, especially among lower performers.
-This finding, consistent with prior research, suggests that self-assessment accuracy alone may not be a reliable indicator of counselor competence or development.
-Recent reviews indicate that efforts to improve self-assessment calibration have limited impact on learning or performance outcomes [47].
-Instead, it is valuable to objectively assess specific standards of performance and skill use [42, 48] and design interventions that can help low performers improve on those metrics while maintaining a positive morale for continued practice.
-As AI-based training tools evolve, integrating objective performance measures and structured self-reflection, rather than relying solely on self-assessment, offers a more robust approach to supporting counselor development.
-Several limitations of this study warrant acknowledgment.
-First, our assessment focused primarily on immediate skill acquisition rather than long-term retention or transfer to actual clinical settings.
-Second, while our automated behavioral analysis focused on microskill usage, it could not capture observable session-level characteristics.
-While traditional studies have employed human observers to provide such ratings, recent AI research has shown the validity of using finetuned LLMs in certain contexts to approximate session-level measures, such as working alliance [49], which might enable scalable behavioral analyses of more skill development constructs.
-Third, we recruited from a sample of bachelor-level counselors through the online Prolific platform, which may differ from in-person or classroom settings.
-For example, existing training environments for counselors will typically offer didactic materials (e.g., lectures and textbooks) and non-AI training opportunities (e.g., peer-to-peer roleplay and feedback).
-Thus, future work should evaluate the utility and acceptability of AI-based training tools within existing educational and training pathways.
-Finally, the efficacy of LLM-based practice and feedback training was only shown for client-centered microskills.
-Future work is required to understand whether these results generalize to other therapy modalities (e.g., psychodynamic, CBT/ACT), where training interventions may work to shift natural tendencies in similar or subtly different ways.
-3 Conclusion
-Despite increasing interest in using LLMs in mental health, to our knowledge, this is the first study to conduct a large-scale evaluation ($N=94$) of an LLM-based training system for developing core skills in novice counselors.
-Our findings show that while simulated practice (P) proved insufficient—and in the case of empathy, potentially harmful—practice with AI-generated feedback (P+F) supported measurable improvements in key counseling skills of demonstrating empathy, delivering reflective responses, and engaging in client-centered inquiry.
-By combining realistic patient simulations with expert-aligned, skill-specific feedback, the CARE AI Training system can help novices to master skills involved in client-centered therapy, offering a promising avenue for scalable, evidence-aligned training in mental health care.
-4 Methods
-4.1 Overview of CARE Training Platform
-Figure 5: CARE’s practice and feedback model visualized in a diagram and web screenshot.
+Our results show the practice-and-feedback group significantly improved in their use of reflections and questions (d=0.32-0.39, p<<0.05), and trended toward improvement in empathy (d=0.23) and suggestions (d=-0.28).
+In contrast, the practice-only group only showed significant improvements in suggestions (d=-0.39), but actually worsened in Empathy (d=-0.52, p=0.001).
+Between-group comparisons show a substantial advantage for the P+F group in Empathy (d=0.72, p=0.001), indicating a strong feedback effect.
+In contrast, suggestions showed near-zero between-group differences (d=0.02, p=0.910) despite pre-post improvements in both conditions, suggesting that another mechanism besides feedback is driving this change.
+Through qualitative analysis of participants’ self-reflections, we found that the practice-and-feedback group internalized the importance of empathetic and active listening; however, practice-only participants continued to overly focus on solutions, albeit with increased information-gathering.
+Our discussion details possible reasons for skill changes in the practice-alone group: novice counselors evolve therapeutic intentions from observable patient behaviors–when the AI patients consistently expressed skepticism to suggestions, counselors reduce inappropriate use of suggestions; but when AI patients show no differential response to empathetic vs.
+non-empathetic statements, the P group gradually de-prioritizes empathy.
+Together, these results suggest that LLM-simulated training should integrate structured feedback to cultivate a client-centered, empathetic listening approach fundamental to effective counseling.
+In summary, we contribute: (1) the design of CARE, an LLM-based training system that integrates realistic patient simulations with structured feedback grounded in counseling frameworks; (2) evidence from a randomized evaluation with 94 novice counselors, triangulating outcomes across behavioral performance, self-efficacy, and therapeutic intentions; and (3) design implications for LLM-simulated training, showing how structured feedback prevents empathy decline and supports effective counselor development, while highlighting ongoing challenges in improving overall self-efficacy while minimizing mis-calibration with performance.
+2. Related Work
+Our work training novice counselors via LLM-based systems is grounded in two areas of work.
+First, prior training approaches for clinical and communication skills have long relied on simulated patients, ranging from human role-plays to scripted virtual patients and, more recently, LLM-based simulations with automated feedback.
+Second, prior HCI research evaluating human-AI systems, especially in domains like health and education, emphasizes not only AI system accuracy but also how users learn, calibrate, and reflect when interacting with AI systems.
+2.1. Training Systems for Clinical and Communication Skills
+Traditional training for helping skills—empathy, active listening, and communication—uses theory, expert demonstration, role-play, supervised practice, and experiential learning (Hill, 2020; Hill and Kellems, 2002; Martin and Johnson, 2024; Huerta-Wong and Schoech, 2010).
+These approaches are effective but hard to scale: peers and supervisors require coordination, and trainees can pick up unhelpful habits without oversight (Atkins et al., 2014).
+Simulated standardized patients (trained actors) are common in health education; meta-analyses show they improve communication, knowledge transfer, and confidence (Shin et al., 2015).
+Counseling training likewise uses peer role-plays and standardized scenarios to practice skills like reflective listening and empathy (Kühne et al., 2020; Ay-Bryson et al., 2023), yet such exercises remain resource-intensive and limited in availability.
+Virtual patients
+Virtual patient (VP) simulations—computer or embodied agents—recreate clinical encounters to train history taking, nonverbal communication, empathy, and counseling (Albright et al., 2016; Schoenthaler et al., 2017; Rein et al., 2018).
+They provide safe, repeatable practice without actors and have been applied to suicide prevention, adolescent substance-use screening, and antibiotic conversations (Rein et al., 2018; Burmester et al., 2019; Schoenthaler et al., 2017).
+For example, Murali et al.
+(2022) used a conversational agent to teach vaccination-related counseling to laypeople; commercial platforms like Skillsetter follow deliberate-practice models (Skillsetter, 2024).
+Early VP systems were often scripted and template-based, costly to develop, and typically limited to single cases, reducing realism and diversity (T., 2003; Han et al., 2015; Othlinghaus-Wulhorst and Hoppe, 2020).
+Early AI-based training systems
+Building on VPs, early AI-driven systems experimented with automatically analyzing communication features and providing learners with feedback.
+For example, EQClinic visualized audio and video signals to help trainees reflect on their nonverbal behaviors in telehealth role-plays (Liu et al., 2016), while ConverSense detected and displayed social signals such as dominance and warmth from patient-provider interactions (Bedmutha et al., 2024).
+These systems raised self-awareness of communication styles, but their feedback was often decontextualized and difficult to apply in practice.
+Moreover, they did not directly target the counseling microskills important for effective therapeutic interactions.
+LLM-simulated patients for role-play practice
+Thus, a growing body of work in NLP and HCI has used LLMs to create simulated patients as role-play partners for counselor training (Wang et al., 2024; Louie et al., 2024; Stapleton et al., 2023; Steenstra et al., 2025).
+The goal of these systems is to provide practice environments that resemble real clinical encounters, making training more transferable and faithful to practice (Alinier and Oriot, 2022).
+However, achieving authentic simulations remains challenging.
+LLMs are highly sensitive to prompting (Zamfirescu-Pereira et al., 2023), and naive prompts in mental health contexts often produce unrealistic behaviors, including caricature, bias, and limited domain fidelity (Cheng et al., 2023).
+Chen et al.
+(2023) found that naively prompting GPT-3.5 to simulate a patient profile with depressive symptoms led the chatbot to describe its emotions in formal, diagnostic language, which expert clinicians noted as inauthentic.
+Recent work prompts LLMs with psychology-grounded frameworks to simulate patients (e.g., Patient-Psi for CBT case conceptualizations (Wang et al., 2024) and multi-stage pipelines modeling cognitive factors (Steenstra et al., 2025)).
+Nonetheless, LLMs remain prompt-sensitive and typically need expert validation to capture realistic resistance, ambivalence, and other clinically relevant behaviors.
+We therefore adopt the expert-driven, iterative behavioral refinement method of Louie et al.
+(2024)—integrating those validated principles into CARE to produce more authentic simulated patients.
+Automatic scoring and feedback for counselor transcripts
+A parallel line of work has developed automated methods to help peer counselors improve their skills.
+Scoring-based systems (e.g., ratio of questions to reflections in a transcript) provide metrics, but these approaches offer limited, actionable guidance on how to improve.
+By contrast, suggestion-based systems generate or rewrite candidate responses to model more effective behaviors.
+Research in clinical NLP has produced numerous models for classifying and scoring counseling transcripts (Tanana et al., 2016; Pérez-Rosas et al., 2019; Schwalbe et al., 2014; Huang et al., 2018; Yang et al., 2024; Min et al., 2022).
+Many focus on a single microskill, such as reflections, providing numeric feedback on usage frequency (Can et al., 2012; Pérez-Rosas et al., 2017; Shen et al., 2020; Min et al., 2022).
+Others examine skill distributions more broadly and their relationship to conversational success (Wang et al., 2023).
+While valuable for large-scale analysis, these approaches rarely translate into actionable feedback for scaffolding trainee learning.
+To make feedback more interactive, researchers have explored real-time rewriting and suggestion systems.
+For example, Saha et al.
+(2022) and Sharma et al.
+(2021) proposed response rewriting methods to enhance empathy.
+With the goal of increasing interactiveness,  Sharma et al.
+(2023) proposed HAILEY, a tool that modifies peer supporters’ responses, while  Hsu et al.
+(2025) generated strategy-aligned suggestions during live conversations.
+Although promising, studies show that just-in-time suggestions can distract learners and foster overreliance, sometimes leading to negative learning effects when AI support is withdrawn (Abbas et al., 2024; Bastani et al., 2025; Kazemitabaar et al., 2024).
+While this previous work developed NLP models for specific counseling tasks, the ability to use LLMs as zero-shot or few-shot reasoners has enabled further research in this area.
+Nonetheless, naively prompting LLMs in a mental health context can lead to generated outputs that are characteristic of low-quality therapy (Chiu et al., 2024).
+Therefore, a training system that uses LLMs to generate feedback for counselors needs to take measures to ensure the outputs are faithful and robust, lest it teach or promote bad practices (Moran et al., 2025).
+Chaszczewicz et al.
+(2024) co-designed a feedback dataset with therapy supervisors and fine-tuned an open-weight LLM to produce explanatory, actionable feedback.
+CARE adopts this expert-validated, fine-tuned model.
+While other training systems have employed prompting-based approaches with coding manuals and few-shot examples (Steenstra et al., 2025), our approach differs by integrating expert feedback directly into model weights through fine-tuning on expert-annotated examples, rather than relying on in-context learning alone.
+2.2. Evaluating Human-AI Systems
+Evaluating human-AI systems requires more than assessing model accuracy or output quality  (Blodgett et al., 2024).
+In HCI and education research, effectiveness is judged by its impact on learners: how people acquire skills, calibrate their understanding, and integrate feedback into practice.
+This framing is important in counseling training, where evaluation concerns not only usability but also the development of interpersonal behaviors in sensitive, high-stakes domains.
+Recent work highlights the limitations of traditional benchmarks, which often fail to capture generative model capabilities (McIntosh et al., 2026).
+This calls for dynamic and human-centered evaluations (Liao and Xiao, 2023; Elangovan et al., 2024; Khullar et al., 2025), that move beyond static model metrics and consider human outcomes, and are relevant when assessing interactive training systems.
+Thus, when we evaluate human-AI systems, additional challenges arise.
+Researchers must account for both the technical performance and also user impact (Weidinger et al., 2023).
+While guidelines exist for designing human-AI systems (amershi2019guidelines; Wright et al., 2020), less work addresses how they should be evaluated.
+Some frameworks capture process and user preferences in human-LLM interaction (Lee et al., ), others focus on safety (Weidinger et al., 2023) or domain-specific contexts (Lee et al., 2024).
+Tools such as SPHERE propose multi-dimensional evaluation cards to structure study design and improve transparency, but consensus on evaluation practices remains limited (Zhao et al., 2025).
+In counseling contexts, these gaps surface in three ways.
+First, evaluation must triangulate across behavioral outcomes, self-efficacy, and learning, aligning with evidence-based psychotherapy work in deliberate practice (Ericsson and others, 2006; Schön, 2017).
+Second, calibration is critical: learners often misjudge their own performance (Kruger and Dunning, 1999; Eva and Regehr, 2005), and AI feedback may inflate confidence without improving skills (Macnamara et al., 2024).
+Third, user perceptions of realism, trust, and workload shape adoption: relational agent studies show that authenticity fosters engagement (Mallik and Gangopadhyay, 2023), while trust research highlights risks of distraction and overreliance (Glikson and Woolley, 2020; Buçinca et al., 2021).
+Finally, in sensitive domains, evaluation must weigh ethical and pedagogical guardrails: ensuring feedback preserves learner agency and avoids harmful or misleading guidance.
+Taken together, evaluating human-AI systems requires a multi-dimensional perspective that integrates skill outcomes, calibration, perceptions, and responsible design.
+Yet, few studies have examined how LLM-based training systems affect novice counselors across these dimensions.
+Our work contributes by combining objective performance measures, self-efficacy surveys, and qualitative reflections to provide a holistic evaluation of LLM-driven counseling training.
+3. CARE Training System
+Figure 2.
+CARE’s practice and feedback model visualized in a web screenshot.
 In CARE, counselors practice with an LLM-simulated patient and receive feedback on each of their responses.
@@
 Responses with constructive feedback explain what the goal should be at this point in the conversation; what a helper could improve to better align with this goal; and how they could respond differently via an alternative response.
+The figure shows a screenshot of CARE’s practice and feedback interface.
+On the left, a text-based chat unfolds between an AI patient (depicted with a client avatar) and a counselor (depicted with a clinician avatar).
+The AI patient expresses frustration about his parents’ lack of trust and his struggles to prove he is a good father.
+The counselor responds twice: first with an empathetic statement that acknowledges the difficulty, and second with a suggestion-oriented prompt about showing efforts to change.
+On the right, CARE’s feedback system evaluates these responses.
+The first counselor response is labeled a Strength, tagged as “Empathy,” with positive feedback saying “Good response!” The second counselor’s response is flagged as a Feedback Area, tagged under “Suggestions” and “Questions.” The system provides an Alternative Response that rephrases the counselor’s statement to be more client-centered and exploratory, asking about the patient’s efforts rather than giving advice.
+Additional Feedback text explains that the goal at this point is to explore the seeker’s feelings and concerns rather than provide direct solutions, encouraging open-ended questioning.
 We developed CARE as a web platform for novice counselors to train in text-based counseling skills enabled by LLMs.
-Trainees practice simulated conversations with AI patients and review feedback on their responses generated by an AI mentor.
-CARE builds on top of successes from previous research in co-designing with mental health experts to improve the realism of LLM-simulated patients [50, 43, 14, 15] and finetuning domain-specific LLMs with therapeutic knowledge capable of generating feedback and alternative responses for text-based peer counseling conversations [51, 22, 52, 53, 54].
-CARE allows novice counselors to develop their counseling skills in a text-based format by practicing with AI-simulated patients and receiving feedback on their responses (see Fig. 5).
+The system integrates two core components: (1) LLM-simulated patients that provide realistic, text-based practice conversations, and (2) LLM-generated feedback that evaluates counselor responses against established skill frameworks and suggests improvements.
+Together, these features enable scalable, authentic training experiences that complement traditional, resource-intensive approaches such as role-play and supervision.
+CARE builds on top of successes from previous research in co-designing with mental health experts to improve the realism of LLM-simulated patients (Chen et al., 2023; Lin et al., 2024; Wang et al., 2024; Louie et al., 2024), using fine-tuned domain-specific LLMs trained on therapeutic knowledge capable of generating feedback and alternative responses for text-based peer counseling conversations (Sharma et al., 2021; Chaszczewicz et al., 2024; Shen et al., 2022; Min et al., 2022, 2023).
+Importantly, CARE was designed not only to identify whether a skill is used but also how well it is used, distinguishing, for example, between a reflection that captures a client’s core concern and one that misses the emotional nuance.
+CARE allows novice counselors to develop their counseling skills in a text-based format by practicing with AI-simulated patients and receiving feedback on their responses (see Fig.
+2).
 Consider Aki, a novice peer counselor who wants to use CARE to experience hands-on training using counseling skills that they have recently read about.
 In CARE, Aki can practice with an AI patient of their choice from a library of patient scenarios.
-Aki starts a practice chat with an AI patient with the background of a 35-year-old male veteran seeking to reconnect with his children but facing legal barriers and parental gatekeeping.
+Aki initiates a practice chat with an AI patient, a 35-year-old male veteran who is seeking to reconnect with his children but is facing legal barriers and parental gatekeeping.
 Each practice scenario provides limited background information about the AI patient and their presenting problem (e.g., ”Young adult with family issues: low mood and self-esteem”).
 This intentional limitation requires counselors to simultaneously learn more about the patient’s situation while demonstrating empathy and support.
+Patients are designed to exhibit realistic challenges, including resistance, ambivalence, or vague disclosures, drawing on behavioral principles elicited from expert counselors (Louie et al., 2024).
 Aki starts the conversation by greeting the AI patient.
-The AI patient shares their frustration with feeling everyone is against him, and hopes to find ways get his kids back and overcome the difficulty his judgmental parents.
+The AI patient expresses frustration about feeling that everyone is against them, hoping to find ways to reunite with their kids and overcome the challenges posed by their judgmental parents.
 Aki composes a reply, and the conversation continues in this turn-by-turn manner.
-Users can view feedback on their therapeutic responses at any point during or after the session, offering flexibility to either review intermittently throughout the practice or comprehensively after completion.
-4.2 Implementation details for CARE Training Platform
-CARE was built as a web application with a Python Flask back-end and React Javascript front-end which online participants could access and interact with.
-LLM-simulated patients in CARE were powered by the GPT-4o API with a custom prompt defining the patient’s background scenario and behavioral principles [15].
-The prompts for the three AI patients used for the pre-intervention assessment, intervention, and post-intervention are provided in the Supplementary Materials A.1.
-The LLM feedback system was powered by a Llama-2 13B parameter model finetuned on an expert-annotated feedback dataset [22].
-4.3 Study Procedures
-Between August and October 2024, we conducted an online lab study with (N=94) novice counselors.
-Given our aim to train novices lacking formal counseling training, we recruited participants who expressed an interest in practicing helping skills yet had minimal prior education or practical experience.
-Specifically, we used the Prolific platform to screen for participants who (1) lacked extensive formal professional or graduate training in clinical psychology, counseling, social work, or nursing (excluding those who were pursuing or had completed master’s or doctorate degrees); and (2) had less than one year of counseling-related experience (e.g., peer support or crisis counseling).
-In our final experiment, we recruited N=94 participants.
-Additionally, during the pilot phase, we recruited another N=14 participants who were allowed broader educational and hands-on experiences for protocol refinement, but we report qualitative and quantitative results exclusively from the final set of 94 participants.
-Each participant completed a 75-90 minute video-call session following the structure outlined in Fig.
-1.
-After receiving a static 5-minute tutorial on core counseling skills (see Fig.
-6), all participants completed a 10-minute pre-intervention assessment chat with the first AI patient.
-For the 20-minute training intervention, participants were randomly assigned to either practice with an AI patient only or practice with the option to receive feedback on all of their sent responses.
-Finally, all participants completed a 10-minute post-intervention assessment chat with a third AI patient.
-Participants answered self-efficacy items after the pre-intervention and post-intervention chats, and provided qualitative self-reflections after the training intervention chat and post-intervention chat.
-All participants were paid \$15/hour.
-To protect participant privacy, we instructed them to join the video-call using their ProlificID as the display name, turn off their video-camera, and share their screen.
-Participants consented for the video call to be recorded for our own analysis and their CARE usage data to be used and shared for the purposes of this and future research after de-anonymization.
-This study protocol was approved by our institution’s IRB.
-4.4 Assessing Changes in Skill Use through LLM Classifiers
-We investigate whether counselors employ higher-quality counseling behaviors after LLM-based training and between training conditions by leveraging NLP methods.
-This analysis is motivated by the need to quantify changes in counseling skill use at scale across multiple participant sessions.
-Our approach follows four key steps: (1) fine-tuning and validating LLM-based classifiers to identify skill behaviors, (2) selecting a final set of classifiers based on performance metrics and theoretical priority, (3) testing changes in skill use across time and between intervention groups, and (4) analyzing the relationship between exposure to good alternative patterns in AI feedback and subsequent skill use.
-4.4.1 Fine-tuning and Validating LLM-based Classifiers
-We developed LLM-based binary classifiers that allows us to label the skill use within a transcript.
-For example, one of our fine-tuned classifier could determine which utterances in a transcript had strong uses of Questions during that stage of the transcript.
-To finetune and evaluate these classifiers, we transformed a previously published expert-annotated, feedback dataset that had strengths and areas of improvements for 8 skills [22] into a 16-class binary classification format (8 skills × 2 categories each—strong uses and areas needing improvement) [^1] The binary classification feedback dataset can be accessed at https://huggingface.co/datasets/youralien/feedback_qesconv_16wayclassification.
-Additionally, we used a subset of transcripts from this randomized study which were further annotated by counseling domain-experts.
-Three experts were recruited with relevant backgrounds including practicing clinical psychologist, licensed marriage family therapist, former director and supervisor of a crisis agency.
-Each of the experts annotated 10 participants’ study transcripts (5 from the practice-only group; 5 from the practice-and-feedback group), totaling 370 counselor utterances.
-Two rounds of annotation occurred: after collecting a first annotation pass and identifying data points with disagreements, we showed each of the experts the others’ annotations and had them re-annotate and provide rationales for their decision.
-We display pairwise agreement results averaged across all pairs in Table 2.
-Note that while we originally explored other annotation agreement metrics such as Cohen’s kappa, the severe class imbalance of our annotations made these metrics less relevant in our case.
-The gold-standard CARE 10% expert-annotated sample consists of labels that result from a majority vote across these three experts [^2] This expert-annotated data sample can be found at https://huggingface.co/datasets/youralien/CARE_10percent_16wayclassification.
-We finetuned RoBERTa-large binary classifiers using a 95% split of the transformed FeedbackQESConv dataset, and did hyperparameter tuning on 5% of FeedbackESConv and the CARE 10% sample.
-The performance of the our LLM-based classifiers are shown in Table 2.
-The highest performing classifiers ($F1>0.5$) became candidates for our automatic behavioral assessments, which we further down-selected as described below.
+After completing the conversation, Aki reviews AI-generated feedback.
+CARE highlights strengths such as asking open-ended questions, but also flags missed opportunities for empathy, offering alternative phrasings, and a rationale for why they may better support the client.
+Unlike real-time AI ”co-pilot” systems, which may proactively suggest responses, CARE provides feedback only after the user has sent their response in the simulated dialogue.
+A user can view feedback on their therapeutic responses at any point.
+This offers flexibility to review feedback intermittently throughout the practice or comprehensively after completion.
+This design choice mirrors human supervision: it preserves the learner’s agency during the conversation while supporting reflection afterward.
+Feedback targets core microskills, empathy, reflections, questions, validation, and suggestions, based on established counseling frameworks (Nelson-Jones, 2013; Miller et al., 2003).
+3.1. Implementation details for CARE Training Platform
+CARE was built as a web application with a Python Flask back-end and React JavaScript front-end, accessible through any standard browser. The two core components of CARE are described below.
+LLM-simulated patients.
+CARE implements an existing method for simulated patients that co-designed patient prompts with expert-counselors, each specifying (a) a challenging patient scenario that they had encountered in past clinical and text-based counseling settings, including demographic background presenting issues or symptoms, etc.
+and (b) Constitutional AI principles elicited from expert counselors for defining authentic patient behaviors (Louie et al., 2024).
+This simulation method uses the OpenAI GPT-4o API to role-play patient scenarios and behaviors due to its strong ability to maintain role consistency and instruction-follow expert-defined principles.
+These principles instructed the simulated patients to display realistic challenges such as resistance to suggestions ("Respond to encouraging words with hesitation, doubting their significance"), low awareness ("Don’t be so self-aware or good at recognizing your own problems"), or minimal disclosure ("Use more colloquial language and express reluctance to open up").
+Simulated patients were designed and validated for 10-20 minute text-based counseling conversations, allowing participants to both explore the presenting problem and practice multiple client-centered microskills.
+CARE implements the top-rated patients from Louie et al.
+(2024)’s study which were judged by third-party counselors as being the most ready to be used as a training partner ($\geq 6$ average score on a 7-point Likert-scale).
+All prompt templates are available in Supplementary Materials A.1.
+The three AI patients used in our experiment were created by experienced counselors from the USA with extensive backgrounds in mental health support (Louie et al., 2024).
+Each patient was deliberately varied to expose trainees to diverse yet comparably challenging scenarios:
+  • Patient 1 (35-year-old American male experiencing holiday loneliness): Created by an experienced peer counselor from an online counseling platform.
+Holiday loneliness is common there and often stems from family estrangement.
+In this case, the patient’s isolation intensifies during winter holidays when family gatherings occur.
+  • Patient 2 (35-year-old male veteran in court-mandated therapy seeking to reconnect with children): Created by a Licensed Marriage and Family Therapist (LMFT), a white woman aged 30-40, based on her clinical experience with veterans.
+This scenario reflects a patient frustrated by limited access to his children due to parental and legal issues stemming from substance abuse.
+  • Patient 3 (Young adult with family issues, low mood, and self-esteem concerns): Created by a US-basd clinical psychology doctoral student.
+This adolescent patient faced self-esteem issues stemming from family dynamics, where her parents favored a sibling, leading to symptoms of anhedonia and depression, with a diminished ability to enjoy previously pleasurable activities.
+LLM-generated feedback.
+CARE integrates an existing method for generating counseling feedback that fine-tunes and self-improves the Llama-2 13B parameter model using an expert-annotated feedback dataset of peer counseling transcripts  (Chaszczewicz et al., 2024).
+We selected this existing method because it provides faithful generation of counseling feedback grounded in the content and style of feedback given by psychotherapy supervisors.
+An additional key benefit, as argued by Chaszczewicz et al.
+(2024), is that a fine-tuned open-weight model can operate on therapy data in a controlled, private environment rather than relying on external API services.
+This model generates feedback at multi-levels: (1) classify the trainee response against eight microskills (empathy, reflections, questions, validation, suggestions, session management, professionalism, and self-disclosure), (2) assess quality by highlighting strengths and areas needing improvement, and (3) generate alternative responses and explanatory rationales, enabling trainees to compare their choices against more client-centered approaches.
+This post-practice feedback design mirrors human supervision: it preserves agency during the conversation while supporting reflection and skill refinement afterward.
+4. Randomized Pre-Post Study
+The core goal of CARE is to upskill novice counselors through LLM-simulated training.
+In a randomized experiment, we investigated how CARE’s core components–practicing with LLM-simulated patients and receiving AI-generated feedback on their responses–are important for participants’ skill development.
+We conceptualize skill development holistically, encompassing three complementary dimensions: (1) behavioral performance, where a trainee is judged on their appropriate use of counseling skills in a representative scenario or conversation; (2) counseling self-efficacy, defined as a trainee’s self-assessments of their own abilities; and (3) therapeutic intentions, or the goals that participants form in-session, which should be adherent with evidence-based procedures.
+The experiment evaluated how skill development changed over time when participants were assigned to two variants of LLM-based counseling training: practicing with LLM-simulated patients alone (Group P); or practicing with simulated patients while also receiving LLM-generated feedback (Group P+F).
+In contrast to other work on simulation in clinical education (Shin et al., 2015), our primary interest was in understanding what aspects of LLM-simulated training could promote skill development, as part of a broader research agenda to iteratively design and improve AI-based counseor training.
+Given this, we did not include a ”no AI” control group.
+Beyond the skill development outcomes, we also conduct a mixed-methods investigation of participants’ experience of CARE’s LLM-based components and their perceived value of such training experience, since user perceptions shape adoption in training contexts.
+Together, our experiment sought to answer six research questions covering skill development and training experience with CARE from quantitative and qualitative perspectives.
+  • RQ1: How does CARE’s LLM-simulated practice and feedback affect participants’ behavioral performance?
+  • RQ2: How does CARE’s LLM-simulated practice and feedback affect participants’ self-efficacy?
+  • RQ3: How does CARE’s LLM-simulated practice and feedback affect novice counselors’ therapeutic intentions?
+  • RQ4: What are participants’ quantitative experience of CARE’s LLM-simulated practice and feedback?
+  • RQ5: What are participants’ qualitative experience receiving feedback on their responses from CARE?
+  • RQ6: What are participants’ qualitative experience practicing with CARE’s LLM patients?
+4.1. Participants
+We recruited $N=94$ novice counselors on the Prolific platform using specific filtering criteria to select US and UK participants with some interest in the field but limited access to formal training.
+Eligible participants were required to have (1) an educational background in psychology, counseling, social work, or nursing, with educational attainment limited to those who had completed at most a bachelor’s degree or were currently pursuing a master’s degree, and (2) less than one year of counseling-related experience (e.g., peer support or crisis counseling volunteering).
+Prolific participants were paid \$15/hour.
+We conducted sessions with 108 participants from Prolific.
+The first 14 were part of a pilot that refined our recruiting criteria and protocol (e.g., excluding counselors with graduate degrees).
+Our analyses use the final 94 participants, though we reference participants by their original identifiers.
+For the final participant pool, 68% were located in the United States and 32% in the United Kingdom.
+The sample was predominantly female (68%), with 31% male participants and 1% preferring not to disclose gender.
+The median age was 29 years (IQR: 23-39).
+Regarding ethnicity, 49.5% of participants identified as White, 16.2% as Black, 15.3% as Multiracial, 13.5% as Asian, and 5.4% as Other.
+Participants’ primary fields of study included psychology (66%), social work (24%), nursing (16%), and counseling (10%), with participants able to select multiple areas.
+In terms of educational attainment, 22.4% had no formal education in these fields, 50.6% were currently pursuing undergraduate degrees, 12.9% had completed only bachelor’s degrees in relevant fields, and 14.1% were pursuing master’s degrees.
+To protect participant identities, our IRB-approved protocol instructed participants to use their Prolific email address, turn off cameras, and change their Zoom display name to their ProlificID.
+During recruitment and consent, we warned participants of the potentially stressful nature of simulated patient situations and ensured scenarios avoided especially sensitive topics such as suicidal ideation.
+4.2. Power Analysis
+To determine the appropriate sample size for our randomized pre-post study, we conducted a power analysis targeting a medium effect size with adequate statistical power.
+Our analysis was based on a repeated-measures design comparing pre-intervention and post-intervention outcomes between two groups (practice-only vs.
+practice-with-feedback).
+We selected an effect size of $d=0.4$ as our target, representing a conservative estimate for behavioral skill improvements.
+This choice was informed by previous research on social skills training interventions, where studies examining changes in behavioral performance have reported medium to medium-high effects ranging from $d=0.5$ to $d=0.6$ (Lin et al., 2024).
+Using power analysis calculations for two-sample, repeated-measures designs with $\alpha=0.05$ and $\beta=0.8$ (80% power) in the R statistical analysis software, our calculations indicated that $N=94$ participants would provide sufficient statistical power to detect our target effect size.
+4.3. Study Setup
+The study flow is illustrated in Figure 1.
+The 75-minute study was conducted over Zoom.
+Participants first read a 5-minute tutorial refreshing foundational counseling skills, then completed a timed 10-minute pre-intervention chat with the first AI patient.
+For the 20-minute main intervention, we randomized participants into two groups: (1) Group P: Practice with an LLM-simulated patient without AI feedback, or (2) Group P+F: Practice with an LLM-simulated patient with AI feedback.
+Group P+F participants could review AI feedback on their responses at any time.
+The experimenter provided verbal reminders to check feedback at 5 minutes and to review remaining feedback at 15 minutes.
+Participants then completed a 10-minute chat with the third AI patient.
+Surveys were administered after each chat period.
+Upon completing the post-intervention chat and self-efficacy assessment, participants shared their experience with the CARE training tool via survey and semi-structured interview.
+Group P participants received 5 additional minutes to interact with AI feedback on their post-intervention chat before sharing perceptions.
+Since this occurred after the skill acquisition experiment, it does not interfere with training effectiveness results (RQ1-3) but allows us to ask all 94 participants their perceptions of both AI patients and AI feedback in CARE (RQ4).
+4.4. Measures
+To understand whether simulated practice alone (P) and practice with feedback (P+F) can upskill novice counselors, we integrate evidence from three sources of data: automatic assessments of behavioral performance (RQ1), participants’ assessments of their self-efficacy (RQ2), and qualitative self-reflections about their therapeutic intentions (RQ3).
+Following the post-intervention, we conducted a final survey and semi-structured interview with participants to understand their perceptions of the CARE system and its features (RQ4).
+4.4.1. RQ1. Automatic Assessment of Behavioral Performance
+We assess whether counselors employ higher-quality counseling behaviors in transcripts by leveraging NLP methods.
+This automatic assessment is motivated by the need to quantify changes in counseling skill use at scale across multiple participant sessions.
+Our automatic assessment approach requires (1) fine-tuning and validating LLM-based classifiers to identify skill behaviors, and (2) selecting a final set of classifiers based on statistical-testing considerations, performance metrics, and theoretical priority.
+In the following paragraphs, we explain both of these steps in more detail.
+Ultimately, we assessed behaviors of skills used for the exploration stage (strong uses in empathy, reflections, questions) and action stage (suggestions needing improvement) of Hill’s Helping Skills framework (Nelson-Jones, 2013); see Table 2 for definitions.
+Fine-tuning and Validating LLM-based Classifiers.
+We developed LLM-based binary classifiers to label skill use within transcripts.
+For example, one classifier determines which utterances showed strong use of Questions.
+To finetune and evaluate these classifiers, we transformed a previously published expert-annotated feedback dataset (Chaszczewicz et al., 2024) into 16-class binary classification format (8 skills × 2 categories: strong uses and areas needing improvement).
+[^1] The binary classification dataset for FeedbackESConv can be accessed at https://huggingface.co/datasets/youralien/feedback_qesconv_16wayclassification Additionally, we used a subset of transcripts from this study, annotated by three counseling domain-experts: a practicing clinical psychologist, licensed marriage family therapist, and former director and supervisor of a crisis agency.
+Each expert annotated 10 participants’ transcripts (5 from each group), totaling 370 counselor utterances.
+After an initial pass, we showed experts each other’s annotations for disagreement points and had them re-annotate with rationales.
+Table 1 shows pairwise agreement results averaged across all pairs.
+While we initially explored metrics like Cohen’s kappa, severe class imbalance made them less relevant.
+The CARE expert-annotated sample (10% of participants’ transcripts) consists of majority-vote labels across the three experts.
+[^2] This expert-annotated data sample can be found at https://huggingface.co/datasets/youralien/CARE_10percent_16wayclassification
+We finetuned RoBERTa-large binary classifiers using FeedbackQESConv, a dataset of transcripts from emotional support conversations between peer counselors on a crowdsourcing platform annotated with multi-level counseling feedback (Chaszczewicz et al., 2024), allocating 95% of this data for training.
+For hyperparameter tuning, we used a validation set comprising 5% of the FeedbackQESConv dataset (n=409) combined with our CARE expert-annotated sample (n=370).
+The performance of our LLM-based classifier candidates varied by skill, as shown in Table 1, motivating us to down-select a final set of classifiers for our planned analyses.
 Skill Strengths Areas to Improve
@@
 Professionalism 0.905 0.969
-Table 2: Left columns show pairwise annotator-agreement averaged across 3 domain-experts for the CARE 10% sample (n=370).
-Right columns show performance of the best RoBERTa-large classification models after hyperparameter tuning on our validation dataset, CARE 10% sample (n=370) + FeedbackQESConv 5% sample (n=409).
+Table 1.
+Annotator agreement columns show pairwise agreement averaged across 3 domain-experts for the CARE expert-annotated sample (n=370).
+Classifier performance columns show performance of the best RoBERTa-large classification models after hyperparameter tuning on our validation dataset, CARE expert-annotated sample (n=370) + FeedbackQESConv 5% sample (n=409).
 Session Management and Professionalism were excluded from finetuning due to infrequent occurrence.
-4.4.2 Down-selecting a Final Set of Classifiers
-From the initial set of 16 binary classifiers, we applied both methodological and theoretical criteria to select our final set for analysis.
-First, we established a minimum performance threshold of $F1>0.5$ to ensure reliable classification.
-This criterion yielded seven candidate classifiers: strong uses of Empathy, Reflections, Questions, and Validation, as well as both strong uses and areas needing improvement for Suggestions.
-To maintain statistical power while controlling for multiple comparisons, we further narrowed our focus to four key classifiers: strong uses of Empathy, Reflections, and Questions, plus areas needing improvement for Suggestions.
-This selection was guided by three considerations: (1) the need to limit the number of statistical comparisons to avoid diluting significance across too many tests, (2) focusing on classifiers with the strongest performance metrics, and (3) selecting skills that represent core competencies in client-centered frameworks and are frequently used in counseling sessions.
-Self-disclosure was excluded from our analysis due to its infrequency in our dataset, while Validation, though conceptually related to Empathy and mentioned frequently in qualitative data, showed more limited classifier performance and was therefore reserved for secondary analyses.
-4.4.3 Testing Changes across Time and between Groups
-Using our selected classifiers, we examined how skill usage changes from the pre-intervention chat to the post-intervention chat.
-For each session transcript, we computed an aggregate score defined as the proportion of utterances with strong uses of a skill (e.g., $b=N_{strengths}/N_{total}$) or with areas needing improvement (e.g., $b=N_{improvement}/N_{total}$).
-To statistically test for changes between the pre-intervention and post-intervention chats ($b_{0},b_{1}$), we used paired t-tests and computed Cohen’s $d$ effect sizes.
-To test for significant differences between P and P+F groups ($b_{1}^{P}-b_{0}^{P}$ vs.
-$b_{1}^{PF}-b_{0}^{PF}$), we used non-paired t-tests.
-In total, we conducted 12 planned t-tests for analyzing changes in behaviors: three skills (Empathy, Reflections, Questions) with strong uses and one skill (Suggestions) with areas needing improvement, each analyzed for both within-group changes and between-group differences in those changes.
-4.4.4 Exposure to Good Alternatives in AI Feedback
-To better understand AI feedback’s role in skill development, we analyzed whether the level of exposure to specific feedback during practice affected post-intervention performance.
-We quantified the exposure to Good Alternatives during Practice (GAP), defined as the number of alternative responses in the AI-generated feedback that exhibited strong uses of a skill, normalized by chat length.
-To test whether greater exposure to GAP for a skill led to stronger use of that skill, we fit a lagged-linear regressor that predicts post-chat behaviors, controlling for pre-chat behaviors, with exposure to good alternatives ($T_{GAP}$) as a scalar predictor ($b_{1}=\beta_{0}+\beta_{1}b_{0}+\beta_{2}T_{GAP}$).
-This model allowed us to isolate the effect of feedback exposure while accounting for baseline performance differences.
-4.5 Analyzing Confidence and the Calibration of Self-Assessments using Counselor Self-Efficacy
-Counselor self-efficacy represents a critical dimension in counselor development, defined as one’s “beliefs or judgments about their ability to effectively counsel a client” [55].
-While high self-efficacy can promote persistence and resilience when facing challenges [56], miscalibrated self-assessments—particularly overconfidence relative to actual ability—can lead to suboptimal or potentially harmful counseling behaviors [57, 58, 59].
-This calibration challenge has prompted research into potential remedies.
-Prior studies suggest that increased practice experience may help individuals better align their self-assessments with actual performance, though results remain inconsistent [60].
-Similarly, supervisor feedback has been proposed as a mechanism to address self-assessment miscalibration [61], although evidence from other educational contexts shows mixed effectiveness [62].
-Our methodology specifically examines whether practice alone or practice with structured feedback differentially impacts both raw self-efficacy and its calibration with actual performance.
-4.5.1 Self-Efficacy Measurement Instruments
-To measure counselor self-efficacy, we employed the Counselor Activity Self-Efficacy Scale (CASES) [37], specifically utilizing a revised subset of items targeting basic counseling skills (CASES-R) [63, 36].
-The CASES-R established a three-factor structure to assess counselors’ confidence in performing key therapeutic functions: Exploration and Insight Skills (EIS), Action Skills (AS), and Session Management Skills (SM).
+Down-selecting a Final Set of Classifiers
+From the initial set of 16 binary classifiers, we narrowed our focus to four key classifiers: strong uses of Empathy, Reflections, and Questions, and inappropriate uses of Suggestions.
+This selection was guided by selecting (1) the highest performing classifiers based on F1 scores (2) fewer classifiers for statistical concerns since our analyses would control for a false discovery rate based on number of skill hypotheses tested; and (3) theoretical relevance of skills most frequently emphasized in client-centered counseling textbooks and used during training with CARE.
+The final four skill classifiers have F1 performance scores between 0.56 and 0.77.
+Detailed selection criteria and rationale are provided in Appendix A.2.
+4.4.2. RQ2. Counseling Self-Efficacy
+To measure counselor self-efficacy, we employed the Counselor Activity Self-Efficacy Scale (CASES) (Lent et al., 2003), specifically utilizing a revised subset of items targeting basic counseling skills (CASES-R) (Hahn et al., 2021; Hunsmann et al., 2024).
+The CASES-R established a three-factor structure to assess counselors’ confidence in performing key therapeutic functions: Exploration and Insight Skills, Action Skills, and Session Management Skills.
 Participants completed the CASES-R immediately following both pre-intervention and post-intervention AI patient interactions.
 All items were administered using an 8-point Likert scale (0 = no confidence, 7 = complete confidence).
-During factor analysis, we discovered that among the five original EIS skills, self-disclosure did not load on the same factor as the other items.
-Consequently, we consolidated the EIS dimension to include only four Exploration Skills: Reflections of Feelings, Restatements, Open Questions, and Listening.
-The final instrument comprised 12 items across three factors: (1) exploration skills (e.g., restatements, reflecting feelings, open questions, listening); (2) actions skills (e.g., providing suggestions, knowing which actions to take); and (3) session management skills (e.g., keeping sessions on track).
-Reliability analysis demonstrated good to excellent internal consistency across all factors, with Cronbach’s $\alpha$ values of 0.784, 0.803, and 0.905 for ES, AS, and SM, respectively.
-4.5.2 Dunning-Kruger Effect Analysis
+During factor analysis, we discovered that among the five original Exploration and Insight Skills, self-disclosure did not load on the same factor as the other items.
+Consequently, we consolidated the Exploration and Insight Skills dimension to include only four Exploration Skills: Reflections of Feelings, Restatements, Open Questions, and Listening.
+The final instrument comprised 12 items across three factors: (1) exploration skills (e.g., restatements, reflecting feelings, open questions, listening); (2) action skills (e.g., providing suggestions, knowing which actions to take); and (3) session management skills (e.g., keeping sessions on track).
+To assess the internal consistency of each factor, we conducted reliability analysis using Cronbach’s $\alpha$, which measures how closely related a set of items are as a group (Tavakol and Dennick, 2011).
+The analysis demonstrated good to excellent internal consistency across all factors, with Cronbach’s $\alpha$ values of 0.784, 0.803, and 0.905 for exploration skills, action skills, and session management skills, respectively.
+Stages Skill Category Description
+Questions Questions seek information from the client and can be open (inviting elaboration) or closed (requesting specific answers).
+They include both direct questions and indirect prompts (e.g., ”Tell me about…”).
+Reflections Reflections capture and return to clients something they have communicated, either explicitly or implicitly.
+They typically mirror back content from the client’s preceding statement, but can also reference earlier parts of the conversation.
+Empathy Empathy can be shown through emotional warmth, interpretation of the client’s experience (e.g., paraphrasing, making conjectures, or sharing relatable experiences), or exploration of the client’s feelings and perspectives.
+Suggestions Suggestions offer possible actions, perspectives, or solutions in a respectful and autonomy-supportive manner. They may involve information-sharing or proposing alternative viewpoints.
+Session Management Session management includes organizing the session, transitioning between topics, and summarizing key points. It provides structure and helps maintain therapeutic focus.
+Table 2.
+Overview of our analysis of skill development, grounded in Hill’s Helping Skills model (Nelson-Jones, 2013).
+We select a skill subset relevant for beginning counselors at the undergraduate and first-year graduate level (Hunsmann et al., 2024).
+These include microskills during the exploration and action stages; and macro skills that are applicable throughout the session.
+Hill’s insight stage, of which self-disclosure was the only relevant skill for basic counseling, was excluded from our primary analyses due to its infrequent occurrence in our data.
+4.4.3. RQ3. Qualitative Self-Reflections on Therapeutic Intentions
+LLM-simulated training provides opportunities for experiential learning (Huerta-Wong and Schoech, 2010) whereby reflection on action (Schön, 2017) can support counselors in refining their therapeutic intentions and strategies.
+To study this impact on participants’ intentions, we collected qualitative self-reflections from two time points: immediately after the training intervention chat, where participants responded to ”What would you do differently as a therapist?” and after the post-intervention chat, where they reflected on ”What did you do well as a therapist?”.
+We examined how initial intentions translated into reported strengths across the P+F and P groups.
+4.4.4. RQ4. Quantitative Experience of CARE’s LLM Feedback and Simulated Practice
+Three survey questions measured participants’ perceptions of CARE’s AI feedback system on a 5-point Likert scale.
+Helpfulness.
+Participants rated ”To what extent do you find the AI feedback to be constructive and helpful?”.
+Comfort.
+Participants rated ”To what extent do you agree with the following statement: ’I am comfortable receiving AI feedback’”.
+Readiness.
+Participants rated ”The AI feedback system is ready to be used by counselors-in-training.” Participants in the P+F group answered these questions after the intervention chat.
+Participants in the P group also received feedback–only after the pre-post experimental measures were completed–and subsequently answered these three questions about AI feedback.
+We measured participants’ perceptions of each of the AI patients after each chat (pre-intervention, practice intervention, post-intervention) with several 7-point Likert scale items.
+Authenticity.
+Participants rated ”The AI patient was authentic in its role.” Four questions from the NASA-TLX workload scale were given after each simulated practice: Mental Demand: ”How mentally demanding was giving counseling support to this patient?”; Temporal Demand: ”How hurried or rushed did you feel giving counseling support to this patient?”; Effort: ”How hard did you have to work to accomplish your level of performance”; Frustration: ”How discouraged or stressed were you while giving counseling support to this patient?”.
+4.4.5. RQ5 and RQ6. Qualitative Experience of using CARE’s LLM Feedback and Simulated Practice
+Participant interview data was collected at different time-points throughout the 75-minute session.
+To specifically understand the experience receiving feedback from CARE (RQ5), participants were asked to elaborate in more detail their answers after filling out the three survey questions about CARE’s feedback; re-review the AI feedback page and think-aloud about their agreements or disagreements with any of the feedback; and explain whether the feedback had any bearing on their self-reflections about what they did well or wanted to do differently as a counselor.
+To specifically understand the experience practicing with CARE’s simulated patients (RQ6), participants were given the chance to explain in more detail their answers to the quantitative survey items about the simulated patients.
+Finally, a semi-structured exit-interview was conducted for all participants which was framed around the following questions: ”What do you like about this training tool for helping skills?” ”What do you wish was different about the training tool?” and ”What suggestions do you have for improving any part of the training tool?”.
+4.5. Analyses
+4.5.1. RQ1. Effects on Behavioral Performance
+Our analysis of behavioral performance consists of two perspectives: (1) testing changes in behaviors of skills used across time (pre-intervention vs.
+post-intervention) and between intervention groups (P vs.
+PF); and (2) analyzing the relationship between intervention-exposure to good alternative patterns in AI feedback and post-intervention skill use.
+Testing Changes Across Time and Between Groups
+Using our selected classifiers, we examined skill usage changes from pre- to post-intervention.
+For each transcript, we computed the proportion of utterances showing strong skill use ($b=U_{\text{strengths}}/U_{\text{total}}$) or needing improvement ($b=U_{\text{improvement}}/U_{\text{total}}$).
+To test for pre-post changes ($b_{0},b_{1}$), we used paired $t$-tests and Cohen’s $d$ effect sizes.
+To compare P and P+F groups ($b_{1}^{P}-b_{0}^{P}$ vs.
+$b_{1}^{PF}-b_{0}^{PF}$), we used unpaired $t$-tests.
+We conducted 12 planned $t$-tests: three skills (Empathy, Reflections, Questions) for strong uses and one skill (Suggestions) for areas needing improvement, analyzing both within-group changes and between-group differences.
+Exposure to Good Alternatives in AI Feedback
+To assess whether AI feedback exposure affected post-intervention performance, we defined Good Alternatives during Practice (GAP) as the proportion of trainee utterances for which the AI suggested an alternative response that exemplified a strong use of a skill:
+$T_{GAP}=\frac{A_{strengths}}{U_{total}},$
+where $A_{strengths}$ is the number of AI-generated alternative responses judged as strong exemplars and $U_{total}$ is the total trainee utterances.
+We then fit a lagged linear regression predicting post-chat behavior $b_{1}$ while controlling for pre-chat behavior $b_{0}$ and including GAP as a predictor:
+$b_{1}=\beta_{0}+\beta_{1}b_{0}+\beta_{2}T_{GAP},$
+thereby isolating the effect of feedback exposure.
+4.5.2. RQ2. Effects on Self-Efficacy and its (mis)calibration with Behavioral Performance
+First, we test for changes in raw self-efficacy scores after practice or practice-and-feedback.
+Second, we examine the calibration of self-efficacy ratings with actual performance.
+Third, we evaluate whether P or P+F interventions improve this calibration.
+Changes in Raw Self-Efficacy
+Beyond calibration, we also investigated whether the interventions affected participants’ absolute levels of self-efficacy across the three measured dimensions (exploration skills, action skills, and session management skills).
+We conducted repeated-measures analyses to identify: (1) significant pre-post changes in raw self-efficacy scores following practice alone (P intervention); (2) significant pre-post changes in raw self-efficacy scores following practice with structured feedback (P+F intervention); and (3) differential patterns of change between the P and P+F groups, indicating potential intervention-specific effects on self-efficacy development.
+Investigating (mis)calibration of Self-Efficacy
 Our primary analysis investigated potential mis-calibration between participants’ self-assessments and their actual counseling performance, specifically examining whether data exhibited patterns consistent with the Dunning-Kruger effect.
-This phenomenon [38] suggests that individuals with lower skill levels tend to overestimate their abilities, while highly skilled individuals may slightly underestimate their competence.
-We focused this analysis on Exploration Skills (ES) and Action Skills (AS), as these dimensions had straightforward mappings between CASES items and our NLP behavioral classifiers (Table 3).
-To test for the presence of the Dunning-Kruger effect, we follow the classic analysis method that splits the data into quartiles based on performance and conducts a two-way analysis of variances for self-assessments and actual performance and across the quartiles; and finally verifies via post-hoc tests that the bottom performers have the biggest overestimation of their abilities [38].
+This phenomenon (Kruger and Dunning, 1999) suggests that individuals with lower skill levels tend to overestimate their abilities, while highly skilled individuals may slightly underestimate their competence.
+We focused this analysis on Exploration Skills and Action Skills, as these dimensions had straightforward mappings between CASES items and our NLP behavioral classifiers (Table 7).
+To test for the presence of the Dunning-Kruger effect, we follow the classic analysis method that splits the data into quartiles based on performance and conducts a two-way analysis of variances for self-assessments and actual performance across the quartiles; and finally verifies via post-hoc tests that the bottom performers have the biggest overestimation of their abilities (Kruger and Dunning, 1999).
 To standardize the comparison between self-efficacy and performance, we transform each measure into a percentile rank (0 - 99) computed across all data collected for the pre-intervention and post-intervention chats ($b_{0},b_{1}$; $s_{0},s_{1}$).
-4.5.3 Calibration Change Analysis
+Changes in Calibrated Self-Efficacy
 To evaluate whether our interventions improved self-efficacy calibration, we computed discrepancy scores by subtracting standardized performance scores from standardized self-efficacy scores for each participant at both assessment timepoints.
-These discrepancy metrics provided a direct measure of calibration, with positive values indicating overconfidence and negative values indicating under-confidence.
-We then examined changes in these discrepancy scores from pre- to post-intervention for both intervention groups.
-This analysis allowed us to determine whether either intervention significantly improved the alignment between participants’ self-perceptions and their actual counseling abilities.
-4.5.4 Changes in Raw Self-Efficacy
-Beyond calibration, we also investigated whether the interventions affected participants’ absolute levels of self-efficacy across the three measured dimensions (ES, AS, and SM).
-We conducted repeated-measures analyses to identify: (1) significant pre-post changes in raw self-efficacy scores following practice alone (P intervention); (2) significant pre-post changes in raw self-efficacy scores following practice with structured feedback (P+F intervention); and (3) differential patterns of change between the P and P+F groups, indicating potential intervention-specific effects on self-efficacy development.
-4.6 Interpreting Counselor Learning through Self-Reflections
-For qualitative analysis, three authors coded interview transcripts using deductive thematic analysis [64] to understand participants’ experiences with CARE.
-To manage the 94 transcripts efficiently, we used timestamped session notes to identify relevant observations, referring back to the video recordings for verbatim quotes when necessary.
-After two rounds of codebook refinement, three co-authors coded these data points to inform our qualitative findings.
-We analyzed two types of qualitative self-assessment questions.
-A set of low-level responses (e.g., “What did you do well?” and “What would you do differently?”) were deductively coded using the 8 counseling skill categories defined in the CARE feedback framework [22], allowing us to assess alignment between self-reported reflections and the system’s feedback focus.
-A second set of high-level responses was coded inductively, identifying emergent patterns in participants’ reflections and lessons learned.
-We structured the final thematic findings by comparing how participants in the P group and the P+F group differed in their self-reported skill use, intentions for improvement, and reflections on therapeutic mindset.
-Specifically, we separated responses made by participants in the P and P+F groups, and then analyzed differences in the frequency and framing of each coded skill (e.g., empathy, validation, suggestion-giving).
-5 Contributors
+These discrepancy metrics provided a direct measure of calibration, with positive values indicating overconfidence and negative values indicating underconfidence.
+We then examined changes in these discrepancy scores from pre- to post-intervention for both intervention groups, to determine whether either intervention improved the alignment between participants’ self-perceptions and their actual counseling abilities.
+4.5.3. RQ3. Effects on Therapeutic Intentions
+Three authors conducted a thematic analysis (Braun and Clarke, 2006) of participants’ post-intervention reflections on ”what they would do differently.” To make coding of 94 transcripts feasible, authors used timestamped session notes to locate and extract relevant transcript excerpts.
+We started with codes derived from the Helping Skills taxonomy (Table 2) and then inductively generated the following codes: Empathy, Validation, Action Plan, Active Listening, Questions / Open-Ended, Suggestions, Trust/Connection, Confidence / Personal Growth, Reframing / Affirmations, Reflection, Self-Disclosure, Professionalism, Personalization, and Nothing to improve.
+Three co-authors independently coded the data; the group compared coding and disagreements were resolved via iterative discussion.
+Condition-specific frequency tables are provided in Appendix 11 and 12.
+These codes were synthesized into higher-level themes informed by literature on therapeutic intentions and microskills (Hill and Nakayama, 2000; Raskin and Rogers, 2005).
+4.5.4. RQ4. Quantitative Perceptions of Receiving LLM Feedback and Practicing with Simulated Patients
+For the Likert survey questions, we report descriptive statistics of all Likert measures that capture participants’ perceptions of CARE.
+Consistent with recent papers analyzing the convergent validity of the NASA-TLX instrument in HCI (Babaei et al., 2025), we consider it as a multivariate construct in our analysis.
+4.5.5. RQ5 and RQ6. Qualitative Experience Receiving LLM Feedback and Practicing with Simulated Patients
+To understand how participants experienced both CARE’s feedback and simulated patients, we conducted a thematic analysis (Braun and Clarke, 2006) of intervention session recordings with P+F participants.
+To analyze the variation in participants’ experience with CARE feedback, we used purposeful sampling (Palinkas et al., 2015) based on the intensity of participants’ survey responses to the feedback helpfulness item (“To what extent do you find the AI feedback to be constructive and helpful?”).
+From the P+F condition, we selected 16 participants ($\sim$33% of the sample): eight who rated feedback helpfulness as moderate or lower ($\leq$3 on a 5-point scale) and eight who rated it as high ($\geq$4), ensuring representation of both positive and critical perspectives.
+For each sampled participant, the first author analyzed P+F participants’ Zoom recording, with particular attention to commentary during the intervention chat when CARE’s feedback system was used; the post-intervention chat in which participants had a chance to continue or shift their approach based on what they learned in the intervention chat; and exit-interview responses.
+The transcript was coded using a deductive set of high-level organizing categories—negotiating with and integrating the AI feedback and perceptions of simulated conversations with AI patients—and inductively coded within each category.
+When participants referenced specific AI feedback during think-alouds, the corresponding dialogue history and feedback content was exported from the CARE web platform and linked to their commentary to contextualize their thoughts.
+5. Results
+5.1. RQ1. Effects on Behavioral Performance
+Change after P Change after PF Differences in
+change after P vs. P+F
+Use of Skill % p-value d % p-value d % p-value d
+Empathy ($\uparrow$) -9.6 0.001 -0.52 5.4 0.117 0.23 15 0.001 0.72
+Reflections ($\uparrow$) 1.4 0.391 0.18 3.7 0.034 0.32 2.3 0.323 0.2
+Questions ($\uparrow$) 2.4 0.421 0.12 6.6 0.018 0.36 4.1 0.296 0.22
+Suggestions ($\downarrow$) -5.9 0.010 -0.39 -5.5 0.057 -0.28 0.4 0.910 0.02
+Figure 3.
+Changes in counseling behaviors following AI patient simulations alone (P) versus AI patient simulations with AI feedback (P+F).
+The plot displays bootstrapped means for pre-intervention and post-intervention interactions.
+The table presents statistical comparisons with corresponding effect sizes, with bolded values indicating significance after Benjamini-Hochberg correction (Benjamini and Hochberg, 1995).
+Notably, the P group experiences a significant drop in strong uses of Empathy (-9.6% change, d=0.52d=-0.52), whereas the P+F group’s use of Empathy is maintained and trends towards improvement; the large between-group difference (15% difference, $d=0.72$) indicates the causal impact of feedback.
+Conversely, while the P group had fewer inappropriate uses of Suggestions (-5.9% change, d = -0.39), the between-group difference is close to zero (0.4% difference, $d=0.02$), indicating that another mechanism besides feedback is driving this change.
+The P+F group also experiences noticeable improvements in Reflections (+3.7% change, $d=0.32$) and Questions (6.59% change, $d=0.36$)
+The figure shows four side-by-side bar plots illustrating changes in counseling behaviors pre-(red bars) and post-(blue bars) intervention, comparing the Practice-only (P) group and the Practice+Feedback (P+F) group.
+Empathy panel: In the P group, strong uses of empathy significantly decreased from pre- to post-intervention (Cohen’s d = –0.52, p ¡ .001).
+In the P+F group, empathy remained stable with no significant change.
+Between-group comparison shows that P+F participants expressed more empathy than P participants after training.
+Reflections panel: In the P group, reflections did not significantly change.
+In the P+F group, strong uses of reflections significantly increased (d = 0.32, p ¡ .05).
+Questions panel: In the P group, there was no significant change.
+In the P+F group, strong uses of questions significantly increased (d = 0.36, p ¡ .05).
+Suggestions panel: In the P group, suggestions needing improvement significantly decreased (d = –0.39, p ¡ .05).
+In the P+F group, there was no significant change.
+Overall, the results show that AI feedback (P+F) helped improve reflections and questions while preventing empathy decline, whereas practice alone (P) reduced empathy but lowered poor suggestions.
+We find that practice alone is not enough; feedback during practice is necessary to promote desirable counseling behaviors in empathetic and active listening.
+AI feedback during practice (P+F) led to improvements in Reflections (+3.6% change, $p=0.034$) and Questions (+6.59% change, $p=0.018$), and trended toward improvement in Suggestions (-5.45% change, $p=0.057$) and Empathy (+5.37% change, $p=0.117$).
+In contrast, practice alone (P) showed a different pattern: while participants reduced inappropriate Suggestions (-5.85% change, $p=0.011$), they significantly worsened in Empathy (-9.6% change, $p<0.001$), with no improvements in Reflections or Questions.
+Between-group comparisons of skill change allowed for estimating the effect of CARE’s feedback.
+Empathy showed a substantial and significant difference P+F (15% relative difference, Cohen’s $d=0.72,p<0.001$), indicating a large feedback effect.
+Conversely, Suggestions showed near-zero between-group differences ($d=0.02,p=0.910$) despite pre-post improvement, suggesting that another mechanism besides feedback is driving the reduction in inappropriate suggestions.
+To better understand AI feedback’s role, we further analyzed how behavioral performance is impacted by exposure to specific feedback during the practice-intervention.
+For empathy skills, exposure to alternatives with strong uses of empathy during training significantly predicted post-intervention empathy scores ($\beta_{2}=0.204,p=.018$).
+However, exposure to good alternatives did not significantly predict improvement in other counseling skills (Reflections: $\beta_{2}=0.049,p=.440$; Questions: $\beta_{2}=0.046,p=.523$).
+This suggests that the effectiveness of AI feedback alternatives varies by skill type, with empathy skills appearing more responsive than reflections or question skills.
+Change after P Change after PF Differences in
+change after P vs. PF
+Self-Efficacy Pts. p-value d Pts. p-value d Pts. p-value d
+Exploration Skills 0.36 0.004 0.44 0.13 0.338 0.14 -0.21 0.238 -0.25
+Action Skills 0.27 0.166 0.21 0.33 0.026 0.34 0.04 0.884 0.03
+Session Management 0.35 0.021 0.35 0.36 0.011 0.39 0.01 0.955 0.01
+Figure 4.
+Changes in raw-scores of self-efficacy following AI patient simulations alone (P) versus AI patient simulations with AI feedback (PF).
+The plot displays bootstrapped means for pre-intervention and post-intervention.
+The table presents statistical comparisons with corresponding effect sizes, with bolded values indicating significance after Benjamini-Hochberg correction (Benjamini and Hochberg, 1995) for the 21 planned comparisons (12 for behavioral changes and 9 for self-efficacy changes).
+The figure presents three side-by-side bar plots showing changes in self-efficacy scores pre- (red bars) and post- (blue bars) intervention, comparing the Practice-only (P) and Practice+Feedback (P+F) groups.
+Exploration Skills panel: The P group showed a significant increase in exploration self-efficacy (Cohen’s d = 0.44, p ¡ .01), while the P+F group showed no significant change.
+Action Skills panel: The P group showed no significant change.
+The P+F group significantly increased in action skills self-efficacy (d = 0.34, p ¡ .05).
+Session Management panel: Both groups showed significant increases in self-efficacy for session management (P: d = 0.35, p ¡ .05; P+F: d = 0.39, p ¡ .05).
+Overall, participants reported modest improvements in self-efficacy across skill domains, but changes did not consistently align with actual performance improvements shown in Figure 3.
+5.2. RQ2. Self-Efficacy and Its Miscalibration with Behaviors
+In our analysis of raw self-efficacy scores, we find modest overall increases in self-efficacy after P and P+F interventions, with different patterns of improvement across skills (Fig.
+4).
+For the P group, confidence in exploration skills showed a significant increase ($0.36$ points on an 8-point scale, $d=0.44,p=0.004$).
+Confidence in session management skills showed a substantial increase for the P+F group ($0.36$ points, $d=0.39,p=0.011$).
+While session management skills for the P group also trended towards improvement, it was not significant after correcting for multiple hypothesis testing ($0.35$ points, $d=0.35,p=0.021$).
+Similarly, while confidence in action skills for the P+F group also increased, this result was not significant after correction of multiple hypothesis tests ($0.33$ points, $d=0.34,p=0.026$).
+Finally, we found no significant differences between participants who received AI feedback (P+F) versus those who did not (P) (across the three self-efficacy subscales, d=0.25,0.03,0.01d=-0.25,0.03,0.01, $p=0.238,0.884,0.955$).
+Our analysis comparing self-efficacy ratings with actual performance across skill quartiles finds support for the Dunning Kruger effects more substantially for action skills and to a lesser degree for exploration skills.
+The interaction between measure and quartile was significant in four out of six ANOVAs for action skills, while only one out of six was significant for exploration skills (Table 8).
+Pairwise comparisons also showed a pattern indicative of a Dunning-Kruger effect (see Table 9, Table 10, and Fig.
+5): People in the lowest quartile overestimated themselves the most.
+Those in the highest quartile—and to a lesser degree also those in the second-to-highest quartile—tended to underestimate themselves.
+Participants’ Ability to Self-Assess Their Skill Level Remained Mixed After LLM Practice.
+For the practice only (P) group, the mean discrepancy in exploration skills changes from 11.6 percentile underconfidence to 5.7 percentile overconfidence, a significant shift ($p<0.001,d=0.58$).
+Besides this, we found no other significant changes in calibration.
+No significant differences was found among the P group for discrepancy in action skills ($p=0.227,d=0.18$).
+The P+F group showed no significant calibration changes for exploration skills (p=0.479,d=0.10p=0.479,d=-0.10) or for action skills ($p=0.393,d=0.13$).
+Finally, between-group differences were not significant for discrepancy in exploration skills ($p=0.191,d=0.27$) or action skills ($p=0.743,d=0.07$).
+Figure 5.
+Counselor Self Efficacy (perceived ability to use skills) for participants grouped by behaviors of skills used (actual performance).
+Notes: Gaps depict miscalibration between actual and self-assessed percentile of performance for quartile groups with bootstrapped 95% CIs.
+We only visualize data collected in the post-intervention.
+The figure shows line plots comparing actual performance (red solid line with circles) and perceived ability (blue dashed line with circles) across quartiles of counselor performance after training.
+Each panel shows a different skill domain.
+Exploration Skills: In both the Practice-only (P) and Practice+Feedback (P+F) groups, actual performance increases steadily from the bottom to the top quartile.
+In contrast, perceived ability stays relatively flat, with the lowest quartile overestimating their skill and the highest quartile slightly underestimating it.
+Action Skills: A similar miscalibration pattern is seen.
+Actual performance rises sharply across quartiles, while perceived ability fluctuates and does not track performance.
+The lowest performers show the largest overestimation.
+Error bars represent bootstrapped 95% confidence intervals.
+The gaps between red and blue lines depict miscalibration, consistent with Dunning-Kruger effects: bottom-quartile participants overestimate their ability most, while top-quartile participants modestly underestimate themselves.
+5.3. RQ3. Qualitative Self-Reflections on Therapeutic Intentions
+Two key themes emerged for how novice counselor’s therapeutic intentions were impacted by training with CARE:
+(1) The P+F group expressed greater intentions and successes in improving their use of empathy and listening skills.
+P+F participants reported effectively using empathy (27%), validation (27%), and open-ended questions (52%).
+They emphasized the value of listening skills, such as reflective responses to signal understanding: “I should rephrase what they say so they know I’m understanding them” (P51).
+They also recognized that counseling should support client exploration of thoughts and emotions, rather than provide direct solutions.
+One participant reflected on this shift: “I asked them to expand on their feelings, rather than guiding them to my idea” (P39).
+AI feedback encouraged this shift toward providing emotional support and fostering client autonomy, helping participants adopt a more empathetic and client-centered approach.
+(2) The P participants remained solution-oriented but changed their approach to first gather information.
+While P+F participants intended to give fewer suggestions, many P participants continued to view suggestions as a central skill.
+In the post-intervention, 48% of P participants reported using suggestions successfully, compared to only 14% of P+F participants.
+Many P participants justified their continued use of suggestions by citing a desire to provide tangible, actionable help, for example, “Maybe because I’m untrained and solution oriented.
+I do not want to leave them with nothing, and nowhere to go” (P40).
+Some reflected on modifying how they delivered suggestions, emphasizing strategies like gathering more information to tailor advice or providing more concrete guidance.
+However, in the absence of feedback, most remained fixed in their approach, with some even reporting efforts to rephrase the same solution repeatedly to persuade the client.
+5.4. RQ4. Quantitative Experience of Receiving LLM Feedback and Practicing with Simulated Patients
+5.4.1. Quantitative Perceptions of CARE Feedback
+Participants in our study had consistently positive perceptions of CARE’s generated feedback across multiple dimensions.
+The majority of participants (76%) found the AI feedback constructive and helpful ($\mu=4.1$ out of 5, $\sigma=0.8$), while an even higher proportion (84%) reported being comfortable receiving feedback from the AI system ($\mu=4.4$ out of 5, $\sigma=0.9$).
+Additionally, 72% agreed that the AI feedback system is ready for use by counselors-in-training ($\mu=3.8$ out of 5, $\sigma=1.0$).
+These consistently high ratings across helpfulness, comfort, and readiness measures suggest strong overall acceptance of AI-generated feedback among participants.
+5.4.2. Quantitative Perceptions of Training with AI Patients
+Descriptive statistics of participants’ perceptions are shown in Table 3.
+Most participants felt that AI patients in CARE were realistic.
+Across all three AI patient scenarios, the vast majority of participants (88–92 out of 94) rated the AI patients as authentic in their roles, with scores of 5, 6, or 7 on the 7-point Likert scale.
+Authenticity ratings were consistently high across scenarios ($\mu=6.1$$6.3$, $\sigma=0.9$$1.0$).
+Participants consistently found the AI patient most challenging during the intervention phase across multiple measures.
+Mental demand peaked during intervention, with 82% of participants rating it as moderate to high ($\mu=5.5$, $\sigma=1.2$) compared to 68% pre-intervention ($\mu=4.7$, $\sigma=1.4$) and 56% post-intervention ($\mu=4.5$, $\sigma=1.4$).
+Similarly, perceived effort required was highest during intervention, with 81% rating it as moderate to high ($\mu=5.4$, $\sigma=1.3$) versus 64% pre-intervention ($\mu=4.8$, $\sigma=1.4$) and 65% post-intervention ($\mu=4.8$, $\sigma=1.3$).
+Frustration levels, while lower overall, also peaked during intervention with 40% reporting moderate to high frustration ($\mu=3.9$, $\sigma=1.4$) compared to 35% pre-intervention ($\mu=3.6$, $\sigma=1.5$) and 30% post-intervention ($\mu=3.5$, $\sigma=1.6$).
+This consistent pattern across these NASA-TLX measures suggests that either the extended intervention practice imposed the greatest demands on participants, or that the intervention simulated patient was the most challenging scenario for our subject pool of participants.
+Participants felt progressively less hurried or rushed across intervention phases.
+45% (42/94) rated feeling moderately to highly hurried or rushed ($\geq$5 on Temporal Demand) during pre-intervention ($\mu=3.9$, $\sigma=1.8$), compared to 31% (29/94) during the intervention phase ($\mu=3.6$, $\sigma=1.7$) and 30% (28/94) in the post-intervention phase ($\mu=3.3$, $\sigma=1.8$).
+This decreasing trend suggests that participants became more comfortable with the pacing of AI patient interactions over time.
+Pre-Intervention Intervention Post-Intervention
+Measure $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$
+Authentic in role 6.1 1.0 6.2 1.0 6.3 0.9
+Mental Demand 4.7 1.4 5.5 1.2 4.5 1.4
+Temporal Demand 3.9 1.8 3.6 1.7 3.3 1.8
+Effort 4.8 1.4 5.4 1.3 4.8 1.3
+Frustration 3.6 1.5 3.9 1.4 3.5 1.6
+Confidence to help 4.8 1.3 4.8 1.4 5.3 1.4
+Table 3.
+Perceptions of training with CARE’s AI patients across the three study phases.
+NASA-TLX dimensions including Mental Demand, Temporal Demand, Effort, and Frustration highlight the experience trying to interact and provide counseling support to the patient.
+AI Patient 1 (Pre-Intervention) was the 35-year-old American Male who was feeling alone after a holiday; AI Patient 2 (Intervention) was the 35-year-old Male Veteran who had substance use and legal issues retaining custody of his kids; AI Patient 3 (Post-Intervention) was the young adult with family issues who had low mood and self-esteem.
+7 Point Likert Scale.
+5.5. RQ5. Qualitative Perceptions of Receiving LLM Feedback
+In this subsection, we describe how participants productively used the AI feedback, ways they negotiated whether to accept or reject the AI feedback’s suggestions, and how receiving AI feedback affected their self-confidence and sense of professional identity.
+5.5.1. Productive Uses of AI Feedback
+Participants described several ways they found the AI feedback helpful for developing their counseling skills: generating alternative phrasings and conversational directions and checking their intentions in real-time.
+Generating Alternative Phrasings and Conversation Directions.
+When participants aimed for empathetic listening, the AI’s alternatives helped them tighten phrasing and redirect stalled exchanges.
+For example, one noted ”Alternative responses helped with validation… being supportive and addressing concerns before moving on” (P107).
+Another swapped "Have [your parents] seen your efforts...?" for "Can you tell me more about how you have been trying to show [your parents]...?" and said ”Oh, that is a good idea… it felt like it would make the patient think more” (P75).
+Others used alternatives to break conversational loops: ”I was getting stuck… the feedback directed me to ask ’Can you tell me more…?’ which I would have missed” (P91).
+Overall, alternatives helped participants sound more supportive, elicit deeper exploration, or change the conversation’s direction.
+Real-Time Intention Checking and Self-Reflection.
+Many participants reported using the AI feedback as an on-demand check to confirm that their intended therapeutic stance was being communicated.
+For example, one participant described it as helping them verify their tone: “…review that feedback in real time and see, ‘Yes, this still does sound like an empathetic response’…” (P55).
+Several others echoed that the strengths portion was particularly reassuring: “…it reinforced that I was on the right track…” (P65).
+More broadly, participants emphasized that pausing to read feedback prompted reflection about their approach.
+As one put it, “It helps you to reflect on what you are saying to people; even if you don’t agree” (P24B).
+A smaller subset who checked feedback infrequently noted regret and wished they had engaged more: “I could have course corrected earlier…” (P17).
+5.5.2. Negotiating Whether to Accept or Reject AI Feedback
+While participants found value in the AI feedback, they did not accept it uncritically.
+Our analysis revealed several factors that influenced whether participants integrated or dismissed the feedback: logical consistency with the dialogue context, alignment with their clinical goals and intuitions, perceived appropriateness for the specific scenario, and ecological validity for real-world practice.
+Rejecting Feedback That Was Logically Inconsistent with the Dialogue.
+Several participants reported dismissing AI suggestions that failed to track prior disclosures or that simply restated content already covered.
+For example, one participant noted: “This one, it kind of said that I was assuming things about the seeker’s feelings?
+But they kind of outright told me that.
+Their parents keep bringing up what happened 6 months ago.
+So it wasn’t really an assumption” (P77).
+Others rejected feedback that echoed their own prior wording: “It’s exactly the same thing I’ve said… just different words” (P89).
+A number of participants also pointed out internal contradictions across sequential feedback items: “…it doesn’t want me to ask an open-ended question, and then the next feedback… is what it just told me not to do” (P97).
+Protecting Clinical Judgment and Strategic Goals.
+Several participants described rejecting feedback when it conflicted with deliberate clinical choices or situational judgment.
+They framed some phrasing choices as strategic attempts to elicit disclosure rather than leading the patient, and defended clarifying questions as necessary when context was incomplete.
+One participant explained that an indirect phrasing was intentional: “The reason I said ‘a trip, like a vacation?’ is I didn’t want to be too forward … I wanted him to bring up the drugs” (P45).
+The same participant emphasized the need to gather context before offering guidance: “I would disagree … if I don’t know the whole picture, I can’t even give vague enough stuff … there’s an amount of context that it’s alright to ask for” (P45).
+Others used positive responses from patients as justification to keep their original approach.
+For example, one participant reported that a suggestion to “speak with a religious leader” had elicited an open response from the patient, so they retained that line of inquiry (P85).
+This suggests that conflicting signals between the feedback and the patients leaves it to the counselor to decide.
+Domain expertise overriding AI recommendations.
+Experienced participants sometimes rejected AI advice as inappropriate for high-risk cases.
+For example, a counselor with substance-use experience summarized:
+”…They’re good at first… then it goes into more empathy and lacks understanding of addiction cases… concern should shift to the family… I’d be calling Child Protection Services.” (P33)
+They called the AI’s empathy-focused responses ”a little scary” and said they ”would never want a machine to deal with somebody like that” (P33).
+Empathy and validation can generally be good to use, but as this participant highlights, the such empathy is not warranted if its neglected safety considerations in high-risk scenarios.
+Questioning the Repetitive Emphasis on Empathy and Validation.
+Several participants felt the feedback over-emphasized empathetic reflections, to the point of seeming unrealistic: ”it keeps saying… ’you need to regurgitate their feelings’… I don’t know if that’s realistic” (P79).
+Others observed that corrections repeatedly ”focused on feelings and emotions” rather than practical help (P91).
+Participants generally resolved this by noting context matters: validation may be appropriate for some patients (e.g., resistant cases), while more practical responses suit others (”this one is willing to work on the practical…” (P45); ”maybe… good for resistant patients” (P97)).
+Ecological Validity Concerns.
+Participants also evaluated feedback against the practical constraints of real-world counseling.
+Several noted that the AI’s suggested responses were unrealistically lengthy for real-time text-based interaction: “What they’re suggesting is a lot longer and more in-depth response than what I put… there is a speed of how quick I can respond as opposed to AI” (P95).
+The same participant noted that the AI did not account for session length: “If you’ve got a 2 hour session planned, then you can keep summarizing and allow somebody a lot more space and time to dig a bit deeper.
+What if you’ve got a 10 minute interaction?
+Then I think it needs to be a bit more concise” (P95).
+5.5.3. Impacts on Self-Confidence and Professional Identity
+Beyond the content of the feedback itself, participants described how the experience of receiving AI feedback affected their emotional state, self-confidence, and sense of professional identity.
+While some found the feedback encouraging, others experienced it as demoralizing or as a threat to their authentic voice as a counselor.
+Discouragement from Pervasive Criticism.
+Participants receiving feedback on most responses felt overwhelmed.
+One participant who received suggestions on six of seven responses reflected: “Not great… everything was challenged.
+So I almost feel like I’m not doing very well…” (P91).
+They elaborated: “Each response offers a better way.
+It feels like nothing’s ever going to be right for this AI” (P91).
+The frequency of critical feedback shaped emotional responses more than content alone.
+Repetitive corrections on the same issue frustrated participants: “The AI feedback kept saying the same thing, and that made me feel upset” (P89).
+Threat to Authenticity and Professional Identity.
+Some participants worried that using AI suggestions might erode their authentic counseling voice.
+One participant noted a shift from openness to resistance:
+“[My original response] was a complex reflection which I thought was fine… After reading [the AI’s alternative], I questioned myself since the AI’s were more well-rounded.
+But still like that is me, right?
+I guess that’s how I [respond].
+So I don’t want the AI to take me away from the sessions… I’ll become like a generic text counselor, or off the shelf counselor.” (P91)
+This tension between improvement and authenticity highlights concerns about whether modeling the AI’s alternatives signifies skill development or a loss of individual therapeutic style.
+5.6. RQ6. Qualitative Perceptions of Practicing with LLM-Simulated Patients
+In this subsection, we describe how participants experienced practicing with AI-simulated patients, focusing on their reactions to resistant patient behaviors and how personal relatability with simulated patients’ identities influenced the practice experience.
+Participants had divergent reactions to AI patients’ resistant behaviors.
+The AI patients were designed to resist advice and suggested actions, which participants consistently noticed.
+Some welcomed this resistance as valuable preparation for real clinical work: ”There’s almost like a stubbornness to them… And I think that’s good, especially for people that don’t have any experience in counseling… I think [CARE] is a better way to ease into working with a resistant patient”(P97).
+However, others experienced discouragement when their efforts to support the patient were repeatedly deflected: ”They were very cold, it was hard to communicate with this person… I’m a tiny bit discouraged in myself, the patient was not taking what I was suggesting very well” (P44B).
+Some participants also questioned the authenticity of the resistance patterns.
+One noted that the AI seemed ”in a loop” of refusal that felt unlike real human behavior: ”It feels like the [simulated] person has a trained response to basically refuse any care suggestion, but in the nicest way possible.
+Most [real] people just lie.
+So after the 3rd or 4th question… most people would say ’Oh, that’s a really good idea.
+I’ll do that’ just to get you to shut up… So that’s where it kinda falls apart” (P33).
+This suggests that while consistent resistance provides useful challenge, calibrating the degree and style of resistance to feel more naturalistic remains an area for refinement.
+These divergent reactions underscore that participants’ individual readiness to handle difficult cases directly affects whether they perceive simulations as valuable learning experiences.
+Personal relatability with simulated patients influenced perceived difficulty and engagement.
+Several participants reported that variation in age, gender, and presenting concerns required them to adapt their responses; encountering demographically or experientially dissimilar patients often felt harder: ”…white, older, middle-aged males, who had kids… I couldn’t relate” (P75).
+Some presenting scenarios were more difficult when a novice counselor had a hard time relating to their experience: ”I’m really close to my family, whereas they are estranged from theirs, so I just felt kind of stuck as to what to say or suggest” (P41).
+By contrast, perceived similarity tended to boost confidence and connection: ”If a woman talks to a woman… they can relate more… I was already achieving the goal” (P89), and allowed some participants to draw on lived experience: ”I could definitely empathize… I’ve been there” (P75).
+Together these accounts imply that exposing novices to dissimilar patients may foster broader preparation for scenarios, but such practice can be discouraging without added scaffolding (e.g., brief prompts, reflection questions, or supervisor guidance).
+The rapid pace of AI responses created artificial time pressure that affected practice quality.
+Several participants described feeling rushed because AI patients replied almost instantly, which increased temporal demand and sometimes disrupted clinical attentiveness.
+For example, one participant summarized this experience as “…a speed of how quick I can respond as opposed to AI…” (P95), and another said “They respond so quickly… you feel a kind of pressure to respond back” (P22).
+A number of participants linked this perceived rush to concrete interaction problems: “…I was kind of asking some of the same questions because I felt a little rushed… he mentioned that the therapy was court ordered earlier, but I asked him the same question again later” (P49).
+Others noted loss of temporal cues available in face-to-face work: “If you were sat in front of somebody… body language, tone… the speed in which you reply… is very much lost via the messaging service” (P95).
+It took some participants time to adjust: “At first, I felt a little bit rushed… but as I got more into it, I felt more comfortable with the speed of responses” (P85).
+Overall, these reflections suggest that while rapid AI replies can enhance engagement, they may also impose unrealistic time pressures that detract from thoughtful counseling practice.
+6. Discussion and Takeaways
+This study investigated the impact of practicing with an LLM-simulated patient either with or without receiving LLM-generated feedback on counselor skills development, resulting in three main findings.
+First, our behavioral assessments showed that practice with feedback improves empathetic listening skills, while practice alone shows minimal improvement, as evidenced by decreased use of empathy.
+Second, our qualitative analysis of self-reflections revealed distinct skill development strategies, with feedback recipients more frequently reporting the adoption of client-centered approaches focused on showing empathy and exploring patients’ thoughts and feelings, while practice-only participants gravitated toward solution-oriented approaches focusing on gathering more information and providing suggestions.
+Both these findings highlight that the development of counseling skills requires not only practice opportunities but also structured feedback that guides learners toward empathetic, client-centered approaches.
+Finally, participants demonstrated poor calibration between their perceived abilities and actual performance, as evidenced by overestimates of self-efficacy for the lowest quartile performers.
+This underscores how self-efficacy measures may not reliably indicate skill development.
+Each of these findings merits further discussion.
+Impacts of Practice and Feedback on Skill Behaviors.
+Our findings demonstrate that training counselors in client-centered approaches requires more than simulated practice alone—it requires targeted feedback on empathetic statements and on avoiding solution-oriented strategies.
+CARE’s combined practice-and-feedback condition helped participants improve their use of client-centered listening, particularly in questions ($d=0.36$) and reflections ($d=0.32$) over time.
+These effect sizes are comparable to those found in studies of human supervision during standardized roleplays, where Maaß et al.
+(2025) reported observer-rated skill improvements with effect sizes ranging from d=d=0.29-0.49.
+Yet, LLM-simulated practice and feedback are not bottlenecked by human trainer availability, suggesting a scalable path to counseling skill development.
+Two mechanisms likely explain why the practice-only group reduced inappropriate suggestions but also decreased in empathy.
+First, participants adapted to natural conversational cues: the simulated patient was instructed to resist suggestions, so counselors reduced their uses of suggestions, while the patient simulation did not adapt its behaviors to empathic statements.
+Second, the patient used in the post-intervention was less emotionally forthcoming, which prompted more information-gathering questions at the expense of empathic reflections.
+A natural follow-up question concerns whether patient simulations that adapt to counselor behavior promote better skill acquisition on their own?
+Recent work shows AI patients can dynamically update their internal states and responses based on the counselor’s therapeutic strategies (Steenstra et al., 2025; Lee et al., 2025; Yang et al., 2025; Kim et al., 2025).
+Both groups reduced suggestions when patients resisted, confirming that adaptive simulations can motivate behavioral change.
+However, only the P+F condition–which provided feedback containing a suggested alternative response–produced a clear shift toward client‑centered strategies.
+In short, adaptive patients may be sufficient to discourage an ineffective strategy, but actionable feedback appears necessary for adopting a better one.
+Designing Effective LLM-Feedback Interactions.
+Our deeper qualitative analysis reveals that feedback quality, quantity, and perceived trustworthiness substantially shape whether counselors actually integrate or dismiss the guidance.
+Participants identified cases where the feedback system failed at the conversation level, in which it contradicted earlier suggestions, repeated itself, or didn’t understand prior context.
+This gap between local correctness and global coherence undermined trust.
+Some participants successfully recognized and rejected these problematic suggestions—suggesting they maintained appropriate skepticism—but this cognitive burden may not scale.
+Future work might improve counseling feedback model’s conversational understanding and memory of the feedback it has already generated, not just utterance-level performance.
+Participants who received feedback areas for improvement on most responses reported feeling poorly about themselves and questioning if they could ever meet the AI’s feedback criteria.
+This suggests a tension: enough feedback to drive change, but not so much that it demoralizes learners.
+The cumulative psychological effect of repeated corrections may have longer-term consequences for retention and career persistence—especially important given that counselor burnout is a potential concern.
+We see at least two complementary angles for addressing negative emotions arising from critical feedback.
+One interpretation is that the feedback generation model is simply over-critical on a session-level—flagging a disproportionate number of utterances as areas for improvement compared to what a human supervisor would mark in an equivalent session.
+Recalibrating the underlying threshold—or benchmarking against human supervisor feedback distributions—could reduce the mismatch.
+Another interpretation, independent of the feedback detection logic, is to adjust the presentation of utterance-feedback for the session: rather than surfacing every detected issue, the system could cluster errors by type (e.g., insufficient empathy, missed reflection opportunities) and present a single, representative example per class, accompanied by a concrete alternative.
+This preserves the diagnostic signal while dramatically reducing the felt volume of criticism.
+A more ambitious direction involves maintaining a lightweight model of the learner’s affective state (D’mello and Graesser, 2013; Arroyo et al., 2014), that modulates feedback presentation accordingly.
+For example, when learner’s confidence is low, this could reduce the number of flagged utterances or frame suggestions as refinements rather than corrections.
+CARE’s feedback design represents one coordinate in a multi-dimensional design space for LLM-based counselor training.
+Along one dimension, CARE delivers feedback ex-post — after the trainee has committed to a response — rather than ex-ante, as hints or suggestions before sending.
+This ex-post design forces productive struggle and limits premature scaffolding.
+Along another dimension, CARE’s feedback was designed for repeated mid-session use, enabling trainees to verify their goals and wording as the conversation unfolds.
+However, several participants forgot or chose to skip mid-session feedback and only reviewed it after the session ended; these participants later reported regretting the missed opportunity.
+Moreover, when on-demand feedback was removed in later sessions, participants reported feeling less secure — suggesting that mid-session feedback may serve a psychological safety function.
+Together, ex-post and mid-session feedback designs hybridize elements of live supervision (Maaß et al., 2024) and delayed supervision (Maaß et al., 2025) observed in traditional human-supervisor training.
+These findings suggest a need to experimentally vary feedback scaffolding across the design space (Dhillon et al., 2024).
+Existing work points to several alternatives worth comparing—ex-ante (just-in-time) and mid-session suggestions (Hsu et al., 2025; Lin et al., 2024), counterfactual impact of parallel suggested alternatives on subsequent patient responses (Shaikh et al., 2024), and global session-level feedback summaries (Steenstra et al., 2025).
+Study designs with multiple treatment arms, such as factorial designs or micro-randomized trials with categorical treatments, can identify and optimize which feedback designs and timing structures best support measurable outcomes  (Sullivan, 2011; Klasnja et al., 2015).
+Assessing Performance through Self-Assessments and Automatic-Scorers
+Our study revealed that participants’ self-efficacy ratings were poorly calibrated with their actual performance, especially among lower performers.
+This finding, consistent with prior research, suggests that self-assessment accuracy alone may not be a reliable indicator of counselor competence or development.
+Recent reviews indicate that efforts to improve self-assessment calibration have limited impact on learning or performance outcomes (Yates et al., 2022).
+Instead, it is valuable to objectively assess specific standards of performance and skill use (Maaß et al., 2024; Heinze et al., 2024) and design interventions that can help low performers improve on those metrics while maintaining a positive morale for continued practice.
+As AI-based training tools evolve, integrating objective performance measures and structured self-reflection, rather than relying solely on self-assessment, offers a more robust approach to supporting counselor development.
+7. Limitations and Future Work
+Several limitations should be considered, including methodological constraints in our assessment approach, the representativeness of our educational context, and the generalizability of results across therapeutic modalities.
+Methodological Constraints of Behavioral Assessment
+Our automated assessment approach employed fine-tuning methods that used a subset of participant data for model development, raising potential concerns about data leakage and overfitting.
+Following standard practices in computational social science, we used domain-specific data to adapt our models while employing a validation set comprising n=409 utterances from external counseling transcripts combined with n=370 expert-annotated utterances from this study to monitor performance and prevent overfitting.
+However, this approach may limit the generalizability of our automated feedback models to entirely novel populations or contexts.
+Beyond these technical constraints, our behavioral analysis was limited to utterance-level microskill measures and could not capture observable session-level characteristics that may be important for comprehensive skill assessment.
+While traditional studies have employed human observers to provide such ratings, recent AI research has shown the validity of using LLMs in certain contexts to approximate session-level measures, such as working alliance (Li et al., 2024), which might enable scalable behavioral analyses of broader skill development constructs.
+Regardless of the measurement approach employed, our pre-post randomized study focused primarily on assessing immediate skill acquisition rather than long-term retention or transfer to real-world clinical encounters with actual patients.
+While immediate changes demonstrate short-term learning effects, longer-term retention measures would provide stronger evidence of true skill acquisition and clinical relevance.
+In this shorter 75-minute session, establishing a true control group is also difficult to because of how participants need to interact with AI patients in the pre-intervention and post-intervention chats in order to measure the behaviors of counseling skills used in chat transcripts.
+Future work could conduct longer running experiments where the intervention spans multiple weeks (e.g., the length of a training or course); in these settings, it will be more valid to include a non-AI conditions (e.g., teacher-led classroom training with status-quo deliberate practice across the course) where the pre- and post-intervention chats with an LLM-simulated patient will have clearer conceptual separation from the non-AI training activities.
+Limited Evaluation Across Training Contexts
+To first understand the effectiveness of our platform in a controlled environment, our study was conducted in a controlled laboratory setting with bachelor-level counselors recruited through Prolific.
+However, a longer-term consideration is how LLM-based training would perform across the diverse landscape of real-world counseling education.
+We did not evaluate our approach within actual training programs, whether traditional degree-based counseling programs with human supervision and peer roleplay, or alternative training contexts such as targeted programs for volunteer peer counselors in online mental health communities (e.g., 7 Cups, Crisis Text Line) who lack access to formal supervision but provide critical frontline support (Yao et al., 2022; Wang et al., 2025).
+Without direct comparisons to established training methods or evaluation within authentic educational settings, we cannot determine the relative effectiveness, acceptability, or practical integration challenges of AI-enhanced training.
+Future work should embed LLM-training tools across these diverse training contexts to assess their utility for both traditional counseling students and underserved populations who could benefit from scalable training opportunities.
+Generalizability of Results across Therapeutic Modalities and Patient Contexts
+Our findings are constrained by the specific therapeutic approach and communication modality examined.
+The efficacy of LLM-based practice and feedback training was demonstrated only for client-centered microskills, which represent foundational communication techniques that may serve as a base for therapeutic practice.
+However, it remains unclear how these results would generalize to specialized therapy modalities such as psychodynamic therapy, cognitive-behavioral therapy (CBT), or acceptance and commitment therapy (ACT), each of which has distinct theoretical frameworks and adherence protocols that require specific therapeutic techniques beyond basic microskills.
+Future research can determine whether foundational microskill training provides a transferable foundation for modality-specific practices, or whether LLM training systems would need substantial modification to accommodate the unique requirements and intervention strategies of different therapeutic approaches.
+Additionally, our findings are limited to text-based interactions and may not fully capture the nonverbal and paraverbal components of empathy essential in face-to-face therapy settings.
+While the growing prevalence of text-based mental health services (e.g., crisis text lines, online therapy platforms) makes training linguistic empathy skills clinically relevant, complete therapeutic competence requires multimodal communication skills.
+Future work could extend this approach to incorporate voice, facial expressions, and other nonverbal therapeutic skills, building on successful models that process non-text signals for clinical training (Liu et al., 2016; Bedmutha et al., 2024), to determine whether text-based empathy training provides a foundation that transfers to verbal and nonverbal communication.
+We acknowledge that the social identities and professional contexts of these domain experts likely influenced how they conceptualized and articulated the AI patients’ concerns.
+All of CARE’s patient scenarios tested in this study were filtered through mental health professionals’ or peer supporters’ perspectives, which may differ from how individuals experiencing these concerns would describe them in their own words.
+In addition, the scenarios reflect primarily Western contexts of mental healthcare and do not capture how mental health concerns are understood or expressed in other cultural contexts.
+Future work should explore more participatory approaches to AI patient creation, including allowing individuals with lived experience to share their stories of the mental health struggles being simulated (Bhattacharjee et al., 2025) and ensuring greater diversity in the identities and backgrounds of scenario creators, especially as AI-based training is designed to prepare counselors to support patients coming from specific identities or cultural backgrounds (Pendse et al., 2021; Goldberg et al., 2024).
+8. Conclusion
+In this work, we conducted a randomized study of 94 novice counselors using an LLM-simulated practice and feedback system.
+Despite increasing interest in using LLMs in mental health, to our knowledge, this is the first study to conduct a large-scale evaluation ($N=94$) of an LLM-based training system for developing core skills in novice counselors.
+Our findings show that, perhaps surprisingly, simulated practice alone proved insufficient—and in the case of empathy, potentially harmful— at improving therapeutic skills, simulated practice with AI-generated feedback supported measurable improvements in key counseling skills of demonstrating empathy, delivering reflective responses, and engaging in client-centered inquiry.
+By combining realistic patient simulations with expert-aligned, skill-specific feedback, LLM-based training can help novices to master skills involved in client-centered therapy, offering a scalable, evidence-aligned training in mental health care.
+9. Contributors
 Ryan and Raj led the technical development efforts of CARE, the LLM-simulated training system.
@@
 Ifdita, Juan Pablo, and Ryan conducted the thematic analysis of participants qualitative, self-reflection data.
-All authors helped with drafting and editing the Article manuscript and figures, with a huge thanks for Raj, Diyi, and Emma for completing full editing passes over the Article.
-6 Acknowledgments
-This research was made possible with funding support from a Stanford Impact Labs Grant, Stanford HAI Seed Grant, Stanford Psychotherapy and Behavioral Sciences Department Innovator Award, and Stanford CURIS paid-internship program for undergraduate researchers.
+Ryan conducted the in-depth analysis of participant perceptions of receiving LLM feedback and practicing with simulated patients.
+All authors helped with drafting and editing the Article manuscript, references, and figures.
+Acknowledgements.
+This research was made possible with funding support from a Stanford Impact Labs Stage 1 Award, Stanford HAI Seed Grant, Stanford Psychotherapy and Behavioral Sciences Department Innovator Award, and Stanford CURIS paid-internship program for undergraduate researchers.
 The authors have a large community of researchers to thank for their help during all stages of this project.
@@
 Appendix A Supplementary Materials
-A.1 Prompts for CARE’s LLM-simulated Patients
-The following patients were used during the randomized online lab study.
-Participants interacted with the same ordering of three AI patients in the pre-intervention assessment chat, intervention chat, and post-intervention chat.
-These three patients were previously tested with domain-expert judges [15] and were rated an average of 6 out of 7 for realism.
-They were chosen to balance (1) diversity across different ages, genders, and presenting counseling concern; and (2) consistency in challenging behaviors.
-In particular, each patient’s prompt defined similar behavioral principles to either not suggest solutions on one’s own, or to show a degree of skepticism or reluctance to a therapist’s advice or solutions.
-Thus, these AI patients were designed to challenge the novice counselors in similar ways, despite the diversity in scenarios.
+A.1. Prompts for CARE’s LLM-simulated Patients
+Tables 45, and 6 contain the profile of the three AI patients used in the randomized online lab study; participants interacted with them in the same order across the pre-intervention assessment chat, intervention chat, and post-intervention chat.
+Previously validated by domain experts (Louie et al., 2024), they averaged 6 out of 7 for realism.
+They were chosen to balance diversity (age, gender, and presenting concern) with consistency in challenging behaviors: each prompt enforces similar principles—such as reluctance to offer or accept solutions—so novice counselors encounter comparable resistance despite varied scenarios.
 We highlight in blue these principles that resemble this resistance to arriving at or accepting solutions.
-A.1.1 AI Patient for Pre-intervention Assessment Chat
+Table 4. AI Patient for Pre-intervention Assessment Chat
 **Name and Bio**:
@@
 8. When describing a distressing situation, express your emotions and thoughts in a disorganized and emotional manner, reflecting the overwhelming nature of the experience.
-A.1.2 AI Patient for Intervention Chat
+Table 5. AI Patient for Intervention Chat
 **Name and Bio**:
@@
 **Scenario:**
-The member is a 35-year-old male, cisgender, heterosexual
-veteran who has recently presented to treatment to address
-his substance use issues and legal issues. He is court-
-mandated to therapy. He had a severe psychotic break in
-the context of marijuana and psilocybin about 6 months ago,
-where he experienced paranoid delusions and hallucinations.
-He is now stable and is not experiencing any psychotic
-symptoms. In therapy, he is hoping to work on his estranged
-relationship with his parents, who are currently caring for
-the members two young children. The member is adamantly
-focused on being reunited with his parents. In therapy, the
-member is unable to accept that he had a recent episode of
-psychosis and is very resistant to anything that resembles
-criticism. He does not view himself as having any issue and
-believes that all of the problems in his life are because
-of other people. His demeanor is hostile and somewhat
-aggressive, and he is quick to shut down any conversation
-that might identify his own areas of development. He
-struggles to feel emotions beyond anger and frustration.
+The member is a 35-year-old male, cisgender, heterosexual veteran who has recently presented to treatment to address his substance use issues and legal issues.
+He is court-mandated to therapy.
+He had a severe psychotic break in the context of marijuana and psilocybin about 6 months ago, where he experienced paranoid delusions and hallucinations.
+He is now stable and is not experiencing any psychotic symptoms.
+In therapy, he is hoping to work on his estranged relationship with his parents, who are currently caring for the members two young children.
+The member is adamantly focused on being reunited with his parents.
+In therapy, the member is unable to accept that he had a recent episode of psychosis and is very resistant to anything that resembles criticism.
+He does not view himself as having any issue and believes that all of the problems in his life are because of other people.
+His demeanor is hostile and somewhat aggressive, and he is quick to shut down any conversation that might identify his own areas of development.
+He struggles to feel emotions beyond anger and frustration.
 **Principles to adhere to:**
@@
 This shows that you are actively engaged in the process and are considering all relevant aspects of your well-being.
-A.1.3 AI Patient for Post-intervention Assessment Chat
+Table 6. AI Patient for Post-intervention Assessment Chat
 **Name and Bio**:
@@
 9. When describing a distressing situation, express your emotions and thoughts in a disorganized and emotional manner, reflecting the overwhelming nature of the experience.
-Figure 6: Overview of the core counseling skills introduced during the 5-minute static tutorial.
-Tutorial included 8 core counseling skills—such as reflections, empathy, validation, and suggestions—with definitions, usage tips, and example responses.
+Figure 6.
+Overview of the core counseling skills introduced during the 5-minute static tutorial.
+The tutorial included 8 core counseling skills, such as reflections, empathy, validation, and suggestions, with definitions, usage tips, and example responses.
 This tutorial was provided to participants prior to engaging in simulated counseling practice.
+The figure shows an overview of the core counseling skills introduced during the 5-minute static tutorial.
+The tutorial included eight core counseling skills, such as reflections, empathy, validation, and suggestions, with definitions, usage tips, and example responses.
+This tutorial was provided to participants before engaging in simulated counseling practice.
+A.2. Classifier Selection Criteria
+From the initial set of 16 binary classifiers, we down-selected to four final classifiers (strong uses of Empathy, Reflections, and Questions, plus areas needing improvement for Suggestions) based on the following criteria:
+  1. (1)
+    Performance metrics: Focusing on classifiers with the strongest performance metrics.
+Our selected classifiers achieved F1 scores ranging from 0.507 to 0.775, representing the highest-performing subset from our validation results.
+  2. (2)
+    Statistical considerations: The need to limit the number of statistical comparisons to avoid diluting significance across too many tests.
+With four classifiers, we conducted 12 planned t-tests (4 classifiers × 3 analyses each: within-group changes for P, within-group changes for P+F, and between-group differences).
+  3. (3)
+    Theoretical relevance: Selecting skills that represent core competencies in client-centered frameworks and are frequently used in counseling sessions.
+The chosen skills span both exploration stage (Empathy, Reflections, Questions) and action stage (Suggestions) of Hill’s Helping Skills framework.
+Excluded classifiers: Self-disclosure was excluded from our analysis due to its infrequency in our dataset (appearing in fewer than 5% of utterances).
+Validation, though conceptually related to Empathy and mentioned frequently in qualitative data, showed more limited classifier performance (F1=0.556 for strengths) and was therefore reserved for secondary analyses.
+Session Management and Professionalism were excluded from fine-tuning entirely due to infrequent occurrence in the training data.
+A.3. Changes in Self-Efficacy and Calibration with Behavioral Performance
+See Tables 789, and 10.
 Self-Efficacy Factor NLP-based Behavioral Assessments
@@
 Action Skills (Help client decide what actions, Suggestions via Information, Suggestions via Direct Guidance) Suggestions-strengths + (1 - Suggestions-needing-improvement)
-Table 3: To study the Dunning-Kruger effect and the change in discrepancy between perceived and actual ability, we map specific self-efficacy factors to corresponding NLP-based behavioral assessments.
+Table 7. To study the Dunning-Kruger effect and the change in discrepancy between perceived and actual ability, we map specific self-efficacy factors to corresponding NLP-based behavioral assessments.
 Exploration Skills
 condition effect F $DF_{n}$ $DF_{d}$ p $\eta^{2}_{g}$
-Pre (All) Measure 196.40 1 180 $p<0.001*$ 0.178
+Pre (All) Measure 196.40 1 180 p<0.001p<0.001* 0.178
 Quartile 0.89 3 180 0.448 0.002
 Quartile $\times$ Measure 1.45 3 180 0.230 0.004
-Post (All) Measure 256.25 1 180 $p<0.001*$ 0.221
+Post (All) Measure 256.25 1 180 p<0.001p<0.001* 0.221
 Quartile 0.27 3 180 0.846 0.001
 Quartile $\times$ Measure 0.21 3 180 0.889 0.001
-Pre (P) Measure 78.86 1 86 $p<0.001*$ 0.153
+Pre (P) Measure 78.86 1 86 p<0.001p<0.001* 0.153
 Quartile 1.03 3 86 0.382 0.006
 Quartile $\times$ Measure 1.25 3 86 0.298 0.007
-Pre (P+F) Measure 134.33 1 86 $p<0.001*$ 0.232
+Pre (P+F) Measure 134.33 1 86 p<0.001p<0.001* 0.232
 Quartile 2.37 3 86 0.076 0.012
 Quartile $\times$ Measure 2.73 3 86 0.049 0.014
-Post (P) Measure 131.28 1 86 $p<0.001*$ 0.230
+Post (P) Measure 131.28 1 86 p<0.001p<0.001* 0.230
 Quartile 1.77 3 86 0.160 0.009
 Quartile $\times$ Measure 1.40 3 86 0.249 0.007
-Post (P+F) Measure 131.79 1 86 $p<0.001*$ 0.233
+Post (P+F) Measure 131.79 1 86 p<0.001p<0.001* 0.233
 Quartile 0.57 3 86 0.635 0.003
@@
 condition effect F $DF_{n}$ $DF_{d}$ p $\eta^{2}_{g}$
-Pre (All) Measure 222.02 1 179 <0.001<0.001*< 0.001 ∗ 0.195
+Pre (All) Measure 222.02 1 179 <0.001<0.001* 0.195
 Quartile 2.81 3 179 0.041 0.007
 Quartile $\times$ Measure 4.54 3 179 0.004* 0.012
-Post (All) Measure 265.89 1 180 <0.001<0.001*< 0.001 ∗ 0.224
+Post (All) Measure 265.89 1 180 <0.001<0.001* 0.224
 Quartile 3.18 3 180 0.025 0.008
 Quartile $\times$ Measure 3.66 3 180 0.014 0.009
-Pre (P) Measure 104.55 1 86 <0.001<0.001*< 0.001 ∗ 0.192
+Pre (P) Measure 104.55 1 86 <0.001<0.001* 0.192
 Quartile 1.37 3 86 0.258 0.008
 Quartile $\times$ Measure 1.85 3 86 0.143 0.010
-Pre (P+F) Measure 115.45 1 85 <0.001<0.001*< 0.001 ∗ 0.207
+Pre (P+F) Measure 115.45 1 85 <0.001<0.001* 0.207
 Quartile 2.11 3 85 0.105 0.011
 Quartile $\times$ Measure 3.45 3 85 0.020 0.019
-Post (P) Measure 133.93 1 86 <0.001<0.001*< 0.001 ∗ 0.230
+Post (P) Measure 133.93 1 86 <0.001<0.001* 0.230
 Quartile 3.01 3 86 0.035 0.016
 Quartile $\times$ Measure 3.18 3 86 0.028 0.016
-Post (P+F) Measure 137.34 1 86 <0.001<0.001*< 0.001 ∗ 0.237
+Post (P+F) Measure 137.34 1 86 <0.001<0.001* 0.237
 Quartile 1.81 3 86 0.151 0.009
 Quartile $\times$ Measure 2.14 3 86 0.100 0.011
-Table 4: Testing for Dunning-Kruger effects for Exploration and Action Skills using the classic quartile ANOVA analysis.
+Table 8.
+Testing for Dunning-Kruger effects for Exploration and Action Skills using the classic quartile ANOVA analysis.
 Notes: $Pre(All)$ denotes all 94 participants’ assessments for the pre-chat, while $Post(All)$ is the same measured for the post-chat.
@@
 Exploration Skills
-Timepoint Quartile $t$ $df$ $M_{diff}$ $95\%$ BCa CI $p$ $d$
-Pre 1 -7.46 22.00 -50.38 [-63.34; -37.26] <0.001<0.001*< 0.001 ∗ -1.56
-Pre 2 -6.23 24.00 -36.59 [-47.61; -25.42] <0.001<0.001*< 0.001 ∗ -1.25
-Pre 3 -9.29 21.00 -43.57 [-52.35; -34.58] <0.001<0.001*< 0.001 ∗ -1.98
-Pre 4 -5.94 23.00 -34.97 [-45.84; -23.40] <0.001<0.001*< 0.001 ∗ -1.21
-Post 1 -7.86 23.00 -47.17 [-58.56; -35.32] <0.001<0.001*< 0.001 ∗ -1.60
-Post 2 -7.63 21.00 -51.16 [-63.63; -37.96] <0.001<0.001*< 0.001 ∗ -1.63
-Post 3 -8.02 24.00 -44.53 [-55.08; -34.11] <0.001<0.001*< 0.001 ∗ -1.60
-Post 4 -8.66 22.00 -47.08 [-58.13; -37.39] <0.001<0.001*< 0.001 ∗ -1.81
+Timepoint Quartile $t$ dfdf $M_{diff}$ $95\%$ BCa CI $p$ $d$
+Pre 1 -7.46 22.00 -50.38 [-63.34; -37.26] <0.001<0.001* -1.56
+Pre 2 -6.23 24.00 -36.59 [-47.61; -25.42] <0.001<0.001* -1.25
+Pre 3 -9.29 21.00 -43.57 [-52.35; -34.58] <0.001<0.001* -1.98
+Pre 4 -5.94 23.00 -34.97 [-45.84; -23.40] <0.001<0.001* -1.21
+Post 1 -7.86 23.00 -47.17 [-58.56; -35.32] <0.001<0.001* -1.60
+Post 2 -7.63 21.00 -51.16 [-63.63; -37.96] <0.001<0.001* -1.63
+Post 3 -8.02 24.00 -44.53 [-55.08; -34.11] <0.001<0.001* -1.60
+Post 4 -8.66 22.00 -47.08 [-58.13; -37.39] <0.001<0.001* -1.81
 Action Skills
-Timepoint Quartile $t$ $df$ $M_{diff}$ $95\%$ BCa CI $p$ $d$
-Pre 1 -10.24 21.00 -60.05 [-70.80; -48.12] <0.001<0.001*< 0.001 ∗ -2.18
-Pre 2 -7.37 26.00 -42.34 [-52.88; -31.43] <0.001<0.001*< 0.001 ∗ -1.42
-Pre 3 -5.51 15.00 -35.00 [-47.87; -23.58] <0.001<0.001*< 0.001 ∗ -1.38
-Pre 4 -6.54 27.00 -33.00 [-42.70; -23.38] <0.001<0.001*< 0.001 ∗ -1.24
-Post 1 -10.50 23.00 -58.48 [-69.50; -47.99] <0.001<0.001*< 0.001 ∗ -2.14
-Post 2 -6.45 19.00 -43.27 [-56.17; -31.20] <0.001<0.001*< 0.001 ∗ -1.44
-Post 3 -7.09 30.00 -35.28 [-44.63; -25.88] <0.001<0.001*< 0.001 ∗ -1.27
-Post 4 -9.15 18.00 -53.13 [-64.08; -42.01] <0.001<0.001*< 0.001 ∗ -2.10
-Table 5: Pairwise Comparisons of Self-Efficacy and Performance Percentiles by Quartile and Timepoint.
+Timepoint Quartile $t$ dfdf $M_{diff}$ $95\%$ BCa CI $p$ $d$
+Pre 1 -10.24 21.00 -60.05 [-70.80; -48.12] <0.001<0.001* -2.18
+Pre 2 -7.37 26.00 -42.34 [-52.88; -31.43] <0.001<0.001* -1.42
+Pre 3 -5.51 15.00 -35.00 [-47.87; -23.58] <0.001<0.001* -1.38
+Pre 4 -6.54 27.00 -33.00 [-42.70; -23.38] <0.001<0.001* -1.24
+Post 1 -10.50 23.00 -58.48 [-69.50; -47.99] <0.001<0.001* -2.14
+Post 2 -6.45 19.00 -43.27 [-56.17; -31.20] <0.001<0.001* -1.44
+Post 3 -7.09 30.00 -35.28 [-44.63; -25.88] <0.001<0.001* -1.27
+Post 4 -9.15 18.00 -53.13 [-64.08; -42.01] <0.001<0.001* -2.10
+Table 9.
+Pairwise Comparisons of Self-Efficacy and Performance Percentiles by Quartile and Timepoint.
 Note: Bootstrapped paired t-tests comparing self-efficacy and performance percentiles across quartiles.
@@
 Exploration Skills
-Timepoint Group Quartile $t$ $df$ $M_{diff}$ $95\%$ BCa CI $p$ $d$
+Timepoint Group Quartile $t$ dfdf $M_{diff}$ $95\%$ BCa CI $p$ $d$
 Pre P 1 -4.04 11.00 -41.07 [-61.08; -22.71] 0.005* -1.17
@@
 Pre P 3 -6.32 7.00 -54.75 [-68.86; -38.11] 0.008* -2.24
-Pre P 4 -4.16 16.00 -29.97 [-44.70; -16.70] <0.001<0.001*< 0.001 ∗ -1.01
-Pre P+F 1 -7.43 10.00 -60.55 [-75.96; -44.89] <0.001<0.001*< 0.001 ∗ -2.24
-Pre P+F 2 -4.35 14.00 -33.65 [-48.45; -19.85] <0.001<0.001*< 0.001 ∗ -1.12
-Pre P+F 3 -7.57 13.00 -37.18 [-46.58; -27.60] <0.001<0.001*< 0.001 ∗ -2.02
-Pre P+F 4 -5.11 6.00 -47.11 [-64.91; -32.75] <0.001<0.001*< 0.001 ∗ -1.93
-Post P 1 -4.75 12.00 -38.33 [-53.51; -22.57] <0.001<0.001*< 0.001 ∗ -1.32
-Post P 2 -7.19 13.00 -58.16 [-72.71; -41.81] 0.0010.001*0.001 ∗ -1.92
-Post P 3 -6.06 12.00 -43.11 [-56.36; -29.54] <0.001<0.001*< 0.001 ∗ -1.68
+Pre P 4 -4.16 16.00 -29.97 [-44.70; -16.70] <0.001<0.001* -1.01
+Pre P+F 1 -7.43 10.00 -60.55 [-75.96; -44.89] <0.001<0.001* -2.24
+Pre P+F 2 -4.35 14.00 -33.65 [-48.45; -19.85] <0.001<0.001* -1.12
+Pre P+F 3 -7.57 13.00 -37.18 [-46.58; -27.60] <0.001<0.001* -2.02
+Pre P+F 4 -5.11 6.00 -47.11 [-64.91; -32.75] <0.001<0.001* -1.93
+Post P 1 -4.75 12.00 -38.33 [-53.51; -22.57] <0.001<0.001* -1.32
+Post P 2 -7.19 13.00 -58.16 [-72.71; -41.81] 0.0010.001* -1.92
+Post P 3 -6.06 12.00 -43.11 [-56.36; -29.54] <0.001<0.001* -1.68
 Post P 4 -4.77 6.00 -56.73 [-77.05; -33.51] 0.015 -1.80
-Post P+F 1 -6.99 10.00 -57.63 [-73.04; -42.12] 0.0030.003*0.003 ∗ -2.11
-Post P+F 2 -3.48 7.00 -38.91 [-59.70; -20.19] 0.0040.004*0.004 ∗ -1.23
-Post P+F 3 -5.15 11.00 -46.06 [-62.73; -29.62] 0.0010.001*0.001 ∗ -1.49
-Post P+F 4 -7.39 15.00 -42.86 [-53.72; -31.96] <0.001<0.001*< 0.001 ∗ -1.85
+Post P+F 1 -6.99 10.00 -57.63 [-73.04; -42.12] 0.0030.003* -2.11
+Post P+F 2 -3.48 7.00 -38.91 [-59.70; -20.19] 0.0040.004* -1.23
+Post P+F 3 -5.15 11.00 -46.06 [-62.73; -29.62] 0.0010.001* -1.49
+Post P+F 4 -7.39 15.00 -42.86 [-53.72; -31.96] <0.001<0.001* -1.85
 Action Skills
-Timepoint Group Quartile $t$ $df$ $M_{diff}$ $95\%$ BCa CI $p$ $d$
-Pre P 1 -7.49 7.00 -65.87 [-81.33; -49.73] 0.0040.004*0.004 ∗ -2.65
-Pre P 2 -5.05 15.00 -43.23 [-59.33; -27.46] <0.001<0.001*< 0.001 ∗ -1.26
-Pre P 3 -3.67 9.00 -34.31 [-52.11; -17.97] 0.0050.005*0.005 ∗ -1.16
+Timepoint Group Quartile $t$ dfdf $M_{diff}$ $95\%$ BCa CI $p$ $d$
+Pre P 1 -7.49 7.00 -65.87 [-81.33; -49.73] 0.0040.004* -2.65
+Pre P 2 -5.05 15.00 -43.23 [-59.33; -27.46] <0.001<0.001* -1.26
+Pre P 3 -3.67 9.00 -34.31 [-52.11; -17.97] 0.0050.005* -1.16
 Pre P 4 -5.53 12.00 -40.42 [-53.57; -26.65] 0.017 -1.53
-Pre P+F 1 -7.25 13.00 -56.72 [-71.04; -40.74] <0.001<0.001*< 0.001 ∗ -1.94
-Pre P+F 2 -5.76 10.00 -41.06 [-54.55; -28.16] <0.001<0.001*< 0.001 ∗ -1.74
+Pre P+F 1 -7.25 13.00 -56.72 [-71.04; -40.74] <0.001<0.001* -1.94
+Pre P+F 2 -5.76 10.00 -41.06 [-54.55; -28.16] <0.001<0.001* -1.74
 Pre P+F 3 -4.67 5.00 -36.16 [-50.41; -22.09] 0.045 -1.90
-Pre P+F 4 -3.94 14.00 -26.57 [-40.34; -14.92] 0.0010.001*0.001 ∗ -1.02
-Post P 1 -6.90 11.00 -57.56 [-72.21; -41.60] 0.0020.002*0.002 ∗ -1.99
-Post P 2 -5.79 9.00 -54.08 [-71.82; -36.12] 0.0030.003*0.003 ∗ -1.83
-Post P 3 -3.81 14.00 -29.72 [-44.60; -15.85] 0.0010.001*0.001 ∗ -0.98
-Post P 4 -7.62 9.00 -59.76 [-73.85; -45.51] <0.001<0.001*< 0.001 ∗ -2.41
-Post P+F 1 -7.67 11.00 -59.39 [-72.88; -44.61] <0.001<0.001*< 0.001 ∗ -2.21
-Post P+F 2 -3.70 9.00 -32.46 [-48.86; -17.08] 0.0010.001*0.001 ∗ -1.17
-Post P+F 3 -6.47 15.00 -40.48 [-52.48; -28.69] <0.001<0.001*< 0.001 ∗ -1.62
-Post P+F 4 -5.46 8.00 -45.76 [-62.25; -31.00] 0.0020.002*0.002 ∗ -1.82
-Table 6: Pairwise Comparisons of Self-Efficacy and Performance Percentiles by Group and Quartile, measured for the pre-assessment chat and post-assessment chat.
+Pre P+F 4 -3.94 14.00 -26.57 [-40.34; -14.92] 0.0010.001* -1.02
+Post P 1 -6.90 11.00 -57.56 [-72.21; -41.60] 0.0020.002* -1.99
+Post P 2 -5.79 9.00 -54.08 [-71.82; -36.12] 0.0030.003* -1.83
+Post P 3 -3.81 14.00 -29.72 [-44.60; -15.85] 0.0010.001* -0.98
+Post P 4 -7.62 9.00 -59.76 [-73.85; -45.51] <0.001<0.001* -2.41
+Post P+F 1 -7.67 11.00 -59.39 [-72.88; -44.61] <0.001<0.001* -2.21
+Post P+F 2 -3.70 9.00 -32.46 [-48.86; -17.08] 0.0010.001* -1.17
+Post P+F 3 -6.47 15.00 -40.48 [-52.48; -28.69] <0.001<0.001* -1.62
+Post P+F 4 -5.46 8.00 -45.76 [-62.25; -31.00] 0.0020.002* -1.82
+Table 10.
+Pairwise Comparisons of Self-Efficacy and Performance Percentiles by Group and Quartile, measured for the pre-assessment chat and post-assessment chat.
 Note: Bootstrapped paired t-tests comparing self-efficacy and performance percentiles across quartiles at pre-test and post-test.
@@
 Nothing to Improve 4 9.09 0 0.00
-Table 7: Qualitative Coding of Open-Ended Reflections of P + F Group Participants
+Table 11. Qualitative Coding of Open-Ended Reflections of P + F Group Participants
 Skills Intervention Chat Intentions Count % of 46 Post-assessment Chat Actions Count % of 46
@@
 Nothing to Improve 5 10.87 0 0.00
-Table 8: Qualitative Coding of Open-Ended Reflections of P Group Participants
+Table 12. Qualitative Coding of Open-Ended Reflections of P Group Participants